A Bibliography of Publications in JavaWorld Magazine

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

19 September 2010 Version 1.05

Abstract

This bibliography records publications in JavaWorld magazine, a Web journal of Java news.

Title word cross-reference

#36 [Dal97]. #37 [Wis97].

$100 [Ste96g].

000 [Ste96f].

1 [Di 96c]. 1.0 [Pet98]. 1.1 [Mit96l]. 1.2 [Ano97-28]. 1/1/97 [Sca97n]. 1/15/97 [Sca97m]. 10/01/97 [Sca97d]. 10/1/96 [Sca96a]. 10/1/98 [Sca98k]. 10/15/97 [Sca97e]. 10/15/98 [Sca98j]. 11/01/97 [Sca97f]. 11/15/97 [Sca97g].

11/15/98 [Sca98l]. 11/2/98 [Sca98m]. 12/15/96 [Sca97o]. 12/18/98 [Sca99]. 12/2/97 [Sca97h]. 12/2/98 [Sca98n]. 15 [Ano96d]. 1996 [Ano96d]. 1997 [Ano96e]. 1Java [Di 97a].

2 [DiG96]. 2.0 [CG98]. 2.1 [Hun98b]. 2.5 [Big98]. 2.6 [McM97n]. 2/1/97 [Sca97p]. 2/15/97 [Sca97i]. 25 [Fer99b]. 2D [Day98d, Day98h].

3 [McM96g]. 3.0 [McC96g]. 3.0.1 [Di 97a]. 3.1 [Law96c]. 3/1/97 [Sca97c]. 3/15/97 [Sca97q]. 3D [Sun97b].

4 [Kal96c]. 4/1/97 [Sca97s]. 4/1/98 [Sca98p]. 4/15/97 [Sca97r]. 4/15/98 [Sca98o]. 40-city [BG97]. 400 [Sca98a].

5 [Ano97j]. 5.0 [Syk97]. 5/1/97 [Sca97t].

5/15/97 [AS97]. 5/15/98 [Sca98y]. 5/4/98 [Sca98q].

6 [Ano97k]. 6.0 [MT98a]. 6/1/97 [Sca97v]. 6/15/97 [Sca97u]. 6/15/98 [Sca98z]. 6/2/98 [Sca98-27].

7/1/97 [Sca97x]. 7/15/96 [AS96d]. 7/15/97 [Sca97w]. 7/15/98 [Sca98r]. 7/2/98 [Sca98s]. 7/3/96 [Ano96o].

8/1/96 [Ste96a]. 8/15/97 [Sca97j]. 8/15/98 [Sca98t]. 8/3/98 [Sca98u]. 8/4/97 [Sca97y].

9/1/98 [Sca98w]. 9/15/96 [Sca96b]. 9/15/97 [Sca97k]. 9/15/98 [Sca98v]. 9/3/97 [Sca97l]. 9/5/96 [Ste96b]. 95 [Ano98n]. ’96 [Mil96a]. ’97 [Fre97, Ess97d]. ’98 [Day98i, MS98]. ’99 [Hol99b].

abilities [Spe98]. abstract [Van97g]. Accelerate [Rou98]. Access [Sma98]. acquisition [McM97j]. across [Hem97c]. action [Alo98b]. Activator [SM98]. ActiveX [Blu96a]. Add [Di 97a]. adding [McM97n]. admin [Mil96d]. Adobe [Ess96b]. adoption [Cul97]. advancements [Arm97c]. advantage [McC97c]. Advertising [Ano96a]. AFC [Bow98a]. again [Zuk98e]. against [Zuk97f]. Agents [Som97a]. aglets [Ven97k]. aims [Nic98k]. ain’t [Van98d]. Alan [Zuk98a, Ste98i]. algorithms [Van98a]. all-Java [Gar97b]. alliance [Kal96e]. along [Mat98b]. alpha [Day98b]. already [Mur97f]. alternative [Mit96f]. alters [McK97c]. America [Sch96a]. analysis [McM97f]. Analyst [Nic98a]. Andreessen [McK97d]. animation [vW96]. anniversary [O’C97]. announce [Ste96g]. announcements [Zuk98a]. announces [McC96a]. Annual [Ano96p]. answer [Mit96d]. answers [Di 97b]. Antialiasing [Day98b]. anymore [Som97a]. anyway [Mit98]. AOL [Gar98c]. API [Hun98b, Rya96a]. APIs [Law96f]. apostles [BG97]. app [WM97b]. appeals [Hei99]. Apple [Bus97e, Kal96e]. Apple/Microsoft [Ste97c]. applet [Mil96c]. AppletAuthor [Mur97b]. applets [vW96]. Appliances [Law96a]. application [Ess96c]. applications [Wes96]. Applied [Mor98a]. approach [RS97]. apps [Cli97]. April [Ano97f]. architecture [Ess96c]. Architectures [Hol98e]. archives [MC98b]. arena [McK97a]. arguments [Fri98b]. arithmetic [Ven96b]. ARM [Kal97d]. arrays [Sau98a]. article [Zuk97e]. articles [Sca97c]. AS/400 [Sca98a]. Asia [UG96]. asked [Mil98a]. Assure [Dye98]. asynchronously [Sun97f]. attention [Ste96f]. audio [WM97a]. augments [Bri98]. August [Ano96d]. authentication [MF97]. Authenticator [Zuk98c]. Authoring [Sha96b]. Automatic [Wes96]. Avoid [Zuk97a]. awarded [Hei99]. Awards [Ano96p]. AWT [Sun96d]. AWTEventMulticaster [Hol99b].

B [McM97p]. back [Ess97d]. Backed [Sca98-28]. backlash [Blu98a]. backwards [Sha96d]. bag [Joh98h]. Ballmer [Bus97b]. Bandwidth [Di 96a]. banks [Bow98c]. Baratz [Ste98i]. barring [MS98]. based [Kal96d]. basic [Har97]. basics [Ven96a, McM97d]. batch [Van98e]. battle [BC97]. be [McK97b]. bean [Joh97c]. BeanBox [Joh97a]. BeanLint [Joh98a]. beans [Sha96h, Law96d]. bears [Bow98c]. become [McM98i]. been [Cor97]. beginner [Sha96c]. beginning [Alo97b]. begins [Nic99]. behavior [Arm97a]. behind [Wis97]. beholder [Blu97g]. being [Sha96a]. benchmark [Nef97]. benders [Arm97c]. Bending [Sha96d]. benefits [Hun98a]. BeOS [Ste98d]. best [DT97]. Beta [Mur97d, Mur97f]. bets [BS98]. bevy [Hug97b]. beyond [Cor97]. bid [Bus97d]. big [Ess97d]. big-wigs [Ess97d]. bigger [Gre98]. bites [Mur98]. bitmap [WM97b]. bits [Zuk98b]. blabbermouth [McG97]. black [Wil98]. blend [Mur96a]. Blundon [Blu96b]. bonanza [Blu97e]. Bond [Som97a]. Book [Rya96b]. books [Van97a]. Boost [Avi98a]. boosts [McM97n]. Borland [McM97a]. bow [Fer99b]. Box [Ven98m]. boxes [Ste98h]. Branding [Sch96a]. breaks [Rod98]. breed [Arm98a]. breeze [Ste96e]. brew [Law96a, Nel98]. brews [Law96g]. briefs [O’C96b]. bring [Sha97]. brings [Law96c]. broadening [Blu97c]. browser [McG97]. bug [McC96h]. bugs [McK97c]. Build [McM97g]. Building [Kad96a]. built [McC96f]. built-in [McC96f]. Bull [Taa98]. bulls [Bow98c]. Bunny [LaG96b]. business [Cli97, Sid97a]. business-critical [Cli97]. businesses [Gar98c]. buttons [Kad97c]. buy [DD98]. Bytecode [Ven96a]. bytecodes [Mah97].

C [Mur96g]. CA [Fer97a]. cable [Con96]. Café [Big98]. caffeine [Gre97]. Calculating [MM98]. call [Mur96l]. called [McM97m]. Camp [Syk96b]. can [Blu96g]. cancels [Tro97b]. ’canonical [Ven98a]. cans [Di 96d]. capabilities [Mor98c]. Cappuccino [Gre96a]. Capture [Mes98b]. Caravelle [Syk97]. Card [San96, DM98]. cards [TJ97]. CardTerminal [WwRDGSB99]. care [Par97]. Career [Ano96n]. carefully [Kal96f]. Cascading [Mit96g]. Case [Kad96a]. Castanet [Wei98]. celebrates [O’C97]. center [Blu96e]. Central [Knu97a, Knu98g]. centric [Avi98d]. CGI [Dur97]. challenges [Fer99c]. change [Gut97a]. channel [SD98]. channels [McM96g]. character [McM98b]. Chart [Win97a]. chat [Hug97a]. cheap [Coo98c]. check [McM97l]. chip [SS97]. Chips [Fer98b]. chock [Mur97d]. Choice [Ano96p]. choices [Bow98a]. choose [Mak97]. Cisco [Sch97b]. city [BG97]. claims [Tro98b]. clarifies [Pet97a]. class [Ven96d]. classes [Kad96b]. clauses [Ven97l]. cleanup [Ven98g]. clicks [Mit96f]. client [Kad96b]. client-server [Kad96b]. client/server [Sho97a]. closer [Ros97]. Closing [Ven97a]. cloudy [McK97e]. co [Fer99a]. co-opetition [Fer99a]. code [Cha97d]. Code-centric [Avi98d]. coding [Ano98a]. coins [Som97b]. collaborate [Ess96c]. collaborative [Hug98d]. collected [Ven96c]. collection [McM96b]. collections [McM98a]. Column [Mat97b]. Comcast [Han98]. Comdex [Mil96d]. comes [ME96a]. coming [Ess97b]. Commentary [Day97a]. comments [Alo98h]. commerce [Gar98c]. commercial [Kad96a]. commercial-quality [Kad96a]. committee [Bus97d]. common [Hem97c]. communication [McM96g]. communications [Con96]. company [Ess97f]. comparative [Van97a]. compared [Dye97]. Competing [Ste97f]. competition [McK98a]. compiler [McM97n]. complement [SE98]. complete [Zuk97d]. completion [Ale97]. complex [WwRDGSB99]. comply [Nic99]. components [Mur96g]. compositing [Day98b]. composition [Ven98l]. Computer [Sun97b]. computers [Avi98f]. Computing [Sha96e]. Condition [Hol98d]. conference [McC96a, Zuk98b]. confident [Nic98l]. confront [Fer99c]. confusion [Bow98a]. conjures [Rad97b]. Connection [O’C96c]. cons [Jer96]. console [Len97]. constants [Arm97b, Ste98a]. Construction [Hag97]. ’constructor [Smi98]. container [Joh97a]. content [AO98]. contest [Val96]. contract [Bus97b]. Control [Ven97b]. controls [McM96f]. controversy [Bow97d]. conventions [Ano98a]. cookin [Kal96f]. cooking [SC97]. cools [UG96]. copy [Dub98a]. CORBA [McK97e]. Core [Van98e, Wei96]. Corel [Pet97a]. Corner [Blu96b]. corporate [Gre96a]. correctly [Rou99]. costs [Mil96d]. countersues [Fer97c]. countersuit [Fer97b]. Counting [Hol98d]. course [Lie96b]. court [Arm98b]. courtroom [Blu98e]. Cover [Van97e]. coverage [Owe98]. Create [Van97e]. creates [Bow98a]. Creating [vW96]. creator [Day98g]. critical [Cli97]. cross [Mil96b]. cross-platform [Mil96b]. crossware [Som97b]. crucial [Sun99c]. CSX [Bow97a]. Cup [Val96]. currency [Col98]. current [Avi98e]. custom [Win97b]. customization [Joh97c]. Customize [Joh97d]. customizing [Loy99]. cut [Dub98a]. cuts [Mil96d]. Cyberstruction [LaG96b].

daily [Ano96g, Ano98e]. dancing [Bow97c]. data [Van97g]. database [LaG96a]. databases [Sha96f]. datagram [Mit97a]. date [Zuk98a]. day [McC96a, Cli97]. days [Alo98g]. DB [Kri98c]. Deadlock [Hol98b]. deal [Rod98]. death [McM97m]. debate [Bow98a]. Debugging [McC96b]. debut [SO97].

Decaf [Joh97d]. decaffeinated [Avi98e]. December [Ano97a]. decide [Mur96l]. decides [Sca98-29]. decompilers [Dye97]. decouple [Lop97]. decry [Tro97a]. defects [Alo98d]. define [Gar98a]. defined [McC96c]. delegation [Hug97e]. deliver [Van97c]. delivered [Sun98j]. delivers [Rya96c]. Demand [Sca97b]. demands [Bow98a]. Demo [Mat97a]. demonstrated [Ven97l]. Department [Tro98a]. dependent [Arm97a]. Depth [McM96c]. design [Mit96l]. Designing [Ven98d]. desktop [Cor97]. despite [Avi98e]. detailed [Joh98f]. dev [Mur96b]. DevCon [Som97b]. develop [Hem97b]. Developer [Di 96d, O’C96c]. developers [Mur96a]. developing [MF98a]. Development [Mil96a, Day98k]. device [Avi98e]. devices [Ess97a]. Diba [McM97j]. dice [Van98a]. did [Mur96l]. different [Zuk98f]. differing [Nic98i]. difficulties [Mit96a]. Digging [Ste98b]. Direct [McM97m]. disappointments [Arm97c]. discloses [Kal96d]. discuss [Cul97]. discusses [Gut97b]. discussion [Sho97c]. disguises [Car98]. dispute [Kal97b]. distinguished [Som98]. Distributed [Sha96e]. Distributing [Som96a]. distribution [Wes96]. divinely [Alo98c]. division [McK98f, Nic98c]. Do [Joh98d]. does [Kal96b]. don’t [Kal97a]. double [Mit96f]. down [Coo98e]. download [Len97]. Draw [Sun97d]. draws [Kal96c]. dried [Joh98d]. drills [McK97a]. driver [Kri98c]. dropping [MT98b]. DS&A [Van98a]. duke [Arm98b]. dynamic [Sun98f, Kru98]. dynamically [Sun97c].

e-mail [Dal97]. earn [Mor98d]. ease [Mil96b]. easier [McC97a]. easily [Cio98a]. easy [Kad97e]. eciphering [Blu97b]. Editor [Ano96p, O’C96a]. Education [Alo97a]. Educators [Bow97b]. effect [Sch97a]. effective [Sun96c]. efforts [UG96]. EJB [Gar98a]. electronic [Sun97h]. elephant [Bow97c]. eliminate [Don98]. embedded [Mur96i]. embeds [Coo98b]. embrace [Bow97b]. empty [Wil98]. enable [Mit96f]. enabled [SS97]. enabling [Ste96d]. end [Sca98-29]. endorsement [Ste96c]. Energizer [LaG96b]. engine [San96]. engineering [Mah97]. engineering/ decompilation [Mah97]. engines [LaG96a]. English [Gre96b]. enhances [Kri98c]. ENIGMAtic [McM98d]. ensure [McK97c]. enterprise [Law96j, McK97f]. enters [DS97]. enumerated [Arm97b]. environment [DT97, Blu97f]. environments [Hem97c]. escalates [BC97]. Escape [Sma98]. etc [Hol98b]. ETrade [Bow98c]. Europe [Syk98]. Evangelist [Alo97c]. even [Ros97]. event [Sun96a]. events [Van97f]. ever [Rya96b]. Everything [Knu97a]. everywhere [Nic98g]. evolution [Con96]. examine [Gut99]. exceptions [van96]. exclusive [Por97]. exec [Joh97g]. execution [McM97c]. executive [Nic98e]. expand [Di 97d]. experience [Sca98-28]. experts [Wei96]. explain [O’C98]. Explicit [wJDM98]. Explorer [Blu97e]. Expo [Ste98f, Mur96b]. extend [Kad97c]. extends [Mor98c]. extensible [Sun97c]. extension [Ven99a]. Extensions [Ven98p, Ven99d]. extract [MC98b]. extranet [Blu97i]. extreme [Ste98g]. eye [Blu97g]. eyes [Sca98a].

faces [Blu97d]. fails [Van97c]. Farewell [Blu97k]. fast [Bel97]. fat [Owe98]. Feature [You96, Bow97c]. features [McC96g]. February [Ano97j]. feel [FRLS97]. feels [Nic98c]. fever [McM98f]. fields [Ven98f]. file [Ven96d]. files [WM97b]. finalization [Ven98g]. finally [Ven97l]. financial [Bow98d]. finds [Law96b]. finished [Pér98]. fire [Nic98h]. fires [LaG96a]. firewalls [Wis97]. firm [Ste96g]. first [McM97a]. first-ever [Nef97]. first-hand [McM98f]. fisted [GS99]. fistful [Joh98c]. flexibility [Big98]. Floating [Ven96b]. Floating-point [Ven96b]. flow [Ven97b]. Focus [McC98]. focused [Ven98f]. foe [Coo98d]. Follow [Zuk97c]. Follow-up [Zuk97c]. forces [Gar98c]. form [Cha97d]. formal [Fer99b]. forms [McC96d]. forthcoming [Zuk98f]. Fortune [Ste96f]. forum [Sho97c]. forward [Sid97a]. foundation [Gut97c]. Four [Mak97, Sca97a]. fractal [Hug98a]. Frames [Sha96g]. Framework [Day97d, DT98]. frameworks [Van99]. Francisco [Law96j, Ste98f]. freeware [McK98a]. freeze [Joh98d]. freeze-dried [Joh98d]. frequently [McC97b]. Friend [Sch97b]. front [Day98j]. full [Mur96h]. fulla [Mur97d]. fully [Avi98a]. function [Mes98a]. functionality [Mur96h]. functions [McC96c]. Fund [Ste96g]. further [McK97a]. fury [Smi97b]. future [Blu96c].

Gage [Gut97a]. gains [Cha97b]. game [Van97a]. Gang [Sca97a]. garbage [Ven96c]. garbage-collected [Ven96c]. gatekeeper [Bus97d]. Gates [Pér98]. gateway [Mur96g]. gems [Day98k]. generation [SC97]. generator [Ven98j]. Generic [Kad96b]. get [Hun97]. gets [Law96f]. getting [Kad97b]. gives [Cha97a]. Giving [Col98]. glue [Law96d]. GNU [Avi98c]. go [Syk96b]. goes [Mur96f]. good [Van97g]. Gosling [UG96]. gospel [Rya96c]. got [Van98d]. grab [Joh98h]. grant [Dug98]. granted [AN98]. graph [Win97a]. Graphic [Hug98a]. graphics [McC96e]. Grappling [Mil96c]. greater [Gar98b]. grinds [Mur97b]. grip [GS99]. ground [Hem97c]. Group [Ess97d]. grows [Day97c]. growth [Bow99]. GUI [LLZ98]. guide [Sha96c]. guru [Di 97b].

Habanero [Mur96c]. Half [Joh97d]. Hall [Mur96k]. hand [Bus97c]. handle [Sun97e]. handles [Ven97c]. handling [Sun96a]. hands [MB97]. hang [Sun97a]. Haven’t [McM97k]. head [Hug97c]. heading [Ste97d]. heap [Ven96c]. heard [McM97k]. hearing [Mil98a]. heat [FRLS97]. helping [Ste96e]. helps [Sun99d]. here [Ano98w]. heterogeneous [Hem97c]. Hewlett [Coo98a]. hierarchical [Mes98b]. high [Gre97]. high-caffeine [Gre97]. highlight [Apr97]. hill [Sha96h]. Hitachi [Ste98h]. holidays [MM98]. honors [Ste97e]. Hood [Ven96e]. hook [Lea96]. Horse [Pet97a]. Host [Sca97b]. hosted [Di 97a]. hot [Mil96a]. HotJava [DS97]. HotSpot [Arm98a]. house [Alo98g]. house-hunting [Alo98g]. How-to [Sun96a]. HP [Nic98j]. HTML [Syk98]. HTMLEditorKit [Loy99]. hunger [Di 96a]. hunting [Alo98g]. hype [UG96]. HyperTV [Mur96d].

I-commerce [Gar98c]. I/O [Zha97]. IBM [Kal96b]. iButtons [DD98]. IDC [Kal97a]. IDE [McM97a]. identify [Hun98a]. IDEs [Mak96]. IDG.net [Ano97a]. idiom [Ven98j]. if [Van98d]. II [McM96d]. ILOG [Sca98-28]. image [Kad97c]. images [Zuk96]. Immature [Spe98]. impact [Sch97a]. Implement [McM97c]. importance [Kad97b]. Impressions [Ven98n]. improve [Zha97]. Improvements [Blu97f]. in-depth [McM97i]. in-memory [Zuk96]. include [Gar97a]. incompatible [MS98]. Increase [Sho97a]. independence [Gut98a]. industry [ME96a]. Inferno [Law96g]. Infospace [Ess97c]. Inheritance [Ven98l]. initialization [Ven98d]. initializing [Sau98a]. initiatives [Woo97]. injunction [MS98]. inner [McM97b]. Inprise [Zuk98a, Zuk98b]. inroads [Syk98]. ins [Jer96]. insider [Sun97e]. insiders [RO97]. inspired [Alo98c]. install [Ven97f]. installation [Mit96a]. Instruction [WM97b]. integrated [Avi98a]. Integrating [Sha96f]. integration [MC98a]. Intel [Sch96c, Han98]. interactive [Con96]. interest [Hol98a]. interface [Sun96c]. interfaces [Kri98a]. International [Val96]. Internationalize [Sun98f]. Internet [Ess96b]. interoperability [MS97]. interpreter [McM97d]. Intertainer [Han98]. Interview [Ano96b]. Interviews [RO97]. intranet [Kal96e]. introduces [McC96h]. Introducing [Hun98b]. introduction [Gre96b]. intros [Hei96]. invest [Han98]. invitation [Kal97b]. invocation [Ven97d]. ISO [Bus97a]. issue [Bus97a]. issues [Cul97]. Iterator [Bis97]. itself [Mur96k]. ITxpo [FRLS97].

J [Mur97e, McM97m]. J/Direct [McM97m]. Jacob [Gre97]. January [Ano96f, Sch97a]. jar [Web98, MC98b]. Java [Di 96d]. Java-based [Kal96d]. Java-centric [Wei98]. Java-enabled [SS97].

Java-integrated [Avi98a]. Java-powered [McM98f]. java.lang.Object [Rou99]. Java/ C [Mur96g]. JavaBean [RB97]. JavaBeans [Mur97b]. JavaBeans-based [Gar97a]. JavaHeads [Alo97a]. Javamation [vW96]. JavaNFS [McM97k]. JavaOne [Ano96g]. JavaOS [ME96c]. JavaReady [Alo98a]. Java’s [Sca98b]. JavaScript [Sha96c]. JavaServer [Cha97d]. JavaSoft [Lie96b]. JavaSpaces [Gut97c]. Javatizes [Mur96k]. JavaWorld [Ano96n]. javax.comm [MD98]. Javits [Gre97]. Javitz [Mur96e]. jazzes [Coo98f]. JBuilder [McM97a]. JDBC [Sha96f]. JDK [Mit96l]. jeopardy [SL98]. jewelry [McM98f]. JFC [San98]. JFC/ Swing [Van98d]. JGL [Van97g]. Jini [Sha98a]. JIT [McM97n]. jive [Kal96b]. JNI [Fri98b]. Jobs [Ess96a]. Joe [Mur96j]. join [Kal97b]. joins [BG97]. jolts [Gre97]. JTree [Mes98b]. Judge [Mil98a]. juicy [Zuk98b]. July [Ano96e]. jumps [Kal96a]. jumpstart [DM98]. June [Ano97k]. junk [Sun97h]. just [Wei96]. Justice [Tro98a]. JViews [Sca98-28]. JVM [McM97n].

Kaiser [Gle98]. Keep [Joh97e]. key [Hun98a]. keynote [Zuk98a]. Keynotes [Rex97b]. keys [MF97]. Kid [Sch97b]. kill [Blu96f]. killer [Mat97a]. Kim [Ano96b]. kind [Avi98c]. Kit [Cha97e]. knew [Tro98b]. Kona [Law96a].

lab [MB97]. lack [McK97c]. land [Mat98b]. landscapes [Hug98a]. lane [Coo98e]. language [Kad96c]. last [Ano96d]. Late [Pie97b]. latest [Van97c]. Latte [Joh97d]. lauds [UG96]. launch [Rou97b]. launcher [Cio98a]. launches [Ess97f]. lawsuit [Zuk97f]. lawyers [Nic98c]. lays [McK97d]. lean [Ven96e]. Learn [Sun97f]. lease [Fer97a]. legacy [Mor98a]. legal [Blu97d]. Lemay [Van97c]. less [Wei98]. lesson [McM98h]. lessons [Nic98i]. let [Mur96g]. lets [Nic98f]. level [Ham99]. Lexical [McM97f]. Liberté [LH97]. license [Joh97h]. licensees [ME96c]. licenses [Ste96e]. licensing [Mil96c]. life [McC97a]. lifestyle [Ven96d]. lighter [Gut99]. like [Ano97-28]. likes [Syk98]. line [Ste97g]. Linux [Gar98d]. listening [Joh97e]. load [Sun97f]. loader [Ven97h]. loaders [McM96e]. loading [WM97b]. Lobby [Alo97c]. Localize [Sun98g]. Lock manager [Hol98c]. long [Pet97a]. long-term [Pet97a]. longer [Sch96c]. longtime [Ven97a]. look [Val96]. loosens [GS99]. lot [Syk97]. Lotus [RS97]. loves [Bul98]. low [Ham99]. low-level [Ham99]. Lucent [Law96g].

Mac [Mak96]. Mac-based [Mak97]. Mach [Hol99b]. machine [Ven96e]. machines [Rou97b]. Macintosh [O’C96a]. MacOS [Ste98e]. Macworld [Ste98f]. madnes [Smi98]. mail [Dal97]. major [Day98f]. make [Sha96d]. makers [OS98]. makes [Van97d]. making [Syk98]. manager [Ven97f]. March [Mak96]. Marimba [Gut97b]. market [DS97]. Marketplace [Ano96c]. masks [Blu98b]. Master [Car98]. Masterpiece [Fer97a]. match [Kal97a]. Matsushita [Gut99]. matters [Sch96c]. may [Blu97e]. McNealy [Mil96d]. me [Syk96a]. mean [Ven96e]. Measure [Bar99]. Media [Day97d, Day97b]. meet [LaG96c]. meets [LaG96b]. memory [Zuk96]. menus [Mit96g]. merge [MC98a]. Merging [Fri98c]. MessageBox [Har98]. messages [Sun98f]. Messaging [Sun99d]. Metamata [McK98b]. method [Ven97d]. methodology [Hem97a]. methods [McC96c]. Microsoft [Blu96b]. Miko [Mat97a]. millennium [Blu98d]. million [Ste96g]. mind [Arm97c]. mine [Sch97b]. minute [Kes96]. model [Sun97e]. MOM [Sho98d]. momentum [Cha97b]. Money [Bul98]. Monitor [Di 96c]. mouth [Pet97a]. Move [Blu97h]. moves [Blu97e]. Moving [Hug97e]. Mozilla [Mur98]. MPAD [Hem97a]. MRJ [Ste98e]. MS [Ste98i]. MS/Sun [Ste98i]. much [Kad96c]. multi [Hem97a]. multi-panel [Hem97a]. multicast [SD98]. multimedia [Kal96e]. multithreading [McM97n]. must [Rya96a]. must-have [Rya96a]. Mutex [Hol98c]. mutually [Por97]. MVC [Sun98h]. My [McM98d]. Mysteries [Hol99b].

names [Nel98]. native [Di 97d]. navigator [Hug97g, McC96g]. NC [Blu97b]. NCs [SS97]. near [Ess97b]. need [Rya96b]. Negative [Bow97d]. NEO [Sha96e]. Nescafé [Joh98d]. ’Net [LaG96c, Knu97a, Ess96c]. Netscape [Sha96g]. NetWare [Mur96h]. Network [Blu97f]. network-based [Sun97f]. networking [Wis97]. networks [Nel97b, Ess97d]. neutral [Pet97b]. News [O’C96b]. next [Sch96b]. next-generation [SC97]. nicely [Bow97c]. niche [Law96b]. Night [Pie97b]. no [Sch96c]. non [Zuk97a]. non-desktop [Avi98f]. non-standard [Zuk97a]. NONE [Blu96c]. nonlicensees [Gar98e]. Note [O’C96a]. nothing [McM97p]. Novell [Mur96h]. NT [Day98f]. NT/95 [Day98f]. number [Mor98d].

O [Zha97]. Object [Sha96e, Mur96b]. Object/Java [Mur96b]. ObjectPool [Dav98a]. objects [McC96f].

ObjectSpace [Van97g]. ObjectWorld [Hei96]. Observable [Sun96b]. observances [MM98]. Observer [Sun96b, Lop97]. Observer/ Observable [Lop97]. Observers [Bis97]. October [Ano97c]. off [MB97]. offer [Mur96h]. offers [Mur96i]. Office [Law96h]. official [MO98]. older [Nic98f]. OMG [Gar98a]. On-Demand [Sca97b]. one [O’C97, Ven98i]. one-year [O’C97]. ones [Cio98a]. Online [Kal96f, Sch96a]. only [Rya96b]. open [Mit96d, Blu97f]. OpenCard [DT98]. Opening [McM98e]. opens [Gut98b]. operating [Kal96a]. opetition [Fer99a]. Opinion [Par97]. Opt [Blu96d]. Optimize [Bel97]. Oracle [Kal96d]. ORB [Bar99]. ORBs [Hei96]. order [Nic99]. oriented [de 98]. OS/2 [Kal96c]. our [Hug97f]. outlines [Blu96c]. Overcoming [Mit96a]. override [Rou99]. own [ME96a].

package [MD98]. packaged [Ano97-28]. Packaging [Bow98b]. Packard [Coo98a]. packets [Mit97a]. packs [Van97g]. PalmPilots [Avi98e]. pane [Loy99]. panel [Hem97a]. panelists [Ste97f]. Papows [RS97]. paradigm [You96]. Parameter [Ste98a]. parameterization [Sau98b]. Part [McM97g, Di 96c]. participant [Rex97a]. Participants [Alo98f]. partners [McK97e]. partnership [Ste97c]. paste [Dub98a]. patch [Loy99]. patients [Gle98]. pattern [Ste98a]. PC [McM96a]. PCs [Nic98f]. peer [Di 97a]. peer-to-peer [Di 97a]. 95 [Day98f]. C [Mur96g]. consumer [Sun97e]. decompilation [Mah97]. Java [Mur96b]. Microsoft [Ste97c]. Observable [Lop97]. redo [Mes98a]. server [Sho97a].

Sun [Ste98i]. Swing [Van98d]. performance [Zha97]. performs [Ven97e]. perils [Dye98]. persistence [Di 97a]. personal [Kal96d]. Personalized [McC96c]. PersonalJava [Avi98f]. Perspective [Arm98b]. Perspectives [Hem97c]. pitch [SE98]. Pitfall [Zuk97e]. pitfalls [Zuk97a]. Plain [Gre96b]. plan [Kal96e]. plans [Kal97e]. platform [Mil96b]. platform-neutral [Pet97b]. platforms [Sha96d]. play [WM97a]. Player [Day97d]. please [McK97e]. plug [Jer96]. Plug-in [Alo98e]. plug-ins [Jer96]. poet [Hug98d]. poetry [Hug98d]. point [Ven96b]. Polese [Ano96b]. policy [Ven97f]. Poll [Ano97-28]. ponder [Alo98d]. portability [Rou97a]. portable [Sha96h]. ports [McG97]. position [Ano98a]. post [TF97]. posted [Ano96o]. POSTing [Mit97c]. posts [McM98j]. potential [Zuk97a]. power [Sho98a]. powered [McM98f]. Practical [Day98i]. praise [RS97]. Predictions [Blu98d]. Preliminary [Mit96l]. premiums [Sch96a]. Prentice [Mur96k]. Prepare [Zuk98f]. president [Ste97h]. prickly [Ste97f]. primer [Di 97c]. Princeton [Bow98d]. print [Van97d]. privacy [McC96h]. problem [Mit96h]. problems [Ven97j]. process [Gar98f]. processes [Gar98e]. processing [Day98d]. processors [Kal97d]. processware [Bow98b]. producer [Sun97e]. producer/consumer [Sun97e]. Product [Air96a]. productivity [Law96h]. products [Mur96j]. Profile [Sca98-28]. program [O’C96c]. programmer [Gre96a]. programmers [Rya96a]. programming [You96]. programs [McC96b]. Progress [Day98j]. project [Alo98a, Law96j]. prominent [Day97b]. promise [Cli97]. promises [Mur96d]. properties [Van97f]. proprietary [Mur96i]. pros [Jer96]. Protect [Mah97]. prove [Sun99c]. provider [Con96]. provides [Hem97c]. proxy [Kur97]. proxy-based [Kur97]. public [Mur96f]. publishers [Kal96f]. pull [MO98]. punch [Van97g]. punches [FRLS97]. pure [Kal97b]. purification [Mur97c]. purity [Bal97]. push [Ste97i]. pushes [Ste97i]. put [OA98]. puts [McM97j]. Puzzle [Syk96a].

Q&A [Mit96c]. quality [Kad96a]. quest [Mat97a]. question [Mit96d]. questions [Di 97b]. quickly [Ano98n].

R [Sun97g]. RandomAccessFile [Ham99]. range [Fer98b]. ranks [Gar98f]. rate [AO98]. Read [Ano97l]. Reader [Ano97-28]. Readers [Sun97g]. readies [SB97]. ready [Sca98x]. ready-to-buy [DD98]. real [Lie96a]. real-time [Mur96i]. Real-world [Cli97]. Reality [McM97l]. really [Mit96b]. realtime [Di 96c]. receives [Zuk97e]. record [Alo98h]. recursive [wJDM98]. red [Mur96c]. red-hot [Mur96c]. Redirecting [Mit96i]. Reduce [Rou97b]. reference [Fri98b]. References [Mit96c]. Reflection [McM97i, Van98e]. Reflections [Spr97]. reiterates [Nel97b]. rejects [Bus97c]. related [Mur96j]. relational [Spr97]. release [Ess97a]. release-date [Zuk98a]. released [Mur97f]. releases [Ste96d]. Reloading [Mit96j]. remain [Mit96d]. Remote [Som96a]. Render [Hug98a]. reorganization [McK98f]. reorganizes [McK98d]. Report [Fre96a]. request [Bus97c]. requests [Alo97c]. resize [Mit96h]. resources [MC98b]. responds [Alo97c]. Restricted [SD98]. Restricted-channel [SD98]. rests [Bus97e]. results [Nef97]. return [Ven97d]. Returning [Fri98b]. reusable [McC97a]. reveal [Nef97]. reveals [Zuk98b]. revenues [Kal97a]. reverse [Mah97]. Review [Rya96b]. revisited [Mit97d]. Revolutionary [Kru98]. Rhetoric [Sid97b]. rid [Wil98]. right [Kad97b]. Ring [McM98f, Cur98]. ritual [Mur97c]. rivals [Gar98g]. River [Ste96e]. RMI [McK97e]. RMI-based [Don98]. road [Kad97b]. Roaster [Ste97b]. robustness [Dav98b]. role [Sca98b]. rolls [Bow97e]. routines [McC97a]. rule [Blu96g]. rules [MF98a]. ruling [Ste98i]. run [Cio98b]. Runnability [HG98]. runs [Mur97e]. Runtime [Ste98e, Ste97d]. rush [GS98]. rushes [SS97].

safety [Van97f]. sale [Coo98f]. Same [Fre97]. San [Ste98f, Law96j]. sandbox [MF97]. Save [Mit96k]. Saving [Dub98b]. says [Mil96d]. Scale [Sho97b]. scales [McK98c]. scanned [McG97]. schedule [McK97c]. Schmidt [Ess97e]. school [Pér98]. scope [Blu97c]. Scott [Ste98i]. scraps [Nic98a]. scripting [LS97]. SD [Ven98n]. SDK [Zuk97a]. search [Avi98d]. Second [Ano96p]. secure [Nel97a]. security [MF97, Sun99b]. see [Ano97-28]. seeds [Gut97a]. seeks [MS98]. sees [Ess97e]. sell [Kal97d]. Semaphore [Hol98c]. Semaphores [Hol98d]. September [Ano97h]. serial [MD98]. Serialization [Joh98g]. served [Sha96a]. server [Di 96c, Cha97c]. server-side [Hun97]. Servers [Sun98c]. serves [Mur96a]. service [Con96]. services [wRDG98]. servlet [Phi98, Cha97e]. servlet-based [Phi98]. servlets [Sho98b]. session [Sho98c]. set [Van97g]. set-top [Ste98h]. sets [Lie96b]. Share [Dal97]. sheds [Zuk98e]. ships [Mes96]. Shockwave [LaG96c]. short [Pet97a]. short- [Pet97a]. Shot [Joh97d]. should [Mak97]. shouting [Blu97a]. shove [Blu97j]. show [Ano96g]. Showcase [Apr97]. showcased [Mur96b]. Siblings [Gar98g]. side [Hun97]. Siemens [TJ97]. Siggraph [Day98e]. Siggraph97 [Day97b]. Signed [Sun98j]. SIGS [Day98k]. simple [WwRDGSB99]. Singletons [MF98b]. sit [Cul97]. site [Hug97g]. skin [Zuk98e]. skinny [Cha97a]. slice [Van98a]. slimline [Gar97a]. slowly [Kri98b]. smart [TJ97]. snag [Mur97e]. software [Wes96, Mil96a]. Solaris [McM97n]. solution [Mur96c]. solutions [McC97b]. Solve [Ven97j]. some [SC97]. soon [Mor98b]. sought [McC97b]. Sound [Smi97b]. Soup [Joh97f]. source [Cha97d]. source-code [Cha97d]. sows [Gut97a]. Speak [Ano97-28, Ano98a]. speaks [Alo97d]. spec [Gut98b]. Specification [Joh98g]. specs [Mit96l]. speed [Dav98a]. speeds [Mat98b]. spheres [Hug98c]. split [Gar98f]. spot [Dic98]. spotlight [Fer99a]. Spring [Mil96d]. sprouting [Kri98b]. squares [Wil98]. stage [MC98a]. standard [TS97, Ven99d]. standardization [Alo97b]. standardized [McK97b]. standards [Bus97a]. Start [Mur96i]. Start-up [Mur96i]. started [Hun97]. state [Arm97a]. state-dependent [Arm97a]. status [Avi98e]. Step [Hug97g]. Stepping [Hug97g]. stirs [Bow97d]. stock [Sch97a]. StockWatch [Sch96b]. Store [Rou97b]. stories [Sca97t]. Story [Van97e]. Strange [Sch96a, Nel98]. Strategic [Sho97e]. strategies [Pet97a]. strategy [Gut97b]. streams [Mit96i].

Street [Mur96f]. string [Di 96d]. strives [Avi98d]. strolling [Coo98e]. structure [McM97e]. structures [Van97g]. struggle [Ven98i]. Studio [Pet98, McM97k]. Study [Kad96a]. stung [Mur98]. Style [Sho98g]. Subject [de 98]. Subject-oriented [de 98]. success [Ale97]. successor [San98]. suddenly [Sca98-29]. suit [FRLS97]. summer [Ste98c]. Summit [Tro97b]. sums [Ste97h]. Sun [McC96a]. Sun-HP [Fer99a]. Sun-Microsoft [Tro97a]. Sun’s [San98]. Sunscape [Blu96b]. SunSoft [Mes96]. SunWorld [McM98j]. supercomputer [Van97e]. support [Sca97a]. surprise [Fer97b]. Survey [Ano96n]. Swedish [Ess97f]. Swing [Sho98e, Ano97-28]. switches [McM97o]. Sybase [Kri98c]. Symantec [Nic98k]. synchronization [Ven97e]. Synchronizing [McM96c]. synergize [Mur96d]. system [Hug97a]. systems [Kal96a, Ste96e].

tabs [McK98b]. Take [McC97c]. takes [Ess96b]. talking [Sun98b]. talks [Mil98b]. taps [Bow98d]. targets [Avi98f]. tasting [BR98]. TCI [Rod98]. team [Cha97a]. Technical [SKC97]. Techniques [Ven98e]. technologies [Sun99c, Sca98-29]. technology [Ste97i]. telecom [Law96b]. telephony [Sha97]. television [Mur96d]. tell [OA98]. Ten [Kes96]. Ten-minute [Kes96]. term [Pet97a]. Test [Kad97e]. testifies [Nic98e]. testing [MB97]. tests [Man98]. text [Sun97d]. textured [Hug98c]. their [SO97]. them [Rou97b]. thin [Blu97k]. thing [Fre97]. think [Sun98a]. those [Ste97g]. thread [Ven97e]. threaded [Sho97c]. Threading [Hol98e]. threads [Dra96, Van97d]. three [McC96a]. three-day [McC96a]. throws [Syk97]. thumbs [SM98]. tiers [Sho97b]. ties [Ros97]. tight [GS99]. tight-fisted [GS99]. time [Mur96i]. Timers [Hol99a]. timing [O’C98]. Tin [Di 96d]. Tip [Por97]. Tips [Mit96e]. title [Van97c]. toaster [McM97j]. Today [Ano96g, Ano96h]. together [Law96d]. tons [Spe98]. too [Sun98a]. tool [Ess97f]. toolkit [LLZ98, Cha97d]. Tools [Bow96]. top [Hei96]. topic [Mil96a]. topics [SKC97]. tour [Gre97, Alo97a]. touts [Ess96a]. transactions [Kri98c]. transfer [Sun98i]. transition [O’C98]. transport [Mit97a]. traps [Rou99]. TreeFrog [Coo98c]. Trends [Ste96f]. trial [Nic98b]. trick [Wis97]. tried [MO98]. tries [McK97e]. troubleshooting [Joh98a]. trust [Sun99b]. Try [Ven97l]. Try-finally [Ven97l]. turn [Kad97e]. tutorial [Ven98k]. TV [Ess97b]. Twelve [MF98a]. two [Sho97b, Blu97f]. two-stage [MC98a]. type [McM98b]. types [Rou97a]. typing [Mit96k].

U.S. [Bus97d]. ubiquitous [Avi98e]. ugly [HM97a]. Ultranet [Mat97b]. undergoes [Mur97c]. Understanding [MF97]. undo [Mes98a]. undo/redo [Mes98a]. Unfurling [Sun98i]. unite [Alo97a]. universe [Blu96e]. Unix [Kal97e]. unruffled [Bal97]. unveiled [Mit96l]. unveils [ME96c]. upcoming [Joh97e]. update [Rad97a]. updated [Ano96d]. Upgrade [Cio98b]. URL [Zuk98d]. use [Mur96g]. Useful [Day98k]. user [Sun96c, Ess97d]. User-defined [McC96c]. users [FRLS97, Mur96f]. uses [Sun97e]. Using [McC96f].

value [Sch97a]. values [Joh98c]. variables [McC97c]. VC [Ste96g]. Vendors [GS98]. verifier [Ven97i]. version [Ste98e]. versions [Joh98e]. versus [Ven98l]. via [Sha96f]. viable [Ste97a]. view [Rex97a]. views [Som96b, DS97]. violate [Bus97b]. virtual [Ven96e]. virtues [Alo98d]. virus [Nel98]. vision [Blu96c]. visual [McM96f, Mur97e]. VisualAge [SB97]. VJ [Alo98h]. VM [Nic98j]. volleys [Nic98h]. votes [Alo97b]. VRML [Pie97b, Jac96]. vs [Jer96].

Waging [Sun97h]. waiting [Cor97]. wake [FRLS97]. walking [Joh97b]. Wall [Mur96f]. walls [McK98e]. want [Ste97g]. war [Sun97h]. Warp [Kal96c]. way [Mit96f]. Web [TF97, Di 96c]. WebChart [Gle98]. WebObjects [Ess96a]. Webphones [Joh97h]. WebWatcher [Syk97]. Welcome [Blu97i]. we’ve [Cor97]. where [Ste97h]. Which [DT97, Ven98l]. whips [Fer98a]. whiteboard [Hug97f]. who [Ste97g]. wide [Hol98a]. Wider [Jac96]. widespread [McK98f]. wigs [Ess97d]. Will [Lea96]. Willis [TS97]. Win32 [Fri98c]. WinCE [Kal97e]. Wind [Ste96e]. Windows [Law96c]. winners [Ste97e]. wins [Ste96f]. wish [Dug98]. Wonderland [Blu98c]. won’t [Mil98a]. woo [BG97]. words [Tro98a]. work [Sha96d]. WorkShop [Mur96j]. World [Jac96, Mur97c]. world-class [Sun97i]. would [Ano97-28]. wraps [MT98a]. WReader [Ano98w]. Write [Dur97]. Writers [Sun97g]. WWDC [Bus97e].

XML [Mat98b]. year [O’C97]. York [Ste98c]. you’ll [Rya96b]. yourself [Zuk98f].

Zander [Nel97b]. Zealous [Nic98l]. zip [MC98b]. Zone [LaG96b].

References

Airborne:1996:NNPa

[Air96a] Max Airborne. News and new product briefs. Java-World: IDG’s magazine for the Java community, 1 (3):??, May 1996. CO-DEN ???? ISSN 1091-

                                                      8906. URL http://www.
                                                      javaworld.com/javaworld/
                                                      jw-05-1996/jw-05-newsbriefs.
                                                      htm.

Airborne:1996:NNPb

[Air96b] Max Airborne. News and new product briefs. Java-World: IDG’s magazine for the Java community, 1 (4):??, June 1996. CO-DEN ???? ISSN 1091-

                                                      8906. URL http://www.
                                                      javaworld.com/javaworld/
                                                      jw-06-1996/jw-06-newsbriefs.
                                                      htm.

Alexander:1997:PES

[Ale97] Steve Alexander. Perspectives: Ensure the success — and completion — of your real-world Java apps. Java-World: IDG’s magazine for the Java community, 2(11): ??, November 1997. CO-DEN ???? ISSN 1091-

                                                      8906. URL http://www.
                                                      javaworld.com/javaworld/

jw-11-1997/jw-11-perspectives.
htm.

Aloi:1997:NJE

[Alo97a] Jenni Aloi. News from Java Education World Tour: Java-Heads of the world unite. JavaWorld: IDG’s magazine for the Java community, 2 (3):??, March 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-03-1997/jw-03-worldtour.
htm.

Aloi:1997:NJS

[Alo97b] Jenni Aloi. News: Java standardization votes in — but that’s just the beginning. JavaWorld: IDG’s magazine for the Java community, 2 (8):??, August 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-08-1997/jw-08-standards.
htm.

Aloi:1997:NMJ

[Alo97c] Jenni Aloi. News: Microsoft’s ‘Java Evangelist’ responds to Java Lobby requests. JavaWorld: IDG’s magazine for the Java community, 2(10):??, October 1997. CODEN ???? ISSN 1091-8906. URL http://

www.javaworld.com/javaworld/
jw-10-1997/jw-10-javalobby.
htm.

Aloi:1997:SRJ

[Alo97d] Jenni Aloi. Survey report: The Java community speaks out. JavaWorld: IDG’s magazine for the Java community, 2(7):??, July 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-07-1997/jw-07-surveyresults.
                                        htm.

Aloi:1998:DOS

[Alo98a] Jenni Aloi. Developers offer support for JavaReady PC project. JavaWorld: IDG’s magazine for the Java community, 3(9):??, September 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-09-1998/jw-09-pollresults2.
                                        htm.

Aloi:1998:PRD

[Alo98b] Jenni Aloi. Poll results: Do readers support Sun’s latest action against Microsoft? JavaWorld: IDG’s magazine for the Java community, 3(6):??, June 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-06-1998/jw-06-pollresults2.
                                        htm.

Aloi:1998:PRJ

[Alo98c] Jenni Aloi. Poll results: Java API poll results: A divinely inspired Java API? JavaWorld: IDG’s magazine for the Java community, 3(7):??, July 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1998/jw-07-pollresults.
htm.

Aloi:1998:PRR

[Alo98d] Jenni Aloi. Poll results:

Readers ponder the virtues (and defects) of anonymous inner classes. Java-World: IDG’s magazine for the Java community, 3 (7):??, July 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1998/jw-07-pollresults2.
htm.

Aloi:1998:PRW

[Alo98e] Jenni Aloi. Poll results:

Who’s using Sun’s Java plugin — and why or why not? JavaWorld: IDG’s magazine for the Java community, 3(6):??, June 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-06-1998/jw-06-pollresults.
htm.

Aloi:1998:RJP

[Alo98f] Jenni Aloi. Report from JavaOne: Participants speak their mind. JavaWorld: IDG’s magazine for the Java community, 3(4):??, April 1998. CODEN ???? ISSN 1091-8906. URL http://

www.javaworld.com/javaworld/

                                        jw-04-1998/jw-04-javatalk.
                                        htm.

Aloi:1998:SHH

[Alo98g] Jenni Aloi. Swing’s househunting days may be over. JavaWorld: IDG’s magazine for the Java community, 3 (9):??, September 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-09-1998/jw-09-swingchange.
                                        htm.

Aloi:1998:VRP

[Alo98h] Jenni Aloi. VJ++ reader poll draws choice reader comments, record number of votes. JavaWorld: IDG’s magazine for the Java community, 3(5):??, May 1998. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-05-1998/jw-05-pollresults.
                                        htm.

Armstrong:1998:SGI

[AN98] Eric Armstrong and James Niccolai. Sun granted injunction against Microsoft. JavaWorld: IDG’s magazine for the Java community, 3(12):??, 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-12-1998/jw-12-injunction.
                                        htm.

Anonymous:1996:A

[Ano96a] Anonymous. Advertising. JavaWorld: IDG’s magazine for the Java community, 1 (1):??, March 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
common/jw-adsales.htm.

Anonymous:1996:IKP

[Ano96b]  Anonymous.  Interview

with Kim Polese. Java-World: IDG’s magazine for the Java community, 1 (2):??, April 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-04-1996/jw-04-polese.
htm.

Anonymous:1996:JM

[Ano96c] Anonymous. Java marketplace. JavaWorld: IDG’s magazine for the Java community, 1(3):??, May 1996. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
common/jw.market.htm.

Anonymous:1996:JMLa

[Ano96d]  Anonymous.  Java mar-

ketplace (last updated 15 August 1996). Java-World: IDG’s magazine for the Java community, 1(6): ??, August 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
common/jw.market.htm.

Anonymous:1996:JMLb

[Ano96e]  Anonymous.  Java mar-

ketplace (last updated 15 July 1997). JavaWorld: IDG’s magazine for the Java community, 1(7):??, September 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
common/jw.market.htm.

Anonymous:1996:JMLc

[Ano96f]  Anonymous.  Java mar-

ketplace (last updated 2 January 1997). Java-World: IDG’s magazine for the Java community, 1(9): ??, November 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
common/jw.market.htm.

Anonymous:1996:JTSa

[Ano96g] Anonymous. JavaOne today show daily. Java-World: IDG’s magazine for the Java community, 1 (1):??, March 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javaone98/j1-98-toc.htm.

Anonymous:1996:JTSc

[Ano96h] Anonymous. JavaOne Today show daily. Java-World: IDG’s magazine for the Java community, 1 (3):??, May 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javaone98/j1-98-toc.htm.

Anonymous:1996:JTSd

[Ano96i] Anonymous. JavaOne Today show daily. Java-World: IDG’s magazine for the Java community, 1 (4):??, June 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javaone98/j1-98-toc.htm.

Anonymous:1996:JTSe

[Ano96j] Anonymous. JavaOne Today show daily. Java-World: IDG’s magazine for the Java community, 1 (5):??, July 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javaone98/j1-98-toc.htm.

Anonymous:1996:JTSf

[Ano96k] Anonymous. JavaOne Today show daily. Java-World: IDG’s magazine for the Java community, 1(6): ??, August 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javaone98/j1-98-toc.htm.

Anonymous:1996:JTSg

[Ano96l] Anonymous. JavaOne Today show daily. Java-World: IDG’s magazine for the Java community, 1(9): ??, November 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javaone98/j1-98-toc.htm.

Anonymous:1996:JTSh

[Ano96m] Anonymous. JavaOne Today show daily. Java-World: IDG’s magazine for the Java community, 1(9): ??, November 1996. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        javaone98/j1-98-toc.htm.

Anonymous:1996:JCS

[Ano96n] Anonymous. JavaWorld career survey. JavaWorld: IDG’s magazine for the Java community, 1(2):??, April 1996. CODEN ???? ISSN 1091-8906. URL http://

                                        www.javaworld.com/javaworld/
                                        careersurvey-toc.html.

Anonymous:1996:NNP

[Ano96o]  Anonymous.  News and

new product briefs (posted 7/3/96). JavaWorld: IDG’s magazine for the Java community, 1(5):??, July 1996. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-07-1996/jw-07-newsbriefs3.
                                        htm.

Anonymous:1996:SAJ

[Ano96p] Anonymous. Second Annual JavaWorld Editors’ Choice Awards. JavaWorld: IDG’s magazine for the Java community, 1(2):??, April 1996. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/

javaworld/pr/jw-awards.
html.

Anonymous:1997:INL

[Ano97a] Anonymous. IDG.net (last updated 2 December 1997). JavaWorld: IDG’s magazine for the Java community, 2 (12):??, December 1997. CO-DEN ???? ISSN 1091-8906.

URL http://www.idg.net.

Anonymous:1997:JMLf

[Ano97b] Anonymous. Java marketplace (last updated 1 July 1997). JavaWorld: IDG’s magazine for the Java community, 2(7):??, July 1997. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
common/jw.market.htm.

Anonymous:1997:JMLi

[Ano97c]  Anonymous.  Java mar-

ketplace (last updated 1 October 1997). Java-World: IDG’s magazine for the Java community, 2(10): ??, October 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
common/jw.market.htm.

Anonymous:1997:JMLd

[Ano97d] Anonymous. Java marketplace (last updated 15 April 1997). JavaWorld: IDG’s magazine for the Java community, 2(5):??, May 1997. CODEN ???? ISSN 1091-

8906. URL http://www.

                                        javaworld.com/javaworld/
                                        common/jw.market.htm.

Anonymous:1997:JMLj

[Ano97e]  Anonymous.  Java mar-

ketplace (last updated 15 October 1997). Java-World: IDG’s magazine for the Java community, 2(11): ??, November 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        common/jw.market.htm.

Anonymous:1997:JMLc

[Ano97f] Anonymous. Java marketplace (last updated 2 April 1997). JavaWorld: IDG’s magazine for the Java community, 2(4):??, April 1997. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        common/jw.market.htm.

Anonymous:1997:JMLb

[Ano97g] Anonymous. Java marketplace (last updated 3 March 1997). JavaWorld: IDG’s magazine for the Java community, 2(3):??, March 1997. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        common/jw.market.htm.

Anonymous:1997:JMLh

[Ano97h]  Anonymous.  Java mar-

ketplace (last updated 3 September 1997). Java-World: IDG’s magazine for the Java community, 2(9):

??, September 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
common/jw.market.htm.

Anonymous:1997:JMLg

[Ano97i]  Anonymous.  Java mar-

ketplace (last updated 4 August 1997). Java-World: IDG’s magazine for the Java community, 2(8): ??, August 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
common/jw.market.htm.

Anonymous:1997:JMLa

[Ano97j]  Anonymous.  Java mar-

ketplace (last updated 5 February 1997). Java-World: IDG’s magazine for the Java community, 2(2): ??, February 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
common/jw.market.htm.

Anonymous:1997:JMLe

[Ano97k] Anonymous. Java marketplace (last updated 6 June 1997). JavaWorld: IDG’s magazine for the Java community, 2(6):??, June 1997. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
common/jw.market.htm.

Anonymous:1997:JRAa

[Ano97l] Anonymous. JavaOne — read all about it in. Java-World: IDG’s magazine for the Java community, 2 (4):??, April 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        javaone97/jw.index.htm.

Anonymous:1997:JRAb

[Ano97m] Anonymous. JavaOne — read all about it in. Java-World: IDG’s magazine for the Java community, 2 (5):??, May 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        javaone97/jw.index.htm.

Anonymous:1997:JTa

[Ano97n] Anonymous. JavaOne Today. JavaWorld: IDG’s magazine for the Java community, 2(4):??, April 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        javaone97/jw.index.htm.

Anonymous:1997:JTb

[Ano97o] Anonymous. JavaOne Today. JavaWorld: IDG’s magazine for the Java community, 2(5):??, May 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        javaone97/jw.index.htm.

Anonymous:1997:JTSa

[Ano97p] Anonymous. JavaOne Today show daily. Java-World: IDG’s magazine for the Java community, 2(1): ??, January 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javaone98/j1-98-toc.htm.

Anonymous:1997:JTSb

[Ano97q] Anonymous. JavaOne Today show daily. Java-World: IDG’s magazine for the Java community, 2(2): ??, February 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javaone98/j1-98-toc.htm.

Anonymous:1997:JTSc

[Ano97r] Anonymous. JavaOne Today show daily. Java-World: IDG’s magazine for the Java community, 2 (3):??, March 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javaone98/j1-98-toc.htm.

Anonymous:1997:JTSd

[Ano97s] Anonymous. JavaOne Today show daily. Java-World: IDG’s magazine for the Java community, 2 (4):??, April 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javaone98/j1-98-toc.htm.

Anonymous:1997:JTSe

[Ano97t] Anonymous. JavaOne Today show daily. Java-World: IDG’s magazine for the Java community, 2 (5):??, May 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        javaone98/j1-98-toc.htm.

Anonymous:1997:JTSf

[Ano97u] Anonymous. JavaOne Today show daily. Java-World: IDG’s magazine for the Java community, 2 (6):??, June 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        javaone98/j1-98-toc.htm.

Anonymous:1997:JTSg

[Ano97v] Anonymous. JavaOne Today show daily. Java-World: IDG’s magazine for the Java community, 2 (7):??, July 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        javaone98/j1-98-toc.htm.

Anonymous:1997:JTSh

[Ano97w] Anonymous. JavaOne Today show daily. Java-World: IDG’s magazine for the Java community, 2(8): ??, August 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        javaone98/j1-98-toc.htm.

Anonymous:1997:JTSi

[Ano97x] Anonymous. JavaOne Today show daily. Java-World: IDG’s magazine for the Java community, 2(9): ??, September 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javaone98/j1-98-toc.htm.

Anonymous:1997:JTSj

[Ano97y] Anonymous. JavaOne Today show daily. Java-World: IDG’s magazine for the Java community, 2(10): ??, October 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javaone98/j1-98-toc.htm.

Anonymous:1997:JTSk

[Ano97z] Anonymous. JavaOne Today show daily. Java-World: IDG’s magazine for the Java community, 2(11): ??, November 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javaone98/j1-98-toc.htm.

Anonymous:1997:JTSl

[Ano97-27] Anonymous. JavaOne Today show daily. Java-World: IDG’s magazine for the Java community, 2(12): ??, December 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javaone98/j1-98-toc.htm.

Anonymous:1997:RPWa

[Ano97-28] Anonymous. Reader poll: Where would you like to see Swing packaged for use with the 1.2 JDK? speak out! JavaWorld: IDG’s magazine for the Java community, 2(10):??, October 1997. CODEN ???? ISSN 1091-8906. URL

http://nigeria.wpi.com/
cgi-bin/gwpoll/gwpoll/
ballot.htm.

Anonymous:1997:RPWb

[Ano97-29] Anonymous. Reader poll: Where would you like to see Swing packaged for use with the 1.2 JDK? speak out! JavaWorld: IDG’s magazine for the Java community, 2(11):??, November 1997. CODEN ???? ISSN 1091-8906. URL

http://nigeria.wpi.com/
cgi-bin/gwpoll/gwpoll/
ballot.htm.

Anonymous:1997:RPWc

[Ano97-30] Anonymous. Reader poll: Where would you like to see Swing packaged for use with the 1.2 JDK? speak out! JavaWorld: IDG’s magazine for the Java community, 2(12):??, December 1997. CODEN ???? ISSN 1091-8906. URL

http://nigeria.wpi.com/
cgi-bin/gwpoll/gwpoll/
ballot.htm.

Anonymous:1998:JCC

[Ano98a] Anonymous. Java coding conventions — what’s your position? Speak out! Java-World: IDG’s magazine for the Java community, 3(9):??, September 1998. CODEN ???? ISSN 1091-8906. URL

http://nigeria.wpi.com/
cgi-bin/gwpoll/gwpoll/
ballot.htm.

Anonymous:1998:JTSa

[Ano98b] Anonymous. JavaOne Today show daily. Java-World: IDG’s magazine for the Java community, 3(1): ??, January 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javaone98/j1-98-toc.htm.

Anonymous:1998:JTSb

[Ano98c] Anonymous. JavaOne Today show daily. Java-World: IDG’s magazine for the Java community, 3(2): ??, February 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javaone98/j1-98-toc.htm.

Anonymous:1998:JTSc

[Ano98d] Anonymous. JavaOne Today show daily. Java-World: IDG’s magazine for the Java community, 3 (3):??, March 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.

                                        javaworld.com/javaworld/
                                        javaone98/j1-98-toc.htm.

Anonymous:1998:JTSd

[Ano98e] Anonymous. JavaOne Today show daily. Java-World: IDG’s magazine for the Java community, 3 (4):??, April 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        javaone98/j1-98-toc.htm.

Anonymous:1998:JTSe

[Ano98f] Anonymous. JavaOne Today show daily. Java-World: IDG’s magazine for the Java community, 3 (5):??, May 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        javaone98/j1-98-toc.htm.

Anonymous:1998:JTSf

[Ano98g] Anonymous. JavaOne Today show daily. Java-World: IDG’s magazine for the Java community, 3 (6):??, June 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        javaone98/j1-98-toc.htm.

Anonymous:1998:JTSg

[Ano98h] Anonymous. JavaOne Today show daily. Java-World: IDG’s magazine for the Java community, 3 (7):??, July 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javaone98/j1-98-toc.htm.

Anonymous:1998:JTSh

[Ano98i] Anonymous. JavaOne Today show daily. Java-World: IDG’s magazine for the Java community, 3(8): ??, August 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javaone98/j1-98-toc.htm.

Anonymous:1998:JTSi

[Ano98j] Anonymous. JavaOne Today show daily. Java-World: IDG’s magazine for the Java community, 3(9): ??, September 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javaone98/j1-98-toc.htm.

Anonymous:1998:JTSj

[Ano98k] Anonymous. JavaOne Today show daily. JavaWorld: IDG’s magazine for the Java community, 3(10):??, 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javaone98/j1-98-toc.htm.

Anonymous:1998:JTSk

[Ano98l] Anonymous. JavaOne Today show daily. JavaWorld: IDG’s magazine for the Java community, 3(11):??, 1998. CODEN ???? ISSN 1091-

8906. URL http://www.

                                        javaworld.com/javaworld/
                                        javaone98/j1-98-toc.htm.

Anonymous:1998:JTSl

[Ano98m] Anonymous. JavaOne Today show daily. JavaWorld: IDG’s magazine for the Java community, 3(12):??, 1998. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        javaone98/j1-98-toc.htm.

Anonymous:1998:LJA

[Ano98n] Anonymous. Launch Java applications quickly from Windows 95. JavaWorld: IDG’s magazine for the Java community, 3(1):??, January 1998. CODEN ???? ISSN 1091-8906. URL http://

                                        www.javaworld.com/javaworld/
                                        javaworld/javatips/jw-
                                        javatip45.html.

Anonymous:1998:RPWa

[Ano98o] Anonymous. Reader poll: Where would you like to see Swing packaged for use with the 1.2 JDK? speak out! JavaWorld: IDG’s magazine for the Java community, 3(1):??, January 1998. CODEN ???? ISSN 1091-8906. URL

                                        http://nigeria.wpi.com/
                                        cgi-bin/gwpoll/gwpoll/
                                        ballot.htm.

Anonymous:1998:RPWb

[Ano98p] Anonymous. Reader poll: Where would you like to see Swing packaged for use with the 1.2 JDK? speak out! JavaWorld: IDG’s magazine for the Java community, 3(2):??, February 1998. CODEN ???? ISSN 1091-8906. URL

http://nigeria.wpi.com/
cgi-bin/gwpoll/gwpoll/
ballot.htm.

Anonymous:1998:RPWc

[Ano98q] Anonymous. Reader poll:

Where would you like to see Swing packaged for use with the 1.2 JDK? speak out! JavaWorld: IDG’s magazine for the Java community, 3 (3):??, March 1998. CODEN ???? ISSN 1091-8906. URL

http://nigeria.wpi.com/
cgi-bin/gwpoll/gwpoll/
ballot.htm.

Anonymous:1998:RPWd

[Ano98r] Anonymous. Reader poll:

Where would you like to see Swing packaged for use with the 1.2 JDK? speak out! JavaWorld: IDG’s magazine for the Java community, 3 (6):??, June 1998. CODEN ???? ISSN 1091-8906. URL

http://nigeria.wpi.com/
cgi-bin/gwpoll/gwpoll/
ballot.htm.

Anonymous:1998:SSJ

[Ano98s] Anonymous. server-side Java performance article. Java-World: IDG’s magazine for the Java community, 3(8): ??, August 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.

                                        javaworld.com/javaworld/
                                        jw-12-1997/jw-12-volanomark.
                                        htm.

Anonymous:1998:WWYd

[Ano98t] Anonymous. Where would you like to see Swing packaged for use with the 1.2 JDK? speak out! Java-World: IDG’s magazine for the Java community, 3(4): ??, April 1998. CODEN ???? ISSN 1091-8906. URL

                                        http://nigeria.wpi.com/
                                        cgi-bin/gwpoll/gwpoll/
                                        ballot.htm.

Anonymous:1998:WWYe

[Ano98u] Anonymous. Where would you like to see Swing packaged for use with the 1.2 JDK? speak out! Java-World: IDG’s magazine for the Java community, 3(5): ??, May 1998. CODEN ???? ISSN 1091-8906. URL

                                        http://nigeria.wpi.com/
                                        cgi-bin/gwpoll/gwpoll/
                                        ballot.htm.

Anonymous:1998:WDY

[Ano98v] Anonymous. Why do you use servlets? Java-World: IDG’s magazine for the Java community, 3 (11):??, 1998. CODEN ???? ISSN 1091-8906. URL

                                        http://nigeria.wpi.com/
                                        cgi-bin/gwpoll/gwpoll/
                                        ballot.htm.

Anonymous:1998:WPH

[Ano98w] Anonymous. WReader poll: here would you like to see Swing packaged for use with the 1.2 JDK? speak out! JavaWorld: IDG’s magazine for the Java community, 3 (7):??, July 1998. CODEN ???? ISSN 1091-8906. URL

http://nigeria.wpi.com/
cgi-bin/gwpoll/gwpoll/
ballot.htm.

Anonymous:1999:JTS

[Ano99] Anonymous. JavaOne Today show daily. Java-World: IDG’s magazine for the Java community, 4(1): ??, January 1999. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javaone98/j1-98-toc.htm.

Aloi:1998:RRJ

[AO98] Jenni Aloi and Michael O’Connell. Readers rate JavaWorld content. Java-World: IDG’s magazine for the Java community, 3(12): ??, 1998. CODEN ???? ISSN 1091-8906. URL http://

www.javaworld.com/javaworld/
jw-12-1998/jw-12-pollresults.
htm.

April:1997:JTS

[Apr97] Carolyn A. April. Java tools, security solutions highlight Internet Showcase. Java-World: IDG’s magazine for the Java community, 2 (5):??, May 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/

                                        jw-05-1997/jw-05-infoworld.
                                        netshowcase.htm.

Armstrong:1997:HIS

[Arm97a] Eric Armstrong. How to implement state-dependent behavior. JavaWorld: IDG’s magazine for the Java community, 2(8):??, August 1997. CODEN ???? ISSN 1091-8906. URL http://

                                        www.javaworld.com/javaworld/
                                        jw-08-1997/jw-08-stated.
                                        htm.

Armstrong:1997:JDC

[Arm97b] Eric Armstrong. Java in depth: Create enumerated constants in Java. Java-World: IDG’s magazine for the Java community, 2 (7):??, July 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-07-1997/jw-07-enumerated.
                                        htm.

Armstrong:1997:JOA

[Arm97c] Eric Armstrong. JavaOne offers advancements, mind benders, and disappointments. JavaWorld: IDG’s magazine for the Java community, 2(5):??, May 1997. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-05-1997/jw-05-javaone-
                                        armstrong.htm.

Armstrong:1998:CSH

[Arm98a]  Eric Armstrong.  Cover

story: HotSpot: A new breed of virtual machine. JavaWorld: IDG’s magazine for the Java community, 3 (3):??, March 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-03-1998/jw-03-hotspot.
htm.

Armstrong:1998:PSM

[Arm98b] Eric Armstrong. Perspective: Sun and Microsoft duke it out in court. JavaWorld: IDG’s magazine for the Java community, 3(10):??, 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1998/jw-10-injunction.
htm.

Airborne:1996:JNMa

[AS96a] Max Airborne and Jill Steinberg. JavaOne news and more! JavaWorld: IDG’s magazine for the Java community, 1(4):??, June 1996. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-06-1996/jw-06-newsbriefs2.
htm.

Airborne:1996:JNMb

[AS96b] Max Airborne and Jill Steinberg. JavaOne news and more! JavaWorld: IDG’s magazine for the Java community, 1(5):??, July 1996. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/

                                        jw-07-1996/jw-07-newsbriefs2.
                                        htm.

Airborne:1996:NNPc

[AS96c] Max Airborne and Jill Steinberg. News and new product briefs. JavaWorld: IDG’s magazine for the Java community, 1(5):??, July 1996. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-07-1996/jw-07-newsbriefs.
                                        htm.

Airborne:1996:NNPd

[AS96d] Max Airborne and Jill Steinberg. News and new product briefs (7/15/96). Java-World: IDG’s magazine for the Java community, 1(6): ??, August 1996. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-08-1996/jw-08-newsbriefs1.
                                        htm.

Aloi:1997:NNN

[AS97] Jenni Aloi and Jill Steinberg. News: News and new product briefs (5/15/97). JavaWorld: IDG’s magazine for the Java community, 2(6):??, June 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-06-1997/jw-06-newsbriefs.
                                        htm.

Aviram:1998:BYN

[Avi98a] Mariva H. Aviram. Boost your network with fully Javaintegrated NetWare 5. Java-World: IDG’s magazine for the Java community, 3 (4):??, April 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-04-1998/jw-04-netware5.
htm.

Aviram:1998:DCP

[Avi98b] Mariva H. Aviram. Designing a cross-platform look and feel with Java. Java-World: IDG’s magazine for the Java community, 3 (4):??, April 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-04-1998/jw-04-crossplatform.
htm.

Aviram:1998:GON

[Avi98c] Mariva H. Aviram. GNU offers a new kind of Java IDE. JavaWorld: IDG’s magazine for the Java community, 3(7):??, July 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1998/jw-07-freebuilder.
htm.

Aviram:1998:NYC

[Avi98d] Mariva H. Aviram. News you can use: Code-centric search tool strives to reduce Java development time. JavaWorld: IDG’s magazine for the Java community, 3(6):??, June 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-06-1998/jw-06-jcentral.
                                        htm.

Aviram:1998:PUJ

[Avi98e] Mariva H. Aviram. PalmPilots ubiquitous at JavaOne, despite device’s current decaffeinated status. Java-World: IDG’s magazine for the Java community, 3 (4):??, April 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-04-1998/jw-04-palmpilot.
                                        htm.

Aviram:1998:PTN

[Avi98f] Mariva H. Aviram. PersonalJava targets non-desktop computers. JavaWorld: IDG’s magazine for the Java community, 3(4):??, April 1998. CODEN ???? ISSN 1091-8906. URL http://

                                        www.javaworld.com/javaworld/
                                        jw-04-1998/jw-04-pjava.
                                        htm.

Balderston:1997:NJD

[Bal97]  Jim Balderston.  News:

Java developers unruffled by purity question. Java-World: IDG’s magazine for the Java community, 2 (7):??, July 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-07-1997/jw-07-infoworld.
                                        purity.htm.

Barros:1999:MDT

[Bar99] José Barros. Measure data transfer speeds via Sun’s ORB in JDK 1.2 beta 4. JavaWorld: IDG’s magazine for the Java community, 4 (1):??, January 1999. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javatips/jw-javatip65.
htm.

Bowen:1997:NMJ

[BC97] Ted Smalley Bowen and Cara Cunningham. News: Microsoft, JavaSoft Java battle escalates. Java-World: IDG’s magazine for the Java community, 2(8): ??, August 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-08-1997/jw-08-infoworld.
javabattle.htm.

Becker:1998:GSJ

[Bec98a] Dan Becker. Get started with the Java Collections Framework. JavaWorld: IDG’s magazine for the Java community, 3(11):??, 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-11-1998/jw-11-collections.
htm.

Becker:1998:IYP

[Bec98b] Dan Becker. Improve your programs with practical audio. JavaWorld: IDG’s magazine for the Java community, 3(6):??, June 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-06-1998/jw-06-sound.
htm.

Bell:1997:MJF

[Bel97]  Doug Bell.  Make Java

fast: Optimize! Java-World: IDG’s magazine for the Java community, 2 (4):??, April 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-04-1997/jw-04-optimize.
htm.

Balderston:1997:NJE

[BG97] Jim Balderston and Dana Gardner. News from Java Education World Tour: IBM joins Java apostles in 40city tour to woo developers. JavaWorld: IDG’s magazine for the Java community, 2(3):??, March 1997. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-03-1997/jw-03-infoworld.
worldtour.htm.

Biggs:1998:NVC

[Big98] Maggie Biggs. News: Visual Café for Java 2.5 delivers development flexibility. JavaWorld: IDG’s magazine for the Java community, 3(3):??, March 1998. CODEN ???? ISSN 1091-

8906. URL http://www.

javaworld.com/javaworld/
jw-03-1998/jw-03-infoworld.
vcfj.htm.

Bishop:1997:JTT

[Bis97] Philip Bishop. Java tip 38: The trick to ‘Iterator Observers’. Java-World: IDG’s magazine for the Java community, 2(10): ??, October 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javatips/jw-javatip38.
htm.

Blundon:1996:BCJ

[Blu96a] William Blundon. Blundon’s corner: Java and ActiveX. JavaWorld: IDG’s magazine for the Java community, 1 (7):??, September 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1996/jw-09-blundon.
htm.

Blundon:1996:BCM

[Blu96b] William Blundon. Blundon’s corner: Microsoft vs. Sunscape. JavaWorld: IDG’s magazine for the Java community, 1(3):??, May 1996. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-05-1996/jw-05-blundon.
htm.

Blundon:1996:BCN

[Blu96c] William Blundon. Blundon’s corner: Netscape outlines its vision for the future: NONE. JavaWorld: IDG’s magazine for the Java community, 1 (9):??, November 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1996/jw-10-blundon.
htm.

Blundon:1996:BCP

[Blu96d] William Blundon. Blundon’s corner: Plug-ins? opt out! JavaWorld: IDG’s magazine for the Java community, 1(4):??, June 1996. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-06-1996/jw-06-blundon.
htm.

Blundon:1996:BCC

[Blu96e] William Blundon. Blundon’s corner: The center of the universe is a database. JavaWorld: IDG’s magazine for the Java community, 1(5):??, July 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1996/jw-07-blundon.
htm.

Blundon:1996:WCK

[Blu96f] William Blundon. Blundon’s corner: What can kill Java? JavaWorld: IDG’s magazine for the Java community, 1 (6):??, August 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/

jw-08-1996/jw-08-blundon.
htm.

Blundon:1996:HCJ

[Blu96g]  William Blundon.  How

can Java rule the world? JavaWorld: IDG’s magazine for the Java community, 1 (2):??, April 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-04-1996/jw-04-worldrule.
htm.

Blundon:1997:BCA

[Blu97a] William Blundon. Blundon’s corner: All over but the shouting? Java-World: IDG’s magazine for the Java community, 2(10): ??, October 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1997/jw-10-blundon.
htm.

Blundon:1997:BCE

[Blu97b] William Blundon. Blundon’s corner: eciphering the NC world. JavaWorld: IDG’s magazine for the Java community, 2(3):??, March 1997. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-03-1997/jw-03-blundon.
htm.

Blundon:1997:HMB

[Blu97c] William Blundon. Blundon’s corner: How Microsoft is broadening Java’s scope.

JavaWorld: IDG’s magazine for the Java community, 2 (2):??, February 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-02-1997/jw-02-blundon.
htm.

Blundon:1997:MFL

[Blu97d] William Blundon. Blundon’s corner: Microsoft faces a legal fury. Java-World: IDG’s magazine for the Java community, 2(12): ??, December 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-12-1997/jw-12-blundon.
htm.

Blundon:1997:BCM

[Blu97e] William Blundon. Blundon’s corner: Microsoft’s Internet Explorer moves may create a Netscape bonanza. JavaWorld: IDG’s magazine for the Java community, 2 (8):??, August 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-08-1997/jw-08-blundon.
htm.

Blundon:1997:BCN

[Blu97f] William Blundon. Blundon’s corner: ‘Netscape TWO’ — improvements to the Open Network Environment. Java-World: IDG’s magazine for the Java community, 2(1): ??, January 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1997/jw-01-blundon.
htm.

Blundon:1997:BCSa

[Blu97g] William Blundon. Blundon’s corner: Security is in the eye of the beholder. Java-World: IDG’s magazine for the Java community, 2(9): ??, September 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1997/jw-09-blundon.
htm.

Blundon:1997:BCSb

[Blu97h] William Blundon. Blundon’s corner: Sun makes ‘the big move’. Java-World: IDG’s magazine for the Java community, 2(11): ??, November 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-11-1997/jw-11-blundon.
htm.

Blundon:1997:BCW

[Blu97i] William Blundon. Blundon’s corner: Welcome to the extranet. JavaWorld: IDG’s magazine for the Java community, 2(6):??, June 1997. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-06-1997/jw-06-blundon.
htm.

Blundon:1997:WPC

[Blu97j] William Blundon. Blundon’s corner: When ‘push’ comes to shove. JavaWorld: IDG’s magazine for the Java community, 2(4):??, April 1997. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-04-1997/jw-04-blundon.
                                        htm.

Blundon:1997:NFT

[Blu97k] William Blundon. News: Farewell to the thin client. JavaWorld: IDG’s magazine for the Java community, 2(7):??, July 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-07-1997/jw-07-blundon.
                                        htm.

Blundon:1998:YRJ

[Blu98a] William Blundon. Are you ready for the Java backlash? JavaWorld: IDG’s magazine for the Java community, 3 (8):??, August 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-08-1998/jw-08-blundon.
                                        htm.

Blundon:1998:BCB

[Blu98b] William Blundon. Blundon’s corner: Behind the masks of Microsoft and Sun. JavaWorld: IDG’s magazine for the Java community, 3 (4):??, April 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-04-1998/jw-04-blundon.
htm.

Blundon:1998:BCJ

[Blu98c] William Blundon. Blundon’s corner: Java in wonderland. JavaWorld: IDG’s magazine for the Java community, 3(6):??, June 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-06-1998/jw-06-blundon.
html.

Blundon:1998:BCP

[Blu98d] William Blundon. Blundon’s corner: Predictions for the millennium. Java-World: IDG’s magazine for the Java community, 3(2): ??, February 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-02-1998/jw-02-blundon.
htm.

Blundon:1998:JCE

[Blu98e]  William Blundon.  Java:

From the courtroom to the enterprise. JavaWorld: IDG’s magazine for the Java community, 3(12):??, 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-12-1998/jw-12-blundon.
htm.

Blundon:1998:WJM

[Blu98f] William Blundon. Why Jini matters. JavaWorld: IDG’s magazine for the Java community, 3(10):??, 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-10-1998/jw-10-blundon.
                                        htm.

Bowen:1996:TJD

[Bow96] Barry D. Bowen. Tools for the Java developer. Java-World: IDG’s magazine for the Java community, 1 (3):??, May 1996. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-05-1996/jw-05-devenvirons.
                                        htm.

Bowen:1997:CSJ

[Bow97a]  Barry D. Bowen.  CSX

says Java delivers. Java-World: IDG’s magazine for the Java community, 2 (5):??, May 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-05-1997/jw-05-jo-csx.
                                        htm.

Bowen:1997:FEE

[Bow97b] Barry D. Bowen. Feature: Educators embrace Java. JavaWorld: IDG’s magazine for the Java community, 2 (1):??, January 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.

javaworld.com/javaworld/
jw-01-1997/jw-01-education.
htm.

Bowen:1997:FEI

[Bow97c]  Barry D. Bowen.  Fea-

ture: The elephant (IBM) is dancing nicely with Java. JavaWorld: IDG’s magazine for the Java community, 2(7):??, July 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1997/jw-07-ibm.htm.

Bowen:1997:NAN

[Bow97d] Barry D. Bowen. News analysis: Negative developer survey stirs controversy. Java-World: IDG’s magazine for the Java community, 2(12): ??, December 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-12-1997/jw-12-marketdecisions.
htm.

Bowen:1997:NIR

[Bow97e] Ted Smalley Bowen. News:

IBM rolls out Java tools, technology. JavaWorld: IDG’s magazine for the Java community, 2(8):??, August 1997. CODEN ???? ISSN 1091-8906. URL http://

www.javaworld.com/javaworld/
jw-08-1997/jw-08-infoworld-
tools.htm.

Bowen:1998:NJA

[Bow98a] Barry D. Bowen. News:

JFC and AFC debate creates confusion, demands choices. JavaWorld: IDG’s magazine for the Java community, 3 (1):??, January 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1998/jw-01-afc.htm.

Bowen:1998:RWJa

[Bow98b] Barry D. Bowen. Realworld Java: Packaging processware with Java. Java-World: IDG’s magazine for the Java community, 3 (6):??, June 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-06-1998/jw-06-realworld-
crossworlds.htm.

Bowen:1998:RWJb

[Bow98c] Barry D. Bowen. Real-world Java: Why ETrade banks on Java to keep bulls and bears in line. JavaWorld: IDG’s magazine for the Java community, 3(6):??, June 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-06-1998/jw-06-realworld-
etrade.htm.

Bowen:1998:PTJ

[Bow98d] Ted Smalley Bowen. Princeton taps Java to extend its financial systems. JavaWorld: IDG’s magazine for the Java community, 3(10):??, 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/

jw-10-1998/jw-10-iw-princeton.
htm.

Bowen:1999:ISS

[Bow99] Ted Smalley Bowen. IBM sows seeds for Java’s growth. JavaWorld: IDG’s magazine for the Java community, 4 (1):??, January 1999. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1999/jw-01-iw-ibm.
htm.

Bowen:1998:JST

[BR98] Ted Smalley Bowen and Lynda Radosevich. Java on the server is tasting good. JavaWorld: IDG’s magazine for the Java community, 3 (8):??, August 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-08-1998/jw-08-iw-entjava.
htm.

Brier:1998:NIA

[Bri98]  Steven E. Brier.  News:

IBM augments OS/2 with new Java virtual machine. JavaWorld: IDG’s magazine for the Java community, 3 (3):??, March 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-03-1998/jw-03-infoworld.
ibm.htm.

Bowen:1998:IBX

[BS98] Ted Smalley Bowen and Ed Scannell. IBM bets on XML. JavaWorld: IDG’s magazine for the Java community, 3(8):??, August 1998. CODEN ???? ISSN 1091-8906. URL http://

www.javaworld.com/javaworld/
jw-08-1998/jw-08-iw-xml.
htm.

Bull:1998:NML

[Bul98]  Katherine Bull.  News:

Money loves Java. Java-World: IDG’s magazine for the Java community, 3 (3):??, March 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-03-1998/jw-03-idgns.
money.htm.

Busse:1997:NJR

[Bus97a] Torsten Busse. News: Java-Soft responds to ISO community on standards issue. JavaWorld: IDG’s magazine for the Java community, 2 (10):??, October 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1997/jw-10-idgns.
standards.htm.

Busse:1997:NMD

[Bus97b] Torsten Busse. News: Microsoft does not violate Java contract, Ballmer says. Java-World: IDG’s magazine for the Java community, 2(11): ??, November 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/

jw-11-1997/jw-11-idgns.
microsoft.htm.

Busse:1997:NSR

[Bus97c] Torsten Busse. News: Sun rejects request to hand over Java. JavaWorld: IDG’s magazine for the Java community, 2(10): ??, October 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1997/jw-10-idgns.
iso.htm.

Busse:1997:NUI

[Bus97d] Torsten Busse. News: U.S.

ISO committee rejects Sun’s Java gatekeeper bid. Java-World: IDG’s magazine for the Java community, 2(11): ??, November 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-11-1997/jw-11-idgns.
iso.htm.

Busse:1997:NWA

[Bus97e]  Torsten Busse.  News:

WWDC: Apple’s future rests on cross-platform development. JavaWorld: IDG’s magazine for the Java community, 2(6):??, June 1997. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-06-1997/jw-06-idgns.
apple.htm.

Carey:1998:MDM

[Car98] Patrick Carey. Master of disguises: Making JavaBeans look like ActiveX controls. JavaWorld: IDG’s magazine for the Java community, 3 (9):??, September 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-09-1998/jw-09-bridge.
                                        htm.

Chen:1998:JDU

[CG98] Zhiqun Chen and Rinaldo Di Giorgio. Java developer: Understanding Java Card 2.0. JavaWorld: IDG’s magazine for the Java community, 3 (3):??, March 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-03-1998/jw-03-javadev.
                                        html.

Chang:1997:IJW

[Cha97a] Phil Inje Chang. Interview: The Java Web Server team gives you the skinny. JavaWorld: IDG’s magazine for the Java community, 2(7):??, July 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-07-1997/jw-07-serverinterview.
                                        htm.

Chang:1997:JGM

[Cha97b] Phil Inje Chang. JDBC gains momentum. Java-World: IDG’s magazine for the Java community, 2 (5):??, May 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-05-1997/jw-05-jo-jdbc.
htm.

Chang:1997:NJW

[Cha97c] Phil Inje Chang. News: Java Web Server ships! Java-World: IDG’s magazine for the Java community, 2 (6):??, June 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-06-1997/jw-06-webserver.
htm.

Chang:1997:NJT

[Cha97d] Phil Inje Chang. News:

JavaServer Toolkit ships in source-code form. Java-World: IDG’s magazine for the Java community, 2(9): ??, September 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1997/jw-09-javaserver.
htm.

Chang:1997:NJD

[Cha97e] Phil Inje Chang. News: Java-Soft delivers Servlet Dev Kit. JavaWorld: IDG’s magazine for the Java community, 2 (4):??, April 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-04-1997/jw-04-servlet.
htm.

Cioroainu:1998:FLL

[Cio98a]  Andrei Cioroainu.  Fast

launcher lets you start Java apps as easily as your native ones! Java-World: IDG’s magazine for the Java community, 3(8): ??, August 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        javatips/jw-javatip58.
                                        htm.

Cioroianu:1998:JTU

[Cio98b] Andrei Cioroianu. Java tip 51: Upgrade your browser to run JDK 1.1 applets. JavaWorld: IDG’s magazine for the Java community, 3 (4):??, April 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        javatips/jw-javatip51.
                                        htm.

Smith:1997:PN

[CJ97] Cate T. Corcoran, Gregory S. Smith and Wanda Jackson. Product news. Java-World: IDG’s magazine for the Java community, 2 (5):??, May 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        javaone97/java1-97-toc.
                                        html#product.

Clingan:1997:SJC

[Cli97] John Clingan. Real-world Java: At Sun’s Java Computing Day, language delivers on promise of businesscritical apps. JavaWorld: IDG’s magazine for the Java community, 2(1):??, January 1997. CODEN ???? ISSN 1091-8906. URL http://

www.javaworld.com/javaworld/
jw-01-1997/jw-01-javaday.
htm.

Coleman:1998:GCJ

[Col98] Arthur Coleman. Giving currency to the Java Card API. JavaWorld: IDG’s magazine for the Java community, 3 (2):??, February 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-02-1998/jw-02-javacard.
htm.

Connors:1996:ECT

[Con96] Jim Connors. The evolution of cable television to interactive communications service provider. Java-World: IDG’s magazine for the Java community, 1(9): ??, November 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1996/jw-10-connors.
htm.

Cook:1998:AHP

[Coo98a]  Rick Cook.  Analysis

of Hewlett–Packard’s JVM: Java grows up. Java-World: IDG’s magazine for the Java community, 3 (4):??, April 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-04-1998/jw-04-hp-jvm.
                                        htm.

Cook:1998:CSJ

[Coo98b] Rick Cook. Cover story: Java embeds itself in the control market. Java-World: IDG’s magazine for the Java community, 3(1): ??, January 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-01-1998/jw-01-embedded.
                                        htm.

Cook:1998:FCG

[Coo98c] Rick Cook. Fast, cheap and good: Java speeds TreeFrog’s development. JavaWorld: IDG’s magazine for the Java community, 3(11):??, 1998. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-11-1998/jw-11-treefrog.
                                        htm.

Cook:1998:FFW

[Coo98d] Rick Cook. Friend or foe? what does Intel really think about Java? JavaWorld: IDG’s magazine for the Java community, 3(10):??, 1998. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-10-1998/jw-10-sw-intel.
                                        htm.

Cook:1998:JSU

[Coo98e] Rick Cook. Is Java strolling down Unix lane? Java-World: IDG’s magazine for the Java community, 3(9): ??, September 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1998/jw-09-forte.
htm.

Cook:1998:JJP

[Coo98f] Rick Cook. Java jazzes point of sale systems with new API. JavaWorld: IDG’s magazine for the Java community, 3(5):??, May 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-05-1998/jw-05-pos.htm.

Corcoran:1997:WWA

[Cor97] Cate T. Corcoran. What we’ve all been waiting for: Java beyond the desktop. JavaWorld: IDG’s magazine for the Java community, 2(5):??, May 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-05-1997/jw-05-jo-desktop.
htm.

Culla:1997:NMS

[Cul97] Jason Culla. News: Microsoft, Sun sit side by side to discuss Java adoption issues. JavaWorld: IDG’s magazine for the Java community, 2 (9):??, September 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1997/jw-09-javapanel.
htm.

Curry:1998:JDI

[Cur98] Stephen M. Curry. Java developer: An introduction to the Java Ring. Java-World: IDG’s magazine for the Java community, 3 (4):??, April 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-04-1998/jw-04-javadev.
htm.

Daley:1997:JTS

[Dal97]  Tom Daley.  Java tip

#36: Share Java objects using e-mail. JavaWorld: IDG’s magazine for the Java community, 2(9):??, September 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javatips/jw-javatip36.
htm.

Davis:1998:BYO

[Dav98a] Thomas E. Davis. Build your own ObjectPool in Java to boost app speed. JavaWorld: IDG’s magazine for the Java community, 3(6):??, June 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/

jw-06-1998/jw-06-object-
pool.htm.

Davis:1998:IRP

[Dav98b] Thomas E. Davis. Improve the robustness and performance of your ObjectPool. JavaWorld: IDG’s magazine for the Java community, 3 (8):??, August 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-08-1998/jw-08-object-
pool.htm.

Day:1997:CIJ

[Day97a]  Bill Day.  Commentary:

The impact of Java standardization. JavaWorld: IDG’s magazine for the Java community, 2(12): ??, December 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-12-1997/jw-12-iso.commentary.
htm.

Day:1997:JMP

[Day97b] Bill Day. News: Java Media prominent at Siggraph97. JavaWorld: IDG’s magazine for the Java community, 2 (9):??, September 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1997/jw-09-siggraph.
htm.

Day:1997:JG

[Day97c] Bill M. Day, Jr. Java grows up. JavaWorld: IDG’s magazine for the Java community, 2(5):??, May 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-05-1997/jw-05-javaone-
day.htm.

Day:1997:JMF

[Day97d] Bill M. Day, Jr. Java media framework player API. JavaWorld: IDG’s magazine for the Java community, 2 (4):??, April 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-04-1997/jw-04-jmf.htm.

Day:1998:GPJ

[Day98a] Bill Day. 3D graphics programming in Java: Part 1, Java 3D. JavaWorld: IDG’s magazine for the Java community, 3(12):??, 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-12-1998/jw-12-media.
htm.

Day:1998:AIA

[Day98b] Bill Day. Antialiasing, images, and alpha compositing in Java 2D. Java-World: IDG’s magazine for the Java community, 3(8): ??, August 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-08-1998/jw-08-media.
htm.

Day:1998:CSI

[Day98c]  Bill Day.  Cover story:

Introduction to Java media programming. Java-World: IDG’s magazine for the Java community, 3 (6):??, June 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-06-1998/jw-06-media.
htm.

Day:1998:IPJ

[Day98d] Bill Day. Image processing with Java 2D. Java-World: IDG’s magazine for the Java community, 3(9): ??, September 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1998/jw-09-media.
htm.

Day:1998:JNS

[Day98e]  Bill Day.  Java news

from Siggraph ’98. Java-World: IDG’s magazine for the Java community, 3(8): ??, August 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-08-1998/jw-08-siggraph.
htm.

Day:1998:JTW

[Day98f]  Bill Day.  Java tip 55:

Work around a major performance bug in JDK 1.2 Beta 3 for Windows NT/ 95. JavaWorld: IDG’s magazine for the Java community, 3(6):??, June 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javatips/jw-javatip55.
htm.

Day:1998:JCD

[Day98g]  Bill Day.  Java’s cre-

ator discusses technology of Sun–Microsoft lawsuit. JavaWorld: IDG’s magazine for the Java community, 3(10):??, 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1998/jw-10-lawsuit.
htm.

Day:1998:MPG

[Day98h] Bill Day. Media programming: Getting started with Java 2D. JavaWorld: IDG’s magazine for the Java community, 3(7):??, July 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1998/jw-07-media.
htm.

Day:1998:NPJ

[Day98i] Bill Day. News: Practical Java at Software Development ’98. JavaWorld: IDG’s magazine for the Java community, 3(3):??, March 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-03-1998/jw-03-sd98.
day.htm.

Day:1998:PMF

[Day98j] Bill Day. Progress on the media front at Siggraph ’98. JavaWorld: IDG’s magazine for the Java community, 3(10):??, 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1998/jw-10-media.
htm.

Day:1998:UGS

[Day98k] Bill Day. Useful gems from SIGS Conference for Java Development. JavaWorld: IDG’s magazine for the Java community, 3(11):??, 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-11-1998/jw-11-devcon.
htm.

Day:1999:GPJ

[Day99]  Bill Day.  3D graph-

ics programming in Java: Part 2, Java 3D. Java-World: IDG’s magazine for the Java community, 4(1): ??, January 1999. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1999/jw-01-media.
htm.

DiGiorgio:1998:IFR

[DD98] Rinaldo Di Giorgio and Dallas Semiconductor. iButtons: The first ready-to-buy 2.0 Java Card API devices. JavaWorld: IDG’s magazine for the Java community, 3(4):??, April 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-04-1998/jw-04-ibuttons.
htm.

deJong:1998:SOP

[de 98] Gerald A. de Jong. Subjectoriented programming through RMI. JavaWorld: IDG’s magazine for the Java community, 3(8):??, August 1998. CODEN ???? ISSN 1091-8906. URL http://

www.javaworld.com/javaworld/
jw-08-1998/jw-08-subjectop.
htm.

DiGiorgio:1996:JDB

[Di 96a] Rinaldo Di Giorgio. Java developer: Bandwidth hunger. JavaWorld: IDG’s magazine for the Java community, 1 (9):??, November 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1996/jw-10-javadev.
htm.

DiGiorgio:1996:JES

[Di 96b] Rinaldo Di Giorgio. Java developer: Java in embedded systems. Java-World: IDG’s magazine for the Java community, 1(7): ??, September 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1996/jw-09-javadev.
htm.

DiGiorgio:1996:JDMa

[Di 96c] Rinaldo Di Giorgio. Java developer: Monitor your Web server in realtime, Part 1. JavaWorld: IDG’s magazine for the Java community, 1 (2):??, April 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-04-1996/jw-04-javadev.
htm.

DiGiorgio:1996:JDT

[Di 96d] Rinaldo Di Giorgio. Java developer: Tin cans and string. JavaWorld: IDG’s magazine for the Java community, 1 (1):??, March 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-03-1996/jw-03-javadev.
htm.

DiGiorgio:1997:DAP

[Di 97a] Rinaldo Di Giorgio. 1java developer: Add persistence and peer-to-peer computing to Java applets hosted by Netscape 3.0.1. Java-World: IDG’s magazine for the Java community, 2 (5):??, May 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-05-1997/jw-05-javadev.
htm.

DiGiorgio:1997:ISC

[Di 97b] Rinaldo Di Giorgio. Interview: Smart card guru answers questions. Java-World: IDG’s magazine for the Java community, 2(12): ??, December 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-12-1997/jw-12-javadev.
interview.htm.

DiGiorgio:1997:JDJ

[Di 97c] Rinaldo Di Giorgio. Java developer: Java smart card primer. JavaWorld: IDG’s magazine for the Java community, 2(12): ??, December 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-12-1997/jw-12-javadev.
htm.

DiGiorgio:1997:JDU

[Di 97d] Rinaldo Di Giorgio. Java developer: Use native methods to expand the Java environment. JavaWorld: IDG’s magazine for the Java community, 2(7):??, July 1997. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1997/jw-07-javadev.
htm.

Dichter:1998:CSJ

[Dic98] Carl Dichter. Cover story: JavaWorld puts top 6 Java visual IDEs on the spot. JavaWorld: IDG’s magazine for the Java community, 3 (4):??, April 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-04-1998/jw-04-visual-
ides.htm.

DiGiorgio:1996:JDMb

[DiG96] Rinaldo DiGiorgio. Java developer: Monitor your Web server in real time, Part 2. JavaWorld: IDG’s magazine for the Java community, 1(5):??, July 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1996/jw-07-javadev.
htm.

DiGiorgio:1998:JDG

[DM98] Rinaldo Di Giorgio and Michael Montgomery. Java developer: Get a jumpstart on the Java Card. Java-World: IDG’s magazine for the Java community, 3(2): ??, February 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-02-1998/jw-02-javadev.
html.

Donthy:1998:JTH

[Don98] Satheesh A. Donthy. Java tip 56: How to eliminate debugging problems for RMIbased applications. Java-World: IDG’s magazine for the Java community, 3 (7):??, July 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/

                                        javatips/jw-javatip56.
                                        htm.

Drake:1996:IJT

[Dra96] Donald G. Drake. Introduction to Java threads. JavaWorld: IDG’s magazine for the Java community, 1 (2):??, April 1996. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-04-1996/jw-04-threads.
                                        htm.

Doan:1997:NJH

[DS97] Amy Doan and Ed Scannell. News: JavaSoft’s HotJava Views enters Windows market. JavaWorld: IDG’s magazine for the Java community, 2(8):??, August 1997. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-08-1997/jw-08-infoworld-
                                        hotjava.htm.

Dichter:1997:CSW

[DT97] Carl Dichter and Chris Tynes. Cover story: Which Java visual development environment is best for you? JavaWorld: IDG’s magazine for the Java community, 2(6):??, June 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-06-1997/jw-06-devtools-
                                        review.htm.

DiGiorgio:1998:JDS

[DT98] Rinaldo Di Giorgio and Peter Trommler. Java developer: Smart cards and the OpenCard Framework. JavaWorld: IDG’s magazine for the Java community, 3 (1):??, January 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1998/jw-01-javadev.
html.

Dube:1998:LHC

[Dub98a] Jean-Pierre Dubé. Learn how to cut, copy, and paste in Java. JavaWorld: IDG’s magazine for the Java community, 3(10):??, 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javatips/jw-javatip61.
htm.

Dube:1998:SBF

[Dub98b]  Jean-Pierre Dubé.  Sav-

ing bitmap files in Java: a tutorial. JavaWorld: IDG’s magazine for the Java community, 3(9):??, September 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javatips/jw-javatip60.
htm.

Dugan:1998:WJG

[Dug98] Sean Dugan. Will Jini grant Sun’s wish? JavaWorld: IDG’s magazine for the Java community, 3(10):??, 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1998/jw-10-iw-jinitest.
htm.

Durante:1997:WCP

[Dur97] Pat L. Durante. Write CGI programs in Java. Java-World: IDG’s magazine for the Java community, 2(1): ??, January 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1997/jw-01-cgiscripts.
htm.

Dyer:1997:CSJ

[Dye97] Dave Dyer. Cover story: Java decompilers compared. JavaWorld: IDG’s magazine for the Java community, 2(7):??, July 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1997/jw-07-decompilers.
htm.

Dyer:1998:CAS

[Dye98] Dave Dyer. Can Assure save Java from the perils of multithreading? JavaWorld: IDG’s magazine for the Java community, 3(10):??, 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1998/jw-10-assure.
htm.

Essick:1996:JTW

[Ess96a] Kristi Essick. Jobs touts WebObjects. JavaWorld: IDG’s magazine for the Java community, 1(3):??, May 1996. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-05-1996/jw-05-webobjects.
htm.

Essick:1996:NAT

[Ess96b] Kristi Essick. News: Adobe takes on the Internet. Java-World: IDG’s magazine for the Java community, 1 (4):??, June 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-06-1996/jw-06-adobe.
htm.

Essick:1996:NIS

[Ess96c] Kristi Essick. News: IBM, Sun collaborate on Net application architecture. Java-World: IDG’s magazine for the Java community, 1(7): ??, September 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1996/jw-09-ibm.htm.

Essick:1997:IRE

[Ess97a] Kristi Essick. IBM to release embedded Java browser for ’Net devices. Java-World: IDG’s magazine for the Java community, 2 (5):??, May 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-05-1997/jw-05-idgns.
ibmbrowser.htm.

Essick:1997:NFJ

[Ess97b] Kristi Essick. News feature: Java — coming to a TV near you? Java-World: IDG’s magazine for the Java community, 2(11): ??, November 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-11-1997/jw-11-idgns.
javatv.htm.

Essick:1997:NIA

[Ess97c] Kristi Essick. News: Infospace announces new Java applications. JavaWorld: IDG’s magazine for the Java community, 2(1):??, January 1997. CODEN ???? ISSN 1091-8906. URL http://

www.javaworld.com/javaworld/
jw-01-1997/jw-01-idgns.
infospace.htm.

Essick:1997:NIB

[Ess97d] Kristi Essick. News: Networks ’97: Industry bigwigs back new Java User Group. JavaWorld: IDG’s magazine for the Java community, 2(7):??, July 1997. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1997/jw-07-idgns.
jug.htm.

Essick:1997:NNS

[Ess97e] Kristi Essick. News: Novell’s Schmidt sees Java as company’s future. Java-World: IDG’s magazine for the Java community, 2(11): ??, November 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-11-1997/jw-11-infoworld.
schmidt.htm.

Essick:1997:NSC

[Ess97f] Kristi Essick. News: Swedish company launches Java applet tool. JavaWorld: IDG’s magazine for the Java community, 2(8):??, August 1997. CODEN ???? ISSN 1091-8906. URL http://

www.javaworld.com/javaworld/
jw-08-1997/jw-08-idgns.
swede.htm.

Ferranti:1997:NJG

[Fer97a] Marc Ferranti. News: Java gives CA’s masterpiece a new lease on life. Java-World: IDG’s magazine for the Java community, 2(9): ??, September 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1997/jw-09-idgns-
ca.htm.

Ferranti:1997:NMC

[Fer97b] Marc Ferranti. News: Microsoft countersuit against Sun comes as no surprise. JavaWorld: IDG’s magazine for the Java community, 2 (11):??, November 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-11-1997/jw-11-idgns.
countersuit.htm.

Ferranti:1997:NUM

[Fer97c] Marc Ferranti. News: Update: Microsoft countersues Sun on Java licensing. Java-World: IDG’s magazine for the Java community, 2(11): ??, November 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-11-1997/jw-11-idgns.
suitupdate.htm.

Ferranti:1998:IWS

[Fer98a]  Marc Ferranti.  Internet

world: Sun whips up Jini support. JavaWorld: IDG’s magazine for the Java community, 3(10):??, 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1998/jw-10-sw-jini.
htm.

Ferranti:1998:NIL

[Fer98b] Marc Ferranti. News: IBM licenses Sun’s Java chips for range of devices. Java-World: IDG’s magazine for the Java community, 3 (4):??, April 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/

jw-04-1998/jw-04-idgns.
javachips.htm.

Ferranti:1999:JBE

[Fer99a] Marc Ferranti. At Java Business Expo the spotlight’s on Sun-HP “co-opetition”. JavaWorld: IDG’s magazine for the Java community, 4 (1):??, January 1999. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1999/jw-01-idgns-
hpsun.htm.

Ferranti:1999:JTF

[Fer99b] Marc Ferranti. Jini to take a formal bow on January 25. JavaWorld: IDG’s magazine for the Java community, 4(1):??, January 1999. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1999/jw-01-idgns-
jini.htm.

Ferranti:1999:UCJ

[Fer99c] Marc Ferranti. Users confront Java challenges. Java-World: IDG’s magazine for the Java community, 4(1): ??, January 1999. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1999/jw-01-idgns-
jbepanel.htm.

Freedman:1996:JCRa

[Fre96a] Matt Freedman. JavaOne conference report. Java-World: IDG’s magazine for the Java community, 1 (4):??, June 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-06-1996/jw-06-javaone-
freedman.htm.

Freedman:1996:JCRb

[Fre96b] Matt Freedman. JavaOne conference report. Java-World: IDG’s magazine for the Java community, 1 (5):??, July 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1996/jw-07-javaone-
freedman.htm.

Freedman:1997:JCR

[Fre97] Matthew Freedman. JavaOne ’97 conference report: Same thing last year? Java-World: IDG’s magazine for the Java community, 2 (5):??, May 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-05-1997/jw-05-javaone-
freedman.htm.

Friesen:1998:IYS

[Fri98a] Geoff Friesen. Internationalize your software, Part 1. JavaWorld: IDG’s magazine for the Java community, 3(12):??, 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/

jw-12-1998/jw-12-internationalize.
htm.

Friesen:1998:RDR

[Fri98b]  Geoff Friesen.  Java tip

54: Returning data in reference arguments via JNI. JavaWorld: IDG’s magazine for the Java community, 3(6):??, June 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javatips/jw-javatip54.
htm.

Friesen:1998:MJW

[Fri98c] Geoff Friesen. Merging Java and Win32: A new way to develop Windows applications. JavaWorld: IDG’s magazine for the Java community, 3(7):??, July 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1998/jw-07-java-
win32.htm.

Friesen:1999:IYS

[Fri99] Geoff Friesen. Internationalize your software, Part 2. JavaWorld: IDG’s magazine for the Java community, 4 (1):??, January 1999. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1999/jw-01-internationalize.
htm.

Ferranti:1997:NIW

[FRLS97] Marc Ferranti, Lynda Radosevich, Martin LaMonica, and Ed Scannell. News: ITxpo: In wake of suit, McNealy punches, users feel heat. Java-World: IDG’s magazine for the Java community, 2(11): ??, November 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-11-1997/jw-11-idgns.
itxpo.htm.

Gardner:1997:NNP

[Gar97a]  Dana Gardner.  News:

Netscape plans include slimline JavaBeans-based client. JavaWorld: IDG’s magazine for the Java community, 2 (10):??, October 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1997/jw-10-infoworld.
beans.htm.

Gardner:1997:NSN

[Gar97b] Dana Gardner. News: Sun, Netscape to deliver all-Java Web browser. Java-World: IDG’s magazine for the Java community, 2(9): ??, September 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1997/jw-09-infoworld.
javagator.htm.

Gardner:1998:ODC

[Gar98a] Dana Gardner. OMG to define CORBA 3.0 with EJB, visual development support.

JavaWorld: IDG’s magazine for the Java community, 3 (9):??, September 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1998/jw-09-iw-omg.
htm.

Gardner:1998:SSJ

[Gar98b] Dana Gardner. Server-side Java gains greater momentum. JavaWorld: IDG’s magazine for the Java community, 3(11):??, 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-11-1998/jw-11-iw-serverside.
htm.

Gardner:1998:SAJ

[Gar98c] Dana Gardner. Sun and AOL join forces on I-commerce for businesses. JavaWorld: IDG’s magazine for the Java community, 3(12):??, 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-12-1998/jw-12-iw-aolnetscape.
htm.

Gardner:1998:SPJ

[Gar98d] Dana Gardner. Sun ports JDK 1.2 to Linux in push for ubiquitous Java. JavaWorld: IDG’s magazine for the Java community, 3(11):??, 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-11-1998/jw-11-iw-jdklinux.
htm.

Gardner:1998:SOJ

[Gar98e]  Dana Gardner.  Sun to

open Java processes to nonlicensees. JavaWorld: IDG’s magazine for the Java community, 3(12):??, 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-12-1998/jw-12-iw-nonlicensees.
                                        htm.

Gardner:1998:VSR

[Gar98f] Dana Gardner. Vendors split ranks with Sun over standards process for embedded Java. JavaWorld: IDG’s magazine for the Java community, 3(11):??, 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-11-1998/jw-11-iw-standards.
                                        htm.

Gardner:1998:XJS

[Gar98g] Dana Gardner. XML and Java: Siblings or rivals? JavaWorld: IDG’s magazine for the Java community, 3(12):??, 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-12-1998/jw-12-iw-xml.
                                        htm.

Glerum:1998:WSK

[Gle98] Katrina Glerum. WebChart serves up Kaiser patients. JavaWorld: IDG’s magazine for the Java community, 3(11):??, 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-11-1998/jw-11-kaiser.
htm.

Greco:1996:JCC

[Gre96a]  Frank Greco.  Java —

cappuccino for the corporate programmer? Java-World: IDG’s magazine for the Java community, 1 (3):??, May 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-05-1996/jw-05-corppro.
htm.

Greco:1996:JPE

[Gre96b] Frank D. Greco. Java (in plain English): An introduction. JavaWorld: IDG’s magazine for the Java community, 1(1):??, March 1996. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-03-1996/jw-03-java.
intro.htm.

Greco:1997:JJJ

[Gre97] Frank D. Greco. Java jolts Jacob Javits: a high-caffeine tour. JavaWorld: IDG’s magazine for the Java community, 2(1):??, January 1997. CODEN ???? ISSN 1091-8906. URL http://

www.javaworld.com/javaworld/
jw-01-1997/jw-01-iworld3.
htm.

Greanier:1998:BBS

[Gre98] Todd M. Greanier. Building a bigger sandbox. Java-World: IDG’s magazine for the Java community, 3(8): ??, August 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-08-1998/jw-08-sandbox.
                                        htm.

Gardner:1998:VRJ

[GS98] Dana Gardner and Ed Scannell. Vendors rush to be at Java’s (server) side. Java-World: IDG’s magazine for the Java community, 3(9): ??, September 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-09-1998/jw-09-iw-serverside.
                                        htm.

Gardner:1999:SLT

[GS99] Dana Gardner and Ed Scannell. Sun loosens tightfisted grip on Java. Java-World: IDG’s magazine for the Java community, 4(1): ??, January 1999. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-01-1999/jw-01-iw-sunsgrip.
                                        htm.

Guth:1997:IJS

[Gut97a] Rob Guth. Interview: Java sows seeds of change at Sun, Gage says. Java-World: IDG’s magazine for the Java community, 2 (7):??, July 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1997/jw-07-idgns.
gage.htm.

Guth:1997:IMV

[Gut97b]  Rob Guth.  Interview:

Marimba’s van Hoff discusses strategy. Java-World: IDG’s magazine for the Java community, 2(1): ??, January 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1997/jw-01-idgns.
vanhoff.htm.

Guth:1997:NSJ

[Gut97c] Rob Guth. News: Sun’s JavaSpaces is foundation for future distributed systems. JavaWorld: IDG’s magazine for the Java community, 2 (9):??, September 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1997/jw-09-idgns.
javaspaces.htm.

Guth:1998:NSP

[Gut98a] Rob Guth. News: Sun pushes Java’s performance, browser independence. Java-World: IDG’s magazine for the Java community, 3 (3):??, March 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.

                                        javaworld.com/javaworld/
                                        jw-03-1998/jw-03-idgns.
                                        japan.htm.

Guth:1998:SOJ

[Gut98b] Rob Guth. Sun opens Jini spec for review. JavaWorld: IDG’s magazine for the Java community, 3(10):??, 1998. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-10-1998/jw-10-idgns-
                                        jinispec.htm.

Guth:1999:SME

[Gut99] Rob Guth. Sun and Matsushita examine lighter Java. JavaWorld: IDG’s magazine for the Java community, 4 (1):??, January 1999. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-01-1999/jw-01-idgns-
                                        lighterjava.htm.

Hagimoto:1997:CJA

[Hag97]  Junzo Hagimoto.  Con-

struction of Java applications through distributed object technology. Java-World: IDG’s magazine for the Java community, 2(12): ??, December 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-12-1997/jw-12-horb.
                                        htm.

Hamner:1999:URB

[Ham99] Derek Hamner. Use a RandomAccessFile to build a low-level database. Java-World: IDG’s magazine for the Java community, 4(1): ??, January 1999. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1999/jw-01-step.
htm.

Haney:1998:NIC

[Han98] Clare Haney. News: Intel, Comcast invest in Intertainer’s Java-based technology. JavaWorld: IDG’s magazine for the Java community, 3(3):??, March 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-03-1998/jw-03-idgns.
intel-comcast.htm.

Harich:1997:JTF

[Har97] Jack Harich. Java tip 39:

Finally: How to use a basic Java 1.1 network and file class loader. Java-World: IDG’s magazine for the Java community, 2(10): ??, October 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javatips/jw-javatip39.
htm.

Harich:1998:JTH

[Har98] Jack Harich. Java tip 48:

How to create a reusable MessageBox class. Java-World: IDG’s magazine for the Java community, 3 (3):??, March 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        javatips/jw-javatip48.
                                        htm.

Heichler:1996:NIO

[Hei96] Elizabeth Heichler. News: Internet, ORBs top new product intros at Object-World. JavaWorld: IDG’s magazine for the Java community, 1(4):??, June 1996. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-06-1996/jw-06-objectworld.
                                        htm.

Heichler:1999:MAP

[Hei99] Elizabeth Heichler. Microsoft appeals preliminary injunction awarded to Sun. Java-World: IDG’s magazine for the Java community, 4(1): ??, January 1999. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-01-1999/jw-01-idgns-
                                        msappeal.htm.

Hemrajani:1997:CSM

[Hem97a]  Anil Hemrajani.  Cover

story: MPAD: A new design and development methodology for multi-panel applets. JavaWorld: IDG’s magazine for the Java community, 2(5):??, May 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/

jw-05-1997/jw-05-appdesign.
htm.

Hemrajani:1997:HDC

[Hem97b] Anil Hemrajani. How to develop commercial-quality Java programs. Java-World: IDG’s magazine for the Java community, 2(1): ??, January 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1997/jw-01-infobook.
htm.

Hemrajani:1997:PJP

[Hem97c] Anil Hemrajani. Perspectives: Java provides common ground across heterogeneous environments. Java-World: IDG’s magazine for the Java community, 2(8): ??, August 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-08-1997/jw-08-perspectives.
htm.

Hayes:1998:STJ

[HG98] Roger Hayes and Manoochehr Ghiassi. ‘runnability’ testing of Java programs. Java-World: IDG’s magazine for the Java community, 3(8): ??, August 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-08-1998/jw-08-runtest.
htm.

Heichler:1997:SMDa

[HM97a] Elizabeth Heichler and Robert McMillan. Sun–Microsoft dispute gets ugly. Java-World: IDG’s magazine for the Java community, 2(1): ??, January 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-10-1997/jw-10-sunsuit.
                                        html.

Heichler:1997:SMDb

[HM97b] Elizabeth Heichler and Robert McMillan. Sun–Microsoft dispute gets ugly. Java-World: IDG’s magazine for the Java community, 2(11): ??, November 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-10-1997/jw-10-sunsuit.
                                        htm.

Holt:1998:NJD

[Hol98a] Stannie Holt. News: Java draws wide market interest. JavaWorld: IDG’s magazine for the Java community, 3(6):??, June 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-06-1998/jw-06-iw-javamarket.
                                        htm.

Holub:1998:PJTb

[Hol98b]  Allen Holub.  Program-

ming Java threads in the real world, Part 2: Common multithreading pitfalls (deadlock, etc.). JavaWorld: IDG’s magazine for the Java community, 3(10):??, 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1998/jw-10-toolbox.
htm;   http://www.holub.
com/goodies/javaworld/
jw_index.html.

Holub:1998:PJTc

[Hol98c] Allen Holub. Programming Java threads in the real world, Part 3: Semaphore, lock manager, and mutex. JavaWorld: IDG’s magazine for the Java community, 3(11):??, 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-11-1998/jw-11-toolbox.
htm;   http://www.holub.
com/goodies/javaworld/
jw_index.html.

Holub:1998:PJTd

[Hol98d]  Allen Holub.  Program-

ming Java threads in the real world, Part 4: Condition variables and counting semaphores. JavaWorld: IDG’s magazine for the Java community, 3(12):??, 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-12-1998/jw-12-toolbox.
htm;   http://www.holub.
com/goodies/javaworld/
jw_index.html.

Holub:1998:PJTa

[Hol98e]  Allen Holub.  Program-

ming Java threads in the real world: Threading architectures. JavaWorld: IDG’s magazine for the Java community, 3(9):??, September 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-09-1998/jw-09-threads.
                                        htm;   http://www.holub.
                                        com/goodies/javaworld/
                                        jw_index.html.

Holub:1999:PJTa

[Hol99a] Allen Holub. Programming Java threads in the real world, Part 5: Timers. Java-World: IDG’s magazine for the Java community, 4(2):??, February 1999. CODEN ???? ISSN 1091-8906. URL http:

                                        //www.holub.com/goodies/
                                        javaworld/jw_index.html.

Holub:1999:PJTb

[Hol99b]  Allen Holub.  Program-

ming Java threads in the real world, Part 6: Mach ’99: Observer and the mysteries of the AWTEvent-Multicaster. JavaWorld: IDG’s magazine for the Java community, 4(3):??, March 1999. CODEN ???? ISSN 1091-8906. URL http:/

                                        /www.holub.com/goodies/
                                        javaworld/jw_index.html.

Hughes:1997:BIC

[Hug97a] Merlin Hughes. Building an Internet chat system. Java-World: IDG’s magazine for the Java community, 2(1): ??, January 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1997/jw-01-chat.
htm.

Hughes:1997:CSJb

[Hug97b] Merlin Hughes. Cover story:

Java step by step: Build a bevy of beans. Java-World: IDG’s magazine for the Java community, 2(8): ??, August 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-08-1997/jw-08-step.
htm.

Hughes:1997:CSJa

[Hug97c]  Merlin Hughes.  Cover

story: JavaBeans and ActiveX go head to head. JavaWorld: IDG’s magazine for the Java community, 2 (3):??, March 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-03-1997/jw-03-avb-tech.
htm.

Hughes:1997:JSSb

[Hug97d] Merlin Hughes. Java step by step: Building a bevy of beans: Create reusable JavaBeans components. Java-World: IDG’s magazine for the Java community, 2(9): ??, September 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1997/jw-09-step.
htm.

Hughes:1997:JSSa

[Hug97e] Merlin Hughes. Java step by step: Moving to JDK 1.1 — using the delegation event model to create custom AWT components. JavaWorld: IDG’s magazine for the Java community, 2(5):??, May 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-05-1997/jw-05-step.
htm.

Hughes:1997:JSSc

[Hug97f] Merlin Hughes. Java step by step: Networking our whiteboard with Java 1.1. Java-World: IDG’s magazine for the Java community, 2(12): ??, December 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-12-1997/jw-12-step.
html.

Hughes:1997:STS

[Hug97g] Merlin Hughes. Java step by step: Stepping through a site navigator applet. Java-World: IDG’s magazine for the Java community, 2(1): ??, January 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1997/jw-01-step.
htm.

Hughes:1998:GJR

[Hug98a] Merlin Hughes. 3D graphic Java: Render fractal landscapes. JavaWorld: IDG’s magazine for the Java community, 3(8):??, August 1998. CODEN ???? ISSN 1091-8906. URL http://

www.javaworld.com/javaworld/
jw-08-1998/jw-08-step.
htm.

Hughes:1998:AES

[Hug98b] Merlin Hughes. Application and enterprise security with the Java 2 platform. JavaWorld: IDG’s magazine for the Java community, 3(12):??, 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jbe/jw-jbe-security.htm.

Hughes:1998:JSSc

[Hug98c] Merlin Hughes. Java step by step: Draw textured spheres. JavaWorld: IDG’s magazine for the Java community, 3(6):??, June 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-06-1998/jw-06-step.
html.

Hughes:1998:JSSb

[Hug98d] Merlin Hughes. Java step by step: Release your inner poet: Use servlets to develop a collaborative poetry app. JavaWorld: IDG’s magazine for the Java community, 3(4):??, April 1998. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-04-1998/jw-04-step.
                                        html.

Hughes:1999:AES

[Hug99] Merlin Hughes. Application and Enterprise security with the Java 2 platform. JavaWorld: IDG’s magazine for the Java community, 4 (1):??, January 1999. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jbe/jw-jbe-security.htm.

Hunter:1997:HGS

[Hun97] Jason Hunter. How to get started with server-side Java. JavaWorld: IDG’s magazine for the Java community, 2 (3):??, March 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-03-1997/jw-03-ssj.htm.

Hunter:1998:DIK

[Hun98a] Jason Hunter. Developers identify 6+ key benefits of servlets. JavaWorld: IDG’s magazine for the Java community, 3(12):??, 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-12-1998/jw-12-pollresults2.
htm.

Hunter:1998:INS

[Hun98b] Jason Hunter. Introducing the new Servlet API 2.1. JavaWorld: IDG’s magazine for the Java community, 3(12):??, 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-12-1998/jw-12-servletapi.
htm.

Jacob:1996:WW

[Jac96] Karl Jacob. Wider world:

Why Java and VRML? Java-World: IDG’s magazine for the Java community, 1 (1):??, March 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-03-1996/jw-03-javavrml.
htm.

Jernigan:1996:JVP

[Jer96] Rayme Jernigan. Java vs.

plug-ins, the pros and cons. JavaWorld: IDG’s magazine for the Java community, 1(3):??, May 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-05-1996/jw-05-fools.
htm.

Johnson:1997:BSJ

[Joh97a] Mark Johnson. The Bean-Box: Sun’s JavaBeans test container. JavaWorld: IDG’s magazine for the Java community, 2(9):??, September 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-09-1997/jw-09-beanbox.
                                        htm.

Johnson:1997:JWT

[Joh97b] Mark Johnson. JavaBeans: A walking tour of JavaBeans. JavaWorld: IDG’s magazine for the Java community, 2 (8):??, August 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-08-1997/jw-08-beans.
                                        htm.

Johnson:1997:JCB

[Joh97c] Mark Johnson. JavaBeans: Control bean customization. JavaWorld: IDG’s magazine for the Java community, 2 (11):??, November 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-11-1997/jw-11-beans.
                                        htm.

Johnson:1997:JDS

[Joh97d] Mark Johnson. JavaBeans: ‘double shot, half decaf, skinny latte’ — customize your Java. JavaWorld: IDG’s magazine for the Java community, 2(9):??, September 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1997/jw-09-beans.
htm.

Johnson:1997:JKL

[Joh97e] Mark Johnson. JavaBeans:

‘keep listening for upcoming events’. JavaWorld: IDG’s magazine for the Java community, 2(10): ??, October 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1997/jw-10-beans.
htm.

Johnson:1997:JSY

[Joh97f] Mark Johnson. JavaBeans:

Soup up your Java classes with customization. Java-World: IDG’s magazine for the Java community, 2(12): ??, December 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-12-1997/jw-12-beans.
htm.

Johnston:1997:NIW

[Joh97g] Margret Johnston. News: Internet world: Java, Internet are Apple’s future, exec says. JavaWorld: IDG’s magazine for the Java community, 2(6):??, June 1997. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/

                                        jw-06-1997/jw-06-idgns.
                                        applefuture.htm.

Johnston:1997:NSL

[Joh97h] Margret Johnston. News: Sun to license Java for use in webphones. Java-World: IDG’s magazine for the Java community, 2(9): ??, September 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-09-1997/jw-09-idgns.
                                        webphones.htm.

Johnson:1998:BJT

[Joh98a] Mark Johnson. BeanLint: A JavaBeans troubleshooting tool, Part 1. JavaWorld: IDG’s magazine for the Java community, 3(12):??, 1998. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-12-1998/jw-12-beans.
                                        htm.

Johnson:1998:BGE

[Joh98b] Mark Johnson. A beginner’s guide to Enterprise JavaBeans. JavaWorld: IDG’s magazine for the Java community, 3(10):??, 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-10-1998/jw-10-beans.
                                        htm.

Johnson:1998:FV

[Joh98c] Mark Johnson. A fistful of values. JavaWorld: IDG’s magazine for the Java community, 3(11):??, 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-11-1998/jw-11-beans.
htm.

Johnson:1998:JDI

[Joh98d] Mark Johnson. JavaBeans:

Do it the ‘Nescafé’ way — with freeze-dried JavaBeans. JavaWorld: IDG’s magazine for the Java community, 3 (1):??, January 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1998/jw-01-beans.
htm.

Johnson:1998:JCO

[Joh98e] Mark Johnson. JavaBeans:

It’s in the contract! object versions for JavaBeans. JavaWorld: IDG’s magazine for the Java community, 3 (3):??, March 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-03-1998/jw-03-beans.
htm.

Johnson:1998:JJB

[Joh98f] Mark Johnson. JavaBeans:

JavaBeans book review — a detailed look at 3 winners. JavaWorld: IDG’s magazine for the Java community, 3(5):??, May 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/

                                        jw-05-1998/jw-05-beans.
                                        htm.

Johnson:1998:JSJ

[Joh98g] Mark Johnson. JavaBeans: Serialization and the JavaBeans specification. Java-World: IDG’s magazine for the Java community, 3(2): ??, February 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-02-1998/jw-02-beans.
                                        htm.

Johnson:1998:JSG

[Joh98h] Mark Johnson. JavaBeans: Serialization grab bag. Java-World: IDG’s magazine for the Java community, 3 (4):??, April 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-04-1998/jw-04-beans.
                                        htm.

Johnson:1998:JTJ

[Joh98i] Mark Johnson. JavaBeans: Turn Java classes into JavaBeans. JavaWorld: IDG’s magazine for the Java community, 3(6):??, June 1998. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-06-1998/jw-06-beans.
                                        htm.

Johnson:1999:BJT

[Joh99] Mark Johnson. BeanLint: A JavaBeans troubleshooting tool, Part 2. Java-World: IDG’s magazine for the Java community, 4(1): ??, January 1999. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1999/jw-01-beans.
htm.

Kadel:1996:CSB

[Kad96a] Rich Kadel. Case study:

Building a commercial-quality Web application in Java. JavaWorld: IDG’s magazine for the Java community, 1 (9):??, November 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1996/jw-10-appfoundry.
htm.

Kadel:1996:GCS

[Kad96b] Rich Kadel. Generic clientserver classes. JavaWorld: IDG’s magazine for the Java community, 1(7):??, September 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1996/jw-09-client-
server.htm.

Kadel:1996:JMMa

[Kad96c] Rich Kadel. Java: much more than a language. Java-World: IDG’s magazine for the Java community, 1 (4):??, June 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/

                                        jw-06-1996/jw-06-javaone-
                                        kadel.htm.

Kadel:1996:JMMb

[Kad96d] Rich Kadel. Java: much more than a language. Java-World: IDG’s magazine for the Java community, 1 (5):??, July 1996. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-07-1996/jw-07-javaone-
                                        kadel.htm.

Kadel:1997:ANG

[Kad97a] Rich Kadel. AWT: The next generation. Java-World: IDG’s magazine for the Java community, 2 (5):??, May 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-05-1997/jw-05-jo-awt.
                                        htm.

Kadel:1997:IGR

[Kad97b] Rich Kadel. The importance of getting on the right road to a fast Java. JavaWorld: IDG’s magazine for the Java community, 2(5):??, May 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-05-1997/jw-05-jo-performance.
                                        htm.

Kadel:1997:LHE

[Kad97c] Rich Kadel. Learn how to extend the AWT with your own image buttons.

JavaWorld: IDG’s magazine for the Java community, 2 (3):??, March 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-03-1997/jw-03-imagebutton.
htm.

Kadel:1997:PCN

[Kad97d]  Rich Kadel.  The pros

and cons of the new Java. JavaWorld: IDG’s magazine for the Java community, 2(5):??, May 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-05-1997/jw-05-javaone-
kadel.htm.

Kadel:1997:TYA

[Kad97e] Rich Kadel. Test your applet the easy way: turn it into an application. Java-World: IDG’s magazine for the Java community, 2(1): ??, January 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1997/jw-01-applet.
to.app.htm.

Kalin:1996:JJO

[Kal96a] Sari Kalin. Java jumps to operating systems. Java-World: IDG’s magazine for the Java community, 1 (3):??, May 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/

                                        jw-05-1996/jw-05-license.
                                        htm.

Kalin:1996:NIDa

[Kal96b] Sari Kalin. News: IBM does Java jive. JavaWorld: IDG’s magazine for the Java community, 1(4):??, June 1996. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-06-1996/jw-06-ibmport.
                                        htm.

Kalin:1996:NIDb

[Kal96c] Sari Kalin. News: IBM draws on Java for OS/2 Warp 4. JavaWorld: IDG’s magazine for the Java community, 1(9):??, November 1996. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-10-1996/jw-10-os2.htm.

Kalin:1996:NOD

[Kal96d] Sari Kalin. News: Oracle discloses plan for Java-based personal applications. Java-World: IDG’s magazine for the Java community, 1(9): ??, November 1996. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-10-1996/jw-10-oracle.
                                        htm.

Kalin:1996:NSA

[Kal96e] Sari Kalin. News: Sun and Apple plan intranet, multimedia alliance. Java-World: IDG’s magazine for the Java community, 1(9):

??, November 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1996/jw-10-networld.
htm.

Kalin:1996:OPC

[Kal96f] Sari Kalin. Online publishers cookin’ carefully with Java. JavaWorld: IDG’s magazine for the Java community, 1(2):??, April 1996. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-04-1996/jw-04-realworld.
htm.

Kalin:1997:JNR

[Kal97a]  Sari Kalin.  Java, ’Net

revenues don’t yet match hype, IDC says. Java-World: IDG’s magazine for the Java community, 2(1): ??, January 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1997/jw-01-idgns.
java-idc.htm.

Kalin:1997:NIWb

[Kal97b]  Sari Kalin.  News: In-

ternet World: Microsoft, Sun dispute invitation to join pure Java. Java-World: IDG’s magazine for the Java community, 2(1): ??, January 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/

                                        jw-01-1997/jw-01-idgns.
                                        pure.htm.

Kalin:1997:NIWa

[Kal97c] Sari Kalin. News: Internet World: Microsoft, Sun, Oracle, and more. Java-World: IDG’s magazine for the Java community, 2(1): ??, January 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-01-1997/jw-01-idgns.
                                        iworld.htm.

Kalin:1997:NJS

[Kal97d] Sari Kalin. News: JavaSoft to sell JavaOS for ARM processors. JavaWorld: IDG’s magazine for the Java community, 2(3):??, March 1997. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-03-1997/jw-03-idgns.
                                        arm.htm.

Kalin:1997:NMP

[Kal97e] Sari Kalin. News: Microsoft plans Java virtual machine for Unix, WinCE by 1998. JavaWorld: IDG’s magazine for the Java community, 2(5):??, May 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-05-1997/jw-05-infoworld.
                                        microsoftvm.htm.

Kestenbaum:1996:TMG

[Kes96]  Peter Kestenbaum.  The

ten-minute guide to Java.

JavaWorld: IDG’s magazine for the Java community, 1 (1):??, March 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-03-1996/jw-03-10minute.
java.htm.

Knudsen:1997:EJJ

[Knu97a]  Craig Knudsen.  Every-

thing Java: JavaWorld net news central. Java-World: IDG’s magazine for the Java community, 2(1): ??, January 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
netnews/netnews.index.
htm.

Knudsen:1997:JNNg

[Knu97b] Craig Knudsen. JavaWorld net news central. Java-World: IDG’s magazine for the Java community, 2(8): ??, August 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
netnews/netnews.index.
htm.

Knudsen:1997:JNNj

[Knu97c] Craig Knudsen. JavaWorld net news central. Java-World: IDG’s magazine for the Java community, 2(11): ??, November 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/

                                        netnews/netnews.index.
                                        htm.

Knudsen:1997:NJNa

[Knu97d] Craig Knudsen. News: Java-World net news central. JavaWorld: IDG’s magazine for the Java community, 2 (2):??, February 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        netnews/netnews.index.
                                        htm.

Knudsen:1997:NJNb

[Knu97e] Craig Knudsen. News: Java-World net news central. JavaWorld: IDG’s magazine for the Java community, 2 (3):??, March 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        netnews/netnews.index.
                                        htm.

Knudsen:1997:NJNc

[Knu97f] Craig Knudsen. News: Java-World net news central. JavaWorld: IDG’s magazine for the Java community, 2 (4):??, April 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        netnews/netnews.index.
                                        htm.

Knudsen:1997:NJNd

[Knu97g] Craig Knudsen. News: Java-World net news central. JavaWorld: IDG’s magazine for the Java community, 2(5):??, May 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
netnews/netnews.index.
htm.

Knudsen:1997:NJNe

[Knu97h] Craig Knudsen. News: Java-World net news central. JavaWorld: IDG’s magazine for the Java community, 2(6):??, June 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
netnews/netnews.index.
htm.

Knudsen:1997:NJNf

[Knu97i] Craig Knudsen. News: Java-World net news central. JavaWorld: IDG’s magazine for the Java community, 2(7):??, July 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
netnews/netnews.index.
htm.

Knudsen:1997:NJNg

[Knu97j] Craig Knudsen. News: Java-World net news central. JavaWorld: IDG’s magazine for the Java community, 2 (9):??, September 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
netnews/netnews.index.
htm.

Knudsen:1997:NJNh

[Knu97k] Craig Knudsen. News: Java-World net news central. JavaWorld: IDG’s magazine for the Java community, 2 (10):??, October 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        netnews/netnews.index.
                                        htm.

Knudsen:1997:NJNi

[Knu97l] Craig Knudsen. News: Java-World net news central. JavaWorld: IDG’s magazine for the Java community, 2 (12):??, December 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        netnews/netnews.index.
                                        htm.

Knudsen:1998:JNNg

[Knu98a] Craig Knudsen. JavaWorld net news central. Java-World: IDG’s magazine for the Java community, 3 (7):??, July 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        netnews/netnews.index.
                                        htm.

Knudsen:1998:JNNh

[Knu98b] Craig Knudsen. JavaWorld net news central. Java-World: IDG’s magazine for the Java community, 3(8): ??, August 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
netnews/netnews.index.
htm.

Knudsen:1998:JNNi

[Knu98c] Craig Knudsen. JavaWorld net news central. Java-World: IDG’s magazine for the Java community, 3(9): ??, September 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
netnews/netnews.index.
htm.

Knudsen:1998:JNNj

[Knu98d] Craig Knudsen. JavaWorld net news central. JavaWorld: IDG’s magazine for the Java community, 3(10):??, 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
netnews/netnews.index.
htm.

Knudsen:1998:JNNk

[Knu98e] Craig Knudsen. JavaWorld net news central. JavaWorld: IDG’s magazine for the Java community, 3(11):??, 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
netnews/netnews.index.
htm.

Knudsen:1998:JNNl

[Knu98f] Craig Knudsen. JavaWorld net news central. JavaWorld: IDG’s magazine for the Java community, 3(12):??, 1998.

CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        netnews/netnews.index.
                                        htm.

Knudsen:1998:NJNa

[Knu98g] Craig Knudsen. News: Java-World Net News Central. JavaWorld: IDG’s magazine for the Java community, 3 (1):??, January 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        netnews/netnews.index.
                                        htm.

Knudsen:1998:NJNb

[Knu98h] Craig Knudsen. News: Java-World Net News Central. JavaWorld: IDG’s magazine for the Java community, 3 (2):??, February 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        netnews/netnews.index.
                                        htm.

Knudsen:1998:NJNc

[Knu98i] Craig Knudsen. News: Java-World Net News Central. JavaWorld: IDG’s magazine for the Java community, 3 (3):??, March 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        netnews/netnews.index.
                                        htm.

Knudsen:1998:NJNd

[Knu98j] Craig Knudsen. News: Java-World Net News Central. JavaWorld: IDG’s magazine for the Java community, 3 (4):??, April 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
netnews/netnews.index.
htm.

Knudsen:1998:NJNe

[Knu98k] Craig Knudsen. News: Java-World net news central. JavaWorld: IDG’s magazine for the Java community, 3(5):??, May 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
netnews/netnews.index.
htm.

Knudsen:1998:NJNf

[Knu98l] Craig Knudsen. News: Java-World net news central. JavaWorld: IDG’s magazine for the Java community, 3(6):??, June 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
netnews/netnews.index.
htm.

Knudsen:1999:JNN

[Knu99] Craig Knudsen. JavaWorld net news central. Java-World: IDG’s magazine for the Java community, 4(1): ??, January 1999. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        netnews/netnews.index.
                                        htm.

Kriens:1998:PI

[Kri98a]  Peter Kriens.  Perspec-

tives: On interfaces. Java-World: IDG’s magazine for the Java community, 3(2): ??, February 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-02-1998/jw-02-perspectives.
                                        htm.

Krill:1998:JDS

[Kri98b] Paul Krill. Java databases sprouting slowly. Java-World: IDG’s magazine for the Java community, 3 (5):??, May 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-05-1998/jw-05-iw.databases.
                                        htm.

Krill:1998:NSE

[Kri98c] Paul Krill. News: Sybase enhances Java DB driver, Web transactions. Java-World: IDG’s magazine for the Java community, 3 (3):??, March 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-03-1998/jw-03-infoworld.
                                        sybase.htm.

Krumel:1998:RRD

[Kru98] Andy Krumel. Revolutionary RMI: Dynamic class loading and behavior objects. JavaWorld: IDG’s magazine for the Java community, 3(12):??, 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-12-1998/jw-12-enterprise.
htm.

Kujubu:1998:NSZ

[Kuj98] Laura Kujubu. News: Sun’s Zander sees networks everywhere. JavaWorld: IDG’s magazine for the Java community, 3(4):??, April 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-04-1998/jw-04-infoworld.
zander.htm.

Kurr:1997:JTW

[Kur97] Ron Kurr. Java tip 42: Write Java apps that work with proxy-based firewalls. Java-World: IDG’s magazine for the Java community, 2(12): ??, December 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javatips/jw-javatip42.
htm.

LaGrow:1996:CZJ

[LaG96a] Craig LaGrow. Cyberstruction zone: Java fires up its database engines. Java-World: IDG’s magazine for the Java community, 1 (3):??, May 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-05-1996/jw-05-cszone.
htm.

LaGrow:1996:CZ

[LaG96b] Craig LaGrow. Cyberstruction zone: Java meets the Energizer Bunny. Java-World: IDG’s magazine for the Java community, 1 (1):??, March 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-03-1996/jw-03-cyberstruction.
htm.

LaGrow:1996:CZS

[LaG96c] Craig LaGrow. Cyberstruction zone: Shockwave and Java meet on the ’Net. JavaWorld: IDG’s magazine for the Java community, 1 (2):??, April 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-04-1996/lagrow.0496/
jw-04-cszone.htm.

Lawton:1996:ABJ

[Law96a]  George Lawton.  Ap-

pliances brew Java with Kona (JavaOS). Java-World: IDG’s magazine for the Java community, 1 (5):??, July 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/

jw-07-1996/jw-07-appliances-
lawton.htm.

Lawton:1996:JFT

[Law96b]  George Lawton.  Java

finds telecom niche. Java-World: IDG’s magazine for the Java community, 1 (5):??, July 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1996/jw-07-telecom-
lawton.htm.

Lawton:1996:NIB

[Law96c] George Lawton. News: IBM brings Java to Windows 3.1. JavaWorld: IDG’s magazine for the Java community, 1 (9):??, November 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1996/jw-10-ibm.htm.

Lawton:1996:NJBa

[Law96d] George Lawton. News: Java Beans glue applets together. JavaWorld: IDG’s magazine for the Java community, 1(4):??, June 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-06-1996/jw-06-beans.
htm.

Lawton:1996:NJBb

[Law96e] George Lawton. News: Java Beans glue applets together. JavaWorld: IDG’s magazine for the Java community, 1(5):??, July 1996. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-07-1996/jw-07-beans.
                                        htm.

Lawton:1996:NJG

[Law96f]  George Lawton.  News:

Java gets new APIs. Java-World: IDG’s magazine for the Java community, 1 (5):??, July 1996. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-07-1996/jw-07-apis-
                                        lawton.htm.

Lawton:1996:NLB

[Law96g] George Lawton. News: Lucent brews Inferno. Java-World: IDG’s magazine for the Java community, 1(9): ??, November 1996. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-10-1996/jw-10-inferno.
                                        htm.

Lawton:1996:NOPa

[Law96h] George Lawton. News: Office productivity comes to Java. JavaWorld: IDG’s magazine for the Java community, 1(4):??, June 1996. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-06-1996/jw-06-corel.
                                        htm.

Lawton:1996:NOPb

[Law96i] George Lawton. News: Office productivity comes to Java. JavaWorld: IDG’s magazine for the Java community, 1(5):??, July 1996. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1996/jw-07-corel.
htm.

Lawton:1996:NSF

[Law96j] George Lawton. News: San Francisco Project takes Java to the enterprise. Java-World: IDG’s magazine for the Java community, 1(9): ??, November 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1996/jw-10-sfproject.
htm.

Leach:1996:WBH

[Lea96] George W. Leach. Book review: Will this book hook you in? Java-World: IDG’s magazine for the Java community, 1(6): ??, August 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-08-1996/jw-08-bookrev-
hooked.htm.

Lenaham:1997:JTI

[Len97] Peter Lenaham. Java tips:

Improve your applet’s download performance using Navigator’s Java console. Java-World: IDG’s magazine for the Java community, 2 (4):??, April 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javatips/jw-javatip28.
htm.

Lambel:1997:NLF

[LH97] François Lambel and Elizabeth Heichler. News: Liberté, fraternité… et Java? JavaWorld: IDG’s magazine for the Java community, 2 (10):??, October 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1997/jw-10-idgns.
allezjava.htm.

Liederman:1996:JRW

[Lie96a]  Erica Liederman.  Java

in the real world. Java-World: IDG’s magazine for the Java community, 1 (1):??, March 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-03-1996/jw-03-realworld.
htm.

Liederman:1996:JSC

[Lie96b]  Erica Liederman.  Java-

Soft sets its course. Java-World: IDG’s magazine for the Java community, 1 (2):??, April 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/

jw-04-1996/jw-04-javasoft.
htm.

Liang:1998:CSA

[LLZ98] Chihong Liang, Kevin Luo, and Tony Zhang. Cover story: AFC vs. JFC: Which GUI toolkit should you use? JavaWorld: IDG’s magazine for the Java community, 3 (2):??, February 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-02-1998/jw-02-afc.v.
jfc.htm.

Lopez:1997:JTH

[Lop97] Albert Lopez. Java tips: How to decouple the Observer/ Observable object model. JavaWorld: IDG’s magazine for the Java community, 2(5):??, May 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javatips/jw-javatip29.
htm.

Loy:1999:SCS

[Loy99] Marc Loy. Start customizing Swing’s editor pane — patch the Swing HTMLEditorKit. JavaWorld: IDG’s magazine for the Java community, 4 (1):??, January 1999. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1999/jw-01-swing.
htm.

Laird:1997:NSR

[LS97] Cameron Laird and Kathryn Soraiz. News: Sun releases scripting technology for Java. JavaWorld: IDG’s magazine for the Java community, 2 (12):??, December 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-12-1997/jw-12-jacl.
                                        htm.

Mahmoud:1997:JTP

[Mah97] Qusay H. Mahmoud. Java tips: Protect your bytecodes from reverse engineering/decompilation. Java-World: IDG’s magazine for the Java community, 2(1): ??, January 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        javatips/jw-javatip22.
                                        htm.

Makower:1996:MMI

[Mak96]  Dave Makower.  March

of the Mac IDEs. Java-World: IDG’s magazine for the Java community, 1(6): ??, August 1996. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-08-1996/jw-08-mac-ides.
                                        htm.

Makower:1997:FMB

[Mak97]  Dave Makower.  Cover

story: Four Mac-based IDEs compared: Which should you choose? Java-World: IDG’s magazine for the Java community, 2(9): ??, September 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1997/jw-09-mac-ides.
htm.

Mangione:1998:PTS

[Man98] Carmine Mangione. Performance tests show Java as fast as C++. Java-World: IDG’s magazine for the Java community, 3(2): ??, February 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-02-1998/jw-02-jperf.
htm.

Matsumura:1997:DQK

[Mat97a] Miko Matsumura. Miko’s vision: Demo or die! the quest for the killer app. Java-World: IDG’s magazine for the Java community, 2(12): ??, December 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-12-1997/jw-12-miko.
htm.

Matsumura:1997:NCU

[Mat97b]  Miko Matsumura.  New

column: Ultranet, the next network. JavaWorld: IDG’s magazine for the Java community, 2(11):??, November 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-11-1997/jw-11-miko.
                                        htm.

Matsumura:1998:MVR

[Mat98a] Miko Matsumura. Miko’s vision: The real future of Java. JavaWorld: IDG’s magazine for the Java community, 3 (1):??, January 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-01-1998/jw-01-miko.
                                        htm.

Matsumura:1998:MVX

[Mat98b] Miko Matsumura. Miko’s vision: XML speeds along in standards land. Java-World: IDG’s magazine for the Java community, 3(2): ??, February 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-02-1998/jw-02-miko.
                                        htm.

McKay:1997:NSH

[MB97] Niall McKay and Ted Smalley Bowen. News: Sun hands off Java testing to outside lab. JavaWorld: IDG’s magazine for the Java community, 2(11):??, November 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-11-1997/jw-11-infoworld.
                                        testing.htm.

McKay:1998:NSR

[MB98] Niall McKay and Ted Smalley Bowen. News: Sun readies JavaBeans scripting tools. JavaWorld: IDG’s magazine for the Java community, 3 (2):??, February 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-02-1998/jw-02-infoworld.
javabeans.htm.

McKay:1998:SMJ

[MC98a] Niall McKay and Jon Cornetto. Sun to merge Java tools in two-stage integration. JavaWorld: IDG’s magazine for the Java community, 3(7):??, July 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1998/jw-07-iw-mergejava.
htm.

Mitchell:1998:JTH

[MC98b] John D. Mitchell and Arthur Choi. Java tip 49: How to extract Java resources from JAR and zip archives. JavaWorld: IDG’s magazine for the Java community, 3 (3):??, March 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javatips/jw-javatip49.
htm.

McCarthy:1996:SAT

[McC96a] Michael McCarthy. Sun announces three-day Java conference. JavaWorld: IDG’s magazine for the Java community, 1(1):??, March 1996. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-03-1996/jw-03-javaone.
htm.

McComb:1996:JDJ

[McC96b] Gordon McComb. JavaScript: Debugging JavaScript programs. JavaWorld: IDG’s magazine for the Java community, 1(5):??, July 1996. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1996/jw-07-javascript.
htm.

McComb:1996:JPJ

[McC96c] Gordon McComb. JavaScript: ‘personalized JavaScript’: User-defined functions, objects, and methods. Java-World: IDG’s magazine for the Java community, 1(7): ??, September 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1996/jw-09-javascript.
htm.

McComb:1996:JUJb

[McC96d] Gordon McComb. JavaScript: Using JavaScript and forms. JavaWorld: IDG’s magazine for the Java community, 1(4):??, June 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/

jw-06-1996/jw-06-javascript.
htm.

McComb:1996:JUJc

[McC96e] Gordon McComb. JavaScript:

Using JavaScript and graphics. JavaWorld: IDG’s magazine for the Java community, 1(6):??, August 1996. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-08-1996/jw-08-javascript.
htm.

McComb:1996:JUJa

[McC96f] Gordon McComb. JavaScript:

Using JavaScript’s built-in objects. JavaWorld: IDG’s magazine for the Java community, 1(3):??, May 1996. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-05-1996/jw-05-javascript.
htm.

McComb:1996:NJF

[McC96g]  Gordon McComb.  New

JavaScript features in Navigator 3.0. JavaWorld: IDG’s magazine for the Java community, 1(9):??, November 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1996/jw-10-javascript.
htm.

McComb:1996:NNI

[McC96h] Gordon McComb. News:

Netscape introduces new ‘privacy’ bug. JavaWorld:

IDG’s magazine for the Java community, 1(9):??, November 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1996/jw-10-jsbug.
htm.

McComb:1997:JMY

[McC97a] Gordon McComb. JavaScript: Make your life easier: Create reusable routines in JavaScript. JavaWorld: IDG’s magazine for the Java community, 2(2):??, February 1997. CODEN ???? ISSN 1091-8906. URL http:

//www.javaworld.com/javaworld/
jw-02-1997/jw-02-javascript.
htm.

McComb:1997:JMF

[McC97b] Gordon McComb. JavaScript: More frequently sought solutions. JavaWorld: IDG’s magazine for the Java community, 2(3):??, March 1997. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-03-1997/jw-03-javascript.
htm.

McComb:1997:JTA

[McC97c] Gordon McComb. JavaScript: Take advantage of userdefined variables. Java-World: IDG’s magazine for the Java community, 2(1): ??, January 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/

jw-01-1997/jw-01-javascript.
htm.

McCulley:1998:FS

[McC98] Mark McCulley. Focus on Swing. JavaWorld: IDG’s magazine for the Java community, 3(7):??, July 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1998/jw-07-swing-
focus.htm.

McGraw:1997:NYB

[McG97] Gary McGraw. News: Is your browser a blabbermouth? are your ports being scanned? JavaWorld: IDG’s magazine for the Java community, 2(3):??, March 1997. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-03-1997/jw-03-securityholes.
htm.

McKay:1997:NJD

[McK97a] Niall McKay. News: Java drills further into embedded OS arena. Java-World: IDG’s magazine for the Java community, 2(10): ??, October 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1997/jw-10-infoworld.
embedded.htm.

McKay:1997:NJSb

[McK97b] Niall McKay. News: Java to be standardized by ISO. JavaWorld: IDG’s magazine for the Java community, 2 (12):??, December 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-12-1997/jw-12-infoworld.
iso.htm.

McKay:1997:NJA

[McK97c] Niall McKay. News: Java-Soft alters JDK schedule to ensure portability and lack of bugs. Java-World: IDG’s magazine for the Java community, 2(10): ??, October 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1997/jw-10-infoworld.
jdk.htm.

McKay:1997:NNA

[McK97d] Niall McKay. News: Netscape’s Andreessen lays out Java enterprise strategy. Java-World: IDG’s magazine for the Java community, 2 (7):??, July 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1997/jw-07-idgns.
andreessen.htm.

McKay:1997:NRF

[McK97e] Niall McKay. News: RMI’s future cloudy as Sun tries to please CORBA partners. JavaWorld: IDG’s magazine for the Java community, 2(6):??, June 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.

javaworld.com/javaworld/
jw-06-1997/jw-06-idgns.
rmi.htm.

McKay:1997:NSB

[McK97f] Niall McKay. News: Sun to brew Enterprise JavaBeans. JavaWorld: IDG’s magazine for the Java community, 2 (8):??, August 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-08-1997/jw-08-idgns-
brew.htm.

McKay:1998:JFC

[McK98a] Niall McKay. JavaSoft faces competition from freeware. JavaWorld: IDG’s magazine for the Java community, 3(5):??, May 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-05-1998/jw-05-infoworld.
jvm.htm.

McKay:1998:MIK

[McK98b] Niall McKay. Metamata, IBM keep tabs on Java performance. JavaWorld: IDG’s magazine for the Java community, 3(7):??, July 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1998/jw-07-iw-metamata.
htm.

McKay:1998:NIS

[McK98c] Niall McKay. News: IBM scales up Java plans. Java-World: IDG’s magazine for the Java community, 3 (3):??, March 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-03-1998/jw-03-idgns.
ibm.htm.

McKay:1998:SRC

[McK98d] Niall McKay. Sun reorganizes company structure. JavaWorld: IDG’s magazine for the Java community, 3(5):??, May 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-05-1998/jw-05-infoworld.
reorg.htm.

McKay:1998:SWJ

[McK98e] Niall McKay. Sun walls off Java. JavaWorld: IDG’s magazine for the Java community, 3(6):??, June 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-06-1998/jw-06-iw-sun.
htm.

McKay:1998:SJD

[McK98f] Niall McKay. Sun’s JavaSoft division faces widespread reorganization. JavaWorld: IDG’s magazine for the Java community, 3(5):??, May 1998. CODEN ???? ISSN 1091-8906. URL http://

www.javaworld.com/javaworld/
jw-05-1998/jw-05-infoworld.
javasoft.htm.

McManis:1996:JIP

[McM96a]  Chuck McManis.  Java

IDEs for the PC. Java-World: IDG’s magazine for the Java community, 1(6): ??, August 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-08-1996/jw-08-win95-
ides.htm.

McManis:1996:JDUb

[McM96b] Chuck McManis. Java in depth: Not using garbage collection. JavaWorld: IDG’s magazine for the Java community, 1(7):??, September 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1996/jw-09-indepth.
htm.

McManis:1996:JDSa

[McM96c]  Chuck McManis.  Java

in depth: Synchronizing threads in Java. Java-World: IDG’s magazine for the Java community, 1 (2):??, April 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-04-1996/jw-04-synch.
htm.

McManis:1996:JDSb

[McM96d]  Chuck McManis.  Java

in depth: Synchronizing threads in Java, Part II. JavaWorld: IDG’s magazine for the Java community, 1(3):??, May 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-05-1996/jw-05-mcmanis.
htm.

McManis:1996:JDB

[McM96e]  Chuck McManis.  Java

in depth: The basics of Java class loaders. Java-World: IDG’s magazine for the Java community, 1(9): ??, November 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1996/jw-10-indepth.
htm.

McManis:1996:JDT

[McM96f] Chuck McManis. Java in depth: Threads and applets and visual controls. JavaWorld: IDG’s magazine for the Java community, 1(5):??, July 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1996/jw-07-mcmanis.
htm.

McManis:1996:JDUa

[McM96g] Chuck McManis. Java in depth: Using communication channels in applets, Part 3. JavaWorld: IDG’s magazine for the Java community, 1(4):??, June 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/

jw-06-1996/jw-06-mcmanis.
htm.

McManis:1997:JDF

[McM97a] Chuck McManis. Java in depth: A first look at Borland’s JBuilder IDE. Java-World: IDG’s magazine for the Java community, 2(1): ??, January 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-11-1997/jw-11-howto.
html.

McManis:1997:JDLc

[McM97b] Chuck McManis. Java in depth: A look at inner classes. JavaWorld: IDG’s magazine for the Java community, 2(1):??, January 1997. CODEN ???? ISSN 1091-8906. URL http://

www.javaworld.com/javaworld/
jw-10-1997/jw-10-indepth.
html.

McManis:1997:JDB

[McM97c] Chuck McManis. Java in depth: Build an interpreter in Java — implement the execution engine. Java-World: IDG’s magazine for the Java community, 2 (7):??, July 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1997/jw-07-indepth.
html.

McManis:1997:JDHa

[McM97d] Chuck McManis. Java in depth: How to build an interpreter in Java, part 1: The BASICs. JavaWorld: IDG’s magazine for the Java community, 2(5):??, May 1997. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-05-1997/jw-05-indepth.
                                        htm.

McManis:1997:JDHb

[McM97e] Chuck McManis. Java in depth: How to build an interpreter in Java: The structure. JavaWorld: IDG’s magazine for the Java community, 2(5):??, May 1997. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-05-1997/jw-05-indepth.
                                        html.

McManis:1997:JDLa

[McM97f]  Chuck McManis.  Java

in depth: Lexical analysis and Java: Part 1. Java-World: IDG’s magazine for the Java community, 2(1): ??, January 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        /javaworld/jw-01-1997/
                                        jw-01-indepth.html.

McManis:1997:JDLb

[McM97g] Chuck McManis. Java in depth: Lexical analysis, part 2: Build an application.

JavaWorld: IDG’s magazine for the Java community, 2 (2):??, February 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-02-1997/jw-02-indepth.
html.

McManis:1997:JDTa

[McM97h] Chuck McManis. Java in depth: Take a look inside Java classes. Java-World: IDG’s magazine for the Java community, 2(8): ??, August 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-08-1997/jw-08-indepth.
html.

McManis:1997:JDTb

[McM97i] Chuck McManis. Java in depth: Take an in-depth look at the Java Reflection API. JavaWorld: IDG’s magazine for the Java community, 2 (1):??, January 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1997/jw-09-indepth.
html.

McMillan:1997:NDA

[McM97j] Robert McMillan. News:

Diba acquisition puts Sun in toaster business. Java-World: IDG’s magazine for the Java community, 2(8): ??, August 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.

                                        javaworld.com/javaworld/
                                        jw-08-1997/jw-08-diba.
                                        htm.

McMillan:1997:NHH

[McM97k] Robert McMillan. News: Haven’t heard about Sun’s new JavaNFS and ‘Project Studio?’. JavaWorld: IDG’s magazine for the Java community, 2(4):??, April 1997. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-04-1997/jw-04-sugwest.
                                        htm.

McMillan:1997:NJI

[McM97l] Robert McMillan. News: Java Internet Business Expo: Reality check for Java. Java-World: IDG’s magazine for the Java community, 2(9): ??, September 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-09-1997/jw-09-jibe.
                                        bob.htm.

McMillan:1997:NMJ

[McM97m] Robert McMillan. News: Microsoft’s J/Direct called death of Java. Java-World: IDG’s magazine for the Java community, 2 (7):??, July 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-07-1997/jw-07-jdirect.
                                        htm.

McMillan:1997:NSB

[McM97n] Robert McMillan. News:

Sun boosts Java performance, adding JIT compiler and JVM with multithreading to Solaris 2.6. JavaWorld: IDG’s magazine for the Java community, 2(7):??, July 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1997/jw-07-speedway.
htm.

McMillan:1997:NSS

[McM97o] Robert McMillan. News:

Sun switches course with new Java chip. Java-World: IDG’s magazine for the Java community, 2(11): ??, November 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-11-1997/jw-11-microjava.
htm.

McMillan:1997:NSP

[McM97p] Robert McMillan. News:

Sun’s plan B for Java standardization: change nothing. JavaWorld: IDG’s magazine for the Java community, 2(10):??, October 1997. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1997/jw-10-iso.htm.

McManis:1998:DUT

[McM98a] Chuck McManis. In depth:

Using threads with collections, Part 1. Java-World: IDG’s magazine for the Java community, 3 (3):??, March 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-03-1998/jw-03-indepth.
html.

McManis:1998:JDD

[McM98b] Chuck McManis. Java in depth: An in-depth look at Java’s character type. Java-World: IDG’s magazine for the Java community, 3(1): ??, January 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1998/jw-01-indepth.
html.

McManis:1998:JDU

[McM98c] Chuck McManis. Java in depth: Using threads with collections, part 2. Java-World: IDG’s magazine for the Java community, 3 (6):??, June 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-06-1998/jw-06-indepth.
html.

McManis:1998:MEJ

[McM98d] Chuck McManis. My ENIG-MAtic Java ring. Java-World: IDG’s magazine for the Java community, 3(8): ??, August 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.

javaworld.com/javaworld/
jw-08-1998/jw-08-indepth.
htm.

McManis:1998:ONP

[McM98e] Chuck McManis. Opening up new ports to Java. Java-World: IDG’s magazine for the Java community, 3(9): ??, September 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1998/jw-09-indepth.
htm.

McManis:1998:RFFa

[McM98f] Chuck McManis. Ring fever:

A first-hand look at Javapowered jewelry. Java-World: IDG’s magazine for the Java community, 3 (4):??, April 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-04-1998/jw-04-ringfever.
htm.

McManis:1998:RFFb

[McM98g] Chuck McManis. Ring fever:

A first-hand look at Javapowered jewelry. Java-World: IDG’s magazine for the Java community, 3 (5):??, May 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-04-1998/jw-04-ringfever.
htm.

McMillan:1998:BGM

[McM98h] Robert McMillan. Baratz gives Microsoft strategy lesson. JavaWorld: IDG’s magazine for the Java community, 3(6):??, June 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-06-1998/jw-06-baratz.
                                        htm.

McMillan:1998:NDS

[McM98i] Robert McMillan. News: Does Sun want to become the next Microsoft? Java-World: IDG’s magazine for the Java community, 3 (3):??, March 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-03-1998/jw-03-lighthouse.
                                        htm.

McMillan:1998:NWR

[McM98j] Robert McMillan. News: Who is right in the Sun– Microsoft lawsuit? Sun-World posts results of reader survey. Java-World: IDG’s magazine for the Java community, 3(1): ??, January 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-01-1998/jw-01-lawsuit.
                                        results.htm.

Mysore:1998:JDJ

[MD98] Shivaram H. Mysore and Rinaldo Di Giorgio. Java developer: Java gets serial support

with the new javax.comm

package. JavaWorld: IDG’s magazine for the Java community, 3(5):??, May 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-05-1998/jw-05-javadev.
htm.

Mills:1996:NJJa

[ME96a] Elinor Mills and Kristi Essick. News: JavaOne: Java industry comes into its own. JavaWorld: IDG’s magazine for the Java community, 1(4):??, June 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-06-1996/jw-06-java2.
htm.

Mills:1996:NJJb

[ME96b] Elinor Mills and Kristi Essick. News: JavaOne: Java industry comes into its own. JavaWorld: IDG’s magazine for the Java community, 1(5):??, July 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1996/jw-07-java2.
htm.

Mills:1996:NSUa

[ME96c] Elinor Mills and Kristi Essick. News: Sun unveils JavaOS and licensees at JavaOne. JavaWorld: IDG’s magazine for the Java community, 1(4):??, June 1996.

CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-06-1996/jw-06-java1.
                                        htm.

Mills:1996:NSUb

[ME96d] Elinor Mills and Kristi Essick. News: Sun unveils JavaOS and licensees at JavaOne. JavaWorld: IDG’s magazine for the Java community, 1(5):??, July 1996. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-07-1996/jw-07-java1.
                                        htm.

Messmer:1996:NSS

[Mes96] Ellen Messmer. News: Sun-Soft ships Java Workshop. JavaWorld: IDG’s magazine for the Java community, 1 (7):??, September 1996. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-09-1996/jw-09-workshop.
                                        htm.

Meshorer:1998:AUR

[Mes98a] Tomer Meshorer. Add an undo/redo function to your Java application with Swing. JavaWorld: IDG’s magazine for the Java community, 3(6):??, June 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-06-1998/jw-06-undoredo.
                                        htm.

Meshorer:1998:CHS

[Mes98b] Tomer Meshorer. Capture hierarchical structures with Swing’s JTree. JavaWorld: IDG’s magazine for the Java community, 3(10):??, 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1998/jw-10-jtree.
htm.

McGraw:1997:UKJ

[MF97] Gary McGraw and Edward Felten. Understanding the keys to Java security: the sandbox and authentication. JavaWorld: IDG’s magazine for the Java community, 2(5):??, May 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-05-1997/jw-05-security.
htm.

McGraw:1998:TRD

[MF98a] Gary McGraw and Edward Felten. Twelve rules for developing more secure Java code. JavaWorld: IDG’s magazine for the Java community, 3(12):??, 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-12-1998/jw-12-securityrules.
htm.

Mitchell:1998:JTS

[MF98b] John D. Mitchell and Bill Foote. Java tip 52: Singletons vs. class (un)loading.

JavaWorld: IDG’s magazine for the Java community, 3(5):??, May 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javatips/jw-javatip52.
htm.

Mills:1996:JHT

[Mil96a] Elinor Mills. Java hot topic of Software Development ’96. JavaWorld: IDG’s magazine for the Java community, 1 (2):??, April 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-04-1996/jw-04-sd96.
htm.

Mills:1996:JEC

[Mil96b] Elinor Mills. Java in OS to ease cross-platform development. JavaWorld: IDG’s magazine for the Java community, 1(3):??, May 1996. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-05-1996/jw-05-license2.
htm.

Mills:1996:NGA

[Mil96c] Elinor Mills. News: Grappling with applet licensing. JavaWorld: IDG’s magazine for the Java community, 1(3):??, May 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-05-1996/jw-05-jade.
htm.

Mills:1996:SCM

[Mil96d] Elinor Mills. Spring Comdex:

McNealy says Java cuts admin costs. JavaWorld: IDG’s magazine for the Java community, 1(5):??, July 1996. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1996/jw-07-mcnealy-
comdex.htm.

Mills:1998:NJW

[Mil98a] Elinor Mills. News: Judge won’t move up hearing on Java injunction as Sun asked. JavaWorld: IDG’s magazine for the Java community, 3(7):??, July 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1998/jw-07-idgns-
injunction.htm.

Mills:1998:NNA

[Mil98b] Elinor Mills. News: Netscape’s Andreessen talks server, Java strategy. JavaWorld: IDG’s magazine for the Java community, 3(1):??, January 1998. CODEN ???? ISSN 1091-8906. URL http://

www.javaworld.com/javaworld/
jw-01-1998/jw-01-idgns.
andreessen.htm.

Mitchell:1996:JQO

[Mit96a] John Mitchell. Java Q&A:

Overcoming installation difficulties. JavaWorld: IDG’s magazine for the Java community, 1(5):??, July 1996.

CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1996/jw-07-javaqa.
htm.

Mitchell:1996:JQJ

[Mit96b] John D. Mitchell. Java Q&A: Is Java really cross-platform? JavaWorld: IDG’s magazine for the Java community, 1 (7):??, September 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1996/jw-09-javaqa.
htm.

Mitchell:1996:R

[Mit96c] John D. Mitchell. Java Q&A: References. Java-World: IDG’s magazine for the Java community, 1 (3):??, May 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-05-1996/jw-05-javaqa.
htm.

Mitchell:1996:JQA

[Mit96d] John D. Mitchell. Java Q&A: The answer to the question “will Java remain ‘open’?”. JavaWorld: IDG’s magazine for the Java community, 1 (9):??, November 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1996/jw-10-javaqa.
htm.

Mitchell:1996:JT

[Mit96e] John D. Mitchell. Java tips.

JavaWorld: IDG’s magazine for the Java community, 1 (1):??, March 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javatips/jw-javatips.index.
htm.

Mitchell:1996:JTA

[Mit96f] John D. Mitchell. Java tips:

An alternative way to enable double clicks. Java-World: IDG’s magazine for the Java community, 1 (2):??, April 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javatips/jw-javatips.index.
htm.

Mitchell:1996:JTC

[Mit96g] John D. Mitchell. Java tips:

Cascading menus. Java-World: IDG’s magazine for the Java community, 1(6): ??, August 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javatips/jw-javatip12.
htm.

Mitchell:1996:JTM

[Mit96h] John D. Mitchell. Java tips: More about threads and the resize problem. Java-World: IDG’s magazine for the Java community, 1 (4):??, June 1996. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        javatips/jw-javatip9.htm.

Mitchell:1996:JTR

[Mit96i] John D. Mitchell. Java tips: Redirecting streams. Java-World: IDG’s magazine for the Java community, 1(7): ??, September 1996. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        javatips/jw-javatip14.
                                        htm.

Mitchell:1996:RAN

[Mit96j] John D. Mitchell. Java tips: Reloading applets in Netscape. JavaWorld: IDG’s magazine for the Java community, 1(3):??, May 1996. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        javatips/jw-javatip7.htm.

Mitchell:1996:JTS

[Mit96k] John D. Mitchell. Java tips: Save typing time! JavaWorld: IDG’s magazine for the Java community, 1(5):??, July 1996. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        javatips/jw-javatip11.
                                        htm.

Mitchell:1996:NPJ

[Mit96l] John D. Mitchell. News: Preliminary JDK 1.1 design specs unveiled. Java-World: IDG’s magazine for the Java community, 1(7): ??, September 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1996/jw-09-jdknews.
htm.

Mitchell:1997:JTO

[Mit97a] John D. Mitchell. Java tip 40: Object transport via datagram packets. Java-World: IDG’s magazine for the Java community, 2(11): ??, November 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javatips/jw-javatip40.
htm.

Mitchell:1997:JTPb

[Mit97b] John D. Mitchell. Java tip 41:

POSTing via Java revisited. JavaWorld: IDG’s magazine for the Java community, 2 (11):??, November 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javatips/jw-javatip41.
htm.

Mitchell:1997:JTPa

[Mit97c] John D. Mitchell. Java tips: POSTing via Java. JavaWorld: IDG’s magazine for the Java community, 2 (8):??, August 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/

                                        javatips/jw-javatip34.
                                        htm.

Mitchell:1997:JTR

[Mit97d] John D. Mitchell. Java tips: Redirecting streams revisited. JavaWorld: IDG’s magazine for the Java community, 2(7):??, July 1997. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        javatips/jw-javatip33.
                                        htm.

Mitchell:1998:WBA

[Mit98] John D. Mitchell. Java tip 53: So what browser is this, anyway? JavaWorld: IDG’s magazine for the Java community, 3(5):??, May 1998. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        javatips/jw-javatip53.
                                        htm.

Mirsky:1998:JTC

[MM98] Gregory N. Mirsky and John D. Mitchell. Java tip 44: Calculating holidays and their observances. JavaWorld: IDG’s magazine for the Java community, 3 (1):??, January 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        javatips/jw-javatip44.
                                        htm.

McKay:1998:JLS

[MO98] Niall McKay and Michael O’Connell. Java Lobby says Microsoft official tried to ‘pull a fast one’. Java-World: IDG’s magazine for the Java community, 3 (4):??, April 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-04-1998/jw-04-fastone.
htm.

Morgan:1997:CSC

[Mor97] Bryan Morgan. Cover story:

CORBA meets Java. Java-World: IDG’s magazine for the Java community, 2(10): ??, October 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1997/jw-10-corbajava.
htm.

Morgan:1998:ACI

[Mor98a]  Bryan Morgan.  Applied

CORBA: Integrating legacy code with the Web. Java-World: IDG’s magazine for the Java community, 3(1): ??, January 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1998/jw-01-corbalegacy.
htm.

Morgan:1998:DJE

[Mor98b] Bryan Morgan. Distributed Java: Enterprise JavaBeans: Coming soon to a server near you. JavaWorld: IDG’s magazine for the Java community, 3(6):??, June 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-06-1998/jw-06-distributed.
                                        htm.

Morgan:1998:DOJ

[Mor98c] Bryan Morgan. Distributed objects: Java 1.2 extends Java’s distributed object capabilities. JavaWorld: IDG’s magazine for the Java community, 3(4):??, April 1998. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-04-1998/jw-04-distributed.
                                        htm.

Morrison:1998:MVJ

[Mor98d] Cliff Morrison. Make Visual J++ 1.1 earn its version number. JavaWorld: IDG’s magazine for the Java community, 3(3):??, March 1998. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-03-1998/jw-03-vjupgrade.
                                        htm.

McKay:1998:SSC

[MR98] Niall McKay and Lynda Radosevic. Sun the software company? Java-World: IDG’s magazine for the Java community, 3(8): ??, August 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-08-1998/jw-08-iw-sunsoftware.
                                        htm.

McKay:1997:NJSa

[MS97] Niall McKay and Bret Sommers. News: JavaSoft to solve RMI’s interoperability problem. JavaWorld: IDG’s magazine for the Java community, 2(7):??, July 1997. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1997/jw-07-idgns.
rmi.htm.

McKay:1998:NSS

[MS98] Niall McKay and Ed Scannell. News: Sun seeks injunction barring incompatible versions of Java in Windows 98. JavaWorld: IDG’s magazine for the Java community, 3(6):??, June 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-06-1998/jw-06-iw-win98.
htm.

McKay:1998:NMT

[MT98a] Niall McKay and Bob Trott.

News: Microsoft takes wraps off Visual J++ 6.0. Java-World: IDG’s magazine for the Java community, 3 (4):??, April 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-04-1998/jw-04-infoworld.
vj60.htm.

Mills:1998:MDJ

[MT98b] Elinor Mills and Bob Trott.

Microsoft is dropping Java in Mac and Unix Explorer. JavaWorld: IDG’s magazine for the Java community, 3(12):??, 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-12-1998/jw-12-idgns-
macunix.htm.

Murphy:1996:JDC

[Mur96a] Kieron Murphy. Java developers conference serves up a rich blend. Java-World: IDG’s magazine for the Java community, 1 (4):??, June 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-06-1996/jw-06-sigsconf-
murphy.htm.

Murphy:1996:NAD

[Mur96b] Kieron Murphy. News: Applets, dev tools showcased at Object/Java Expo. Java-World: IDG’s magazine for the Java community, 1(7): ??, September 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1996/jw-09-javaexpo.
htm.

Murphy:1996:NHO

[Mur96c] Kieron Murphy. News: Habanero offers a red-hot solution. JavaWorld: IDG’s magazine for the Java community, 1(5):??, July 1996. CODEN ???? ISSN 1091-

8906. URL http://www.

javaworld.com/javaworld/
jw-07-1996/jw-07-habanero.
htm.

Murphy:1996:NHP

[Mur96d] Kieron Murphy. News: HyperTV promises to synergize television and the Web. JavaWorld: IDG’s magazine for the Java community, 1(3):??, May 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-05-1996/jw-05-hypertv.
htm.

Murphy:1996:NJJ

[Mur96e]  Kieron Murphy.  News:

Java at the Javitz. Java-World: IDG’s magazine for the Java community, 1 (5):??, July 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1996/jw-07-pcexpo.
htm.

Murphy:1996:NJU

[Mur96f] Kieron Murphy. News: Java Users conference ‘goes public’ on Wall Street. Java-World: IDG’s magazine for the Java community, 1(7): ??, September 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1996/jw-09-makemoney.
htm.

Murphy:1996:NJC

[Mur96g]  Kieron Murphy.  News:

Java/C++ gateway to let Java use C++ components. JavaWorld: IDG’s magazine for the Java community, 1(3):??, May 1996. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-05-1996/jw-05-twinpeaks.
                                        htm.

Murphy:1996:NNO

[Mur96h] Kieron Murphy. News: Novell to offer full Java functionality in NetWare. Java-World: IDG’s magazine for the Java community, 1 (3):??, May 1996. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-05-1996/jw-05-novell.
                                        htm.

Murphy:1996:NSO

[Mur96i] Kieron Murphy. News: Startup offers proprietary realtime Java for embedded systems. JavaWorld: IDG’s magazine for the Java community, 1(3):??, May 1996. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-05-1996/jw-05-newmonics.
                                        htm.

Murphy:1996:NSA

[Mur96j] Kieron Murphy. News: Sun announces Java WorkShop, Joe, and related products.

JavaWorld: IDG’s magazine for the Java community, 1(3):??, May 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-05-1996/jw-05-sunprods.
htm.

Murphy:1996:PHJ

[Mur96k] Kieron Murphy. ‘real world’ Java: Prentice Hall Javatizes itself. JavaWorld: IDG’s magazine for the Java community, 1(4):??, June 1996. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-06-1996/jw-06-realworld.
htm.

Murphy:1996:WDT

[Mur96l] Kieron Murphy. So why did they decide to call it Java? JavaWorld: IDG’s magazine for the Java community, 1 (9):??, November 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1996/jw-10-javaname.
htm.

Murphy:1997:PCJ

[Mur97a] Kieron Murphy. Cover story:

The pros and cons of JDK 1.1. JavaWorld: IDG’s magazine for the Java community, 2(4):??, April 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-04-1997/jw-04-jdk.htm.

Murphy:1997:IAG

[Mur97b]  Kieron Murphy.  IBM’s

AppletAuthor grinds JavaBeans. JavaWorld: IDG’s magazine for the Java community, 2(1):??, January 1997. CODEN ???? ISSN 1091-8906. URL http://

                                        www.javaworld.com/javaworld/
                                        jw-01-1997/jw-01-appauthor.
                                        htm.

Murphy:1997:JUP

[Mur97c] Kieron Murphy. Java undergoes purification ritual at Internet World. Java-World: IDG’s magazine for the Java community, 2(1): ??, January 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-01-1997/jw-01-iworld2.
                                        htm.

Murphy:1997:NBJ

[Mur97d] Kieron Murphy. News: Beta 3 of JDK 1.1 is chock fulla Beans — and more. Java-World: IDG’s magazine for the Java community, 2(2): ??, February 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-02-1997/jw-02-jdk1.
                                        1b3.htm.

Murphy:1997:NVJ

[Mur97e] Kieron Murphy. News: Visual J++ 1.01 runs into an installation snag. Java-World: IDG’s magazine for the Java community, 2(2): ??, February 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-02-1997/jw-02-visual.
java.htm.

Murphy:1997:WJA

[Mur97f] Kieron Murphy. Why Java-Soft already released another JDK 1.1 beta. Java-World: IDG’s magazine for the Java community, 2(1): ??, January 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1997/jw-01-jdkbug.
htm.

Murphy:1998:FBB

[Mur98]  Kieron Murphy.  First

bug bites Java 1.2 — Mozilla stung. Java-World: IDG’s magazine for the Java community, 3(8): ??, August 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-08-1998/jw-08-12bug.
htm.

Neffenger:1997:JEW

[Nef97]  John Neffenger.  Java-

World exclusive: We reveal results of first-ever JVM server benchmark. Java-World: IDG’s magazine for the Java community, 2(12): ??, December 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.

                                        javaworld.com/javaworld/
                                        jw-12-1997/jw-12-volanomark.
                                        htm.

Neffenger:1998:WJV

[Nef98] John Neffenger. Which Java VM scales best? Java-World: IDG’s magazine for the Java community, 3(8): ??, August 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-08-1998/jw-08-volanomark.
                                        htm.

Nelson:1997:NCS

[Nel97a]  Matthew Nelson.  News:

Cisco to secure Java. Java-World: IDG’s magazine for the Java community, 2(8): ??, August 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-08-1997/jw-08-infoworld.
                                        cisco.htm.

Nelson:1997:NSZ

[Nel97b]  Matthew Nelson.  News:

Sun’s Zander reiterates importance of networks and Java. JavaWorld: IDG’s magazine for the Java community, 2(10):??, October 1997. CODEN ???? ISSN 1091-8906. URL http://

                                        www.javaworld.com/javaworld/
                                        jw-10-1997/jw-10-infoworld.
                                        zander.htm.

Nelson:1998:DCF

[Nel98] Matthew Nelson. Developer creates the first Java virus and names it ‘Strange Brew’. JavaWorld: IDG’s magazine for the Java community, 3 (9):??, September 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1998/jw-09-iw-virus.
htm.

Niccolai:1998:ASS

[Nic98a] James Niccolai. Analyst:

Sun scraps Java chip plans. JavaWorld: IDG’s magazine for the Java community, 3(11):??, 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-11-1998/jw-11-idgns-
javachip.htm.

Niccolai:1998:FMT

[Nic98b] James Niccolai. Focus at Microsoft trial switches to Java. JavaWorld: IDG’s magazine for the Java community, 3(12):??, 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-12-1998/jw-12-idgns-
gosling.htm.

Niccolai:1998:JSD

[Nic98c] James Niccolai. Java Software Division president feels heat from Microsoft lawyers in Java case. JavaWorld: IDG’s magazine for the Java community, 3(10):??, 1998. CODEN ???? ISSN 1091-

8906. URL http://www.

                                        javaworld.com/javaworld/
                                        jw-10-1998/jw-10-idgns-
                                        baratz.htm.

Niccolai:1998:JMR

[Nic98d] James Niccolai. Judge may rule soon in Sun–Microsoft hearing. JavaWorld: IDG’s magazine for the Java community, 3(12):??, 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-12-1998/jw-12-iw-ruling.
                                        htm.

Niccolai:1998:MET

[Nic98e] James Niccolai. Microsoft executive testifies in Sun’s lawsuit. JavaWorld: IDG’s magazine for the Java community, 3(10):??, 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-10-1998/jw-10-idgns-
                                        muglia.htm.

Niccolai:1998:NSS

[Nic98f] James Niccolai. New Sun software lets older PCs run Java. JavaWorld: IDG’s magazine for the Java community, 3(7):??, July 1998. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-07-1998/jw-07-idgns-
                                        oldpcs.htm.

Niccolai:1998:NMT

[Nic98g] James Niccolai. News: Mc-Nealy touts Java everywhere, TCI deal. Java-World: IDG’s magazine for the Java community, 3(2): ??, February 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-02-1998/jw-02-idgns.
mcnealy.htm.

Niccolai:1998:SMF

[Nic98h] James Niccolai. Sun and Microsoft fire closing volleys in Java hearing. JavaWorld: IDG’s magazine for the Java community, 3(10):??, 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1998/jw-10-idgns-
hearing.htm.

Niccolai:1998:SMO

[Nic98i] James Niccolai. Sun and Microsoft offer judge differing lessons on Java. Java-World: IDG’s magazine for the Java community, 3(9): ??, September 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1998/jw-09-idgns-
court.htm.

Niccolai:1998:SJV

[Nic98j] James Niccolai. Sun’s Java VM no good for embedded market, HP says. Java-World: IDG’s magazine for the Java community, 3 (5):??, May 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-05-1998/jw-05-idgns.
                                        hp.htm.

Niccolai:1998:SAE

[Nic98k] James Niccolai. Symantec aims for enterprise with Java development tool. Java-World: IDG’s magazine for the Java community, 3(8): ??, August 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-08-1998/jw-08-idgns-
                                        symantec.htm.

Niccolai:1998:ZDC

[Nic98l] James Niccolai. Zealous developers confident of Java’s future. JavaWorld: IDG’s magazine for the Java community, 3(4):??, April 1998. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-04-1998/jw-04-idgns.
                                        zealous.htm.

Niccolai:1999:MBC

[Nic99] James Niccolai. Microsoft begins to comply with Java order. JavaWorld: IDG’s magazine for the Java community, 4(1):??, January 1999. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-01-1999/jw-01-idgns-
                                        mscomplies.htm.

OConnell:1998:DTS

[OA98] Michael O’Connell and Jenni Aloi. Developers tell Sun where to put Swing for JDK 1.2. JavaWorld: IDG’s magazine for the Java community, 3(9):??, September 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1998/jw-09-pollresults.
htm.

OConnell:1996:ENJ

[O’C96a] Michael O’Connell. Editor’s note: Java and the Macintosh. JavaWorld: IDG’s magazine for the Java community, 1(3):??, May 1996. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-05-1996/jw-05-editorial.
htm.

OConnell:1996:NB

[O’C96b] Michael O’Connell. News briefs. JavaWorld: IDG’s magazine for the Java community, 1(1):??, March 1996. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-03-1996/jw-03-newsbriefs.
htm.

OConnell:1996:NJU

[O’C96c] Michael O’Connell. News:

JavaSoft unveils Developer Connection program. Java-World: IDG’s magazine for the Java community, 1(9):

??, November 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1996/jw-10-devconnect.
htm.

OConnell:1997:JCO

[O’C97] Michael O’Connell. From the editor: JavaWorld celebrates its one-year anniversary! JavaWorld: IDG’s magazine for the Java community, 2(3):??, March 1997. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-03-1997/jw-03-editnote.
htm.

OConnell:1998:DET

[O’C98]  Michael O’Connell.  De-

velopers explain timing of their transition to JDK 1.2. JavaWorld: IDG’s magazine for the Java community, 3(11):??, 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-11-1998/jw-11-pollresults.
htm.

Orchard:1997:NLJ

[Orc97] Dave Orchard. News: A look at Java’s future. Java-World: IDG’s magazine for the Java community, 2 (6):??, June 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-06-1997/jw-06-javafuture.
htm.

OConnell:1998:JLP

[OS98] Michael O’Connell and Kane Scarlett. Java Lobby pushes PC makers to include standard Java. Java-World: IDG’s magazine for the Java community, 3(8): ??, August 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-08-1998/jw-08-javareadypc.
htm.

Owens:1998:TJC

[Owe98] Ross Owens. Thin Java coverage in a fat Java book. JavaWorld: IDG’s magazine for the Java community, 3(5):??, May 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-05-1998/jw-05-iw.bookrev.
htm.

Parr:1997:OWW

[Par97] Terence Parr. Opinion: Why we care about Java. Java-World: IDG’s magazine for the Java community, 2(11): ??, November 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-11-1997/jw-11-portability.
htm.

Perez:1998:NGS

[Pér98] Juan Carlos Pérez. News:

Gates should have finished school, McNealy says. Java-World: IDG’s magazine for the Java community, 3 (7):??, July 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1998/jw-07-idgns-
school.htm.

Petreley:1997:NHM

[Pet97a] Nicholas Petreley. News: Horse’s mouth: Corel clarifies its short- and longterm Java strategies. Java-World: IDG’s magazine for the Java community, 2(11):??, November 1997. CODEN ???? ISSN 1091-8906. URL http://

www.ncworldmag.com/ncw-
10-1997/ncw-10-remalta.
htm.

Petreley:1997:PRB

[Pet97b] Nicholas Petreley. Product review: Building platformneutral Java: Which tool is best? JavaWorld: IDG’s magazine for the Java community, 2(11):??, November 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-11-1997/jw-11-jvtools.
htm.

Petreley:1998:PAJ

[Pet98] Nicholas Petreley. Product analysis: Java Studio 1.0: We want it! (but not right now). Java-World: IDG’s magazine for the Java community, 3(1): ??, January 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1998/jw-01-javastudio.
htm.

Philion:1998:BSB

[Phi98] Paul Philion. Build servletbased enterprise Web applications. JavaWorld: IDG’s magazine for the Java community, 3(12):??, 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-12-1998/jw-12-servlethtml.
htm.

Pietrowicz:1997:BRJ

[Pie97a] Stephen Pietrowicz. Book review: Java for business. JavaWorld: IDG’s magazine for the Java community, 2 (10):??, October 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1997/jw-10-bookreview.
htm.

Pietrowicz:1997:BRL

[Pie97b] Stephen Pietrowicz. Book review: Late Night VRML 2.0 with Java. Java-World: IDG’s magazine for the Java community, 2(11): ??, November 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-11-1997/jw-11-bookreview-
vrml.htm.

Porebski:1997:JTT

[Por97] Andrzej Porebski. Java tips: Tip 31: How to create mutually exclusive menus in Java. JavaWorld: IDG’s magazine for the Java community, 2(6):??, June 1997. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        javatips/jw-javatip31.
                                        htm.

Radosevich:1997:NMM

[Rad97a] Lynda Radosevich. News: Marimba moves beyond Java to update apps. Java-World: IDG’s magazine for the Java community, 2(9): ??, September 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-09-1997/jw-09-infoworld-
                                        marimba.htm.

Radosevich:1997:NMC

[Rad97b] Lynda Radosevich. News: Microsoft conjures Javabased software distribution. JavaWorld: IDG’s magazine for the Java community, 2(7):??, July 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-07-1997/jw-07-infoworld.
                                        package.htm.

Radosevich:1997:NND

[RB97] Lynda Radosevich and Jim Balderston. News: Netscape DevCon: Andreessen pushes JavaBean components. Java-World: IDG’s magazine for the Java community, 2 (7):??, July 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1997/jw-07-infoworld.
devcon-beans.htm.

Retelewski:1997:JTC

[Ret97]  Tom Retelewski.  Java

tips: Create new event types in Java. Java-World: IDG’s magazine for the Java community, 2(8): ??, August 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javatips/jw-javatip35.
htm.

Rex:1997:JPV

[Rex97a] Erica Rex. JavaOne: The participants’ view. Java-World: IDG’s magazine for the Java community, 2 (5):??, May 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-05-1997/jw-05-jo-quotes.
htm.

Rex:1997:K

[Rex97b] Erica Rex. Keynotes. Java-World: IDG’s magazine for the Java community, 2 (5):??, May 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/

                                        javaone97/java1-97-toc.
                                        html#keynote.

Ripps:1996:JES

[Rip96] David L. Ripps. Java for embedded systems. Java-World: IDG’s magazine for the Java community, 1(7): ??, September 1996. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-09-1996/jw-09-ripps.
                                        htm.

Rex:1997:IJI

[RO97] Erica Rex and Dave Orchard. Interviews with Java insiders. JavaWorld: IDG’s magazine for the Java community, 2(5):??, May 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        javaone97/java1-97-toc.
                                        html#intervie.

Rodriguez:1998:NTD

[Rod98] Karen Rodriguez. News: TCI deal breaks new ground for Java developers. Java-World: IDG’s magazine for the Java community, 3(2): ??, February 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-02-1998/jw-02-tci.htm.

Rommel:1998:ILM

[Rom98]  John Rommel.  An in-

side look at Microsoft’s recent Visual J++ conference.

JavaWorld: IDG’s magazine for the Java community, 3(7):??, July 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1998/jw-07-vjconf.
htm.

Rosenthal:1997:NAP

[Ros97]  Steve Rosenthal.  News:

Apple promises even closer JavaOS ties. JavaWorld: IDG’s magazine for the Java community, 2(7):??, July 1997. CODEN ???? ISSN 1091-8906. URL http://

www.javaworld.com/javaworld/
jw-07-1997/jw-07-apple.
htm.

Roulo:1997:JTT

[Rou97a] Mark Roulo. Java’s three types of portability. Java-World: IDG’s magazine for the Java community, 2 (5):??, May 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-05-1997/jw-05-portability.
htm.

Roulo:1997:RLT

[Rou97b] Mark Roulo. Reduce the launch time of your applets: Store them on client machines. JavaWorld: IDG’s magazine for the Java community, 2(6):??, June 1997. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/

                                        jw-06-1997/jw-06-plugins.
                                        htm.

Roulo:1998:AYJ

[Rou98]  Mark Roulo.  Accelerate

your Java apps! Java-World: IDG’s magazine for the Java community, 3(9): ??, September 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-09-1998/jw-09-speed.
                                        htm.

Roulo:1999:HAT

[Rou99] Mark Roulo. How to avoid traps and correctly override methods from java.lang.object. JavaWorld: IDG’s magazine for the Java community, 4 (1):??, January 1999. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-01-1999/jw-01-object.
                                        htm.

Radosevich:1997:NIE

[RS97] Lynda Radosevich and Rebecca Sykes. News: Internet Expo: Lotus’s Papows offers praise for Microsoft and its approach with Java. Java-World: IDG’s magazine for the Java community, 2(9): ??, September 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-09-1997/jw-09-idgns-
                                        praise.htm.

Ryan:1996:BRB

[Rya96a] S. Ryan. Book review: Book on the Java API is a musthave for Java programmers. JavaWorld: IDG’s magazine for the Java community, 1 (6):??, August 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-08-1996/jw-08-bookrev-
api.htm.

Ryan:1996:BRO

[Rya96b] S. Ryan. Book review: The only Java book you’ll ever need? JavaWorld: IDG’s magazine for the Java community, 1(5):??, July 1996. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1996/jw-07-bookrev.
htm.

Ryan:1996:JDGa

[Rya96c] S. Ryan. JavaOne delivers gospel. JavaWorld: IDG’s magazine for the Java community, 1(4):??, June 1996. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-06-1996/jw-06-javaone-
ryan.htm.

Ryan:1996:JDGb

[Rya96d] S. Ryan. JavaOne delivers gospel. JavaWorld: IDG’s magazine for the Java community, 1(5):??, July 1996. CODEN ???? ISSN 1091-

8906. URL http://www.

                                        javaworld.com/javaworld/
                                        jw-07-1996/jw-07-javaone-
                                        ryan.htm.

Saljoughy:1997:OPJ

[Sal97]  Arsalan Saljoughy.  Ob-

ject persistence and Java. JavaWorld: IDG’s magazine for the Java community, 2(5):??, May 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-05-1997/jw-05-persistence.
                                        htm.

Sandvik:1996:CEJ

[San96] Kent Sandvik. Card engine in Java. Java-World: IDG’s magazine for the Java community, 1(6): ??, August 1996. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-08-1996/jw-08-cards.
                                        htm.

Sands:1998:NJD

[San98] John Sands. News: JFC: An in-depth look at Sun’s successor to AWT. Java-World: IDG’s magazine for the Java community, 3(1): ??, January 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-01-1998/jw-01-jfc.htm.

Sauvageau:1998:API

[Sau98a]  Yvon Sauvageau.  Ap-

plet parameterization — initializing arrays. Java-World: IDG’s magazine for the Java community, 3(9): ??, September 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javatips/jw-javatip59.
htm.

Sauvageau:1998:APC

[Sau98b]  Yvon Sauvageau.  Ap-

plet parameterization via class reflection. Java-World: IDG’s magazine for the Java community, 3(8): ??, August 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javatips/jw-javatip57.
htm.

Scannell:1997:NIRb

[SB97] Ed Scannell and Ted Smalley Bowen. News: IBM readies VisualAge for Java. JavaWorld: IDG’s magazine for the Java community, 2(6):??, June 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-06-1997/jw-06-infoworld.
ibm.htm.

Scannell:1997:JCS

[SC97] Ed Scannell and Cara Cunningham. JavaSoft cooking up some next-generation JavaBeans. JavaWorld: IDG’s magazine for the Java community, 2(5):??, May 1997. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-05-1997/jw-05-infoworld.
                                        nextgenbeans.htm.

Scarlett:1996:NN

[Sca96a] Kane Scarlett. News and new product briefs (10/1/96). JavaWorld: IDG’s magazine for the Java community, 1 (9):??, November 1996. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-10-1996/jw-10-newsbriefs2.
                                        htm.

Scarlett:1996:NNP

[Sca96b] Kane Scarlett. News and new product briefs (9/15/96). JavaWorld: IDG’s magazine for the Java community, 1 (9):??, November 1996. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-10-1996/jw-10-newsbriefs.
                                        htm.

Scannell:1997:NGF

[Sca97a] Ed Scannell. News: ‘Gang of Four’ to extend CORBA to support JavaBeans. Java-World: IDG’s magazine for the Java community, 2 (7):??, July 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-07-1997/jw-07-infoworld.
                                        corba.htm.

Scannell:1997:NIL

[Sca97b] Ed Scannell. News: IBM launches beta of Java-based Host On-Demand. Java-World: IDG’s magazine for the Java community, 2(9): ??, September 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1997/jw-09-infoworld-
ibm.htm.

Scarlett:1997:NNPc

[Sca97c] Kane Scarlett. News and new product briefs — 25 new articles! (3/1/97). JavaWorld: IDG’s magazine for the Java community, 2 (3):??, March 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-03-1997/jw-03-newsbriefs2.
htm.

Scarlett:1997:NNPp

[Sca97d] Kane Scarlett. News and new product briefs (10/01/97). JavaWorld: IDG’s magazine for the Java community, 2 (10):??, October 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1997/jw-10-newsbriefs2.
htm.

Scarlett:1997:NNPs

[Sca97e] Kane Scarlett. News and new product briefs (10/15/97). JavaWorld: IDG’s magazine for the Java community, 2 (11):??, November 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-11-1997/jw-11-newsbriefs.
htm.

Scarlett:1997:NNPr

[Sca97f] Kane Scarlett. News and new product briefs (11/01/97). JavaWorld: IDG’s magazine for the Java community, 2 (11):??, November 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-11-1997/jw-11-newsbriefs2.
htm.

Scarlett:1997:NNPu

[Sca97g] Kane Scarlett. News and new product briefs (11/15/97). JavaWorld: IDG’s magazine for the Java community, 2 (12):??, December 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-12-1997/jw-12-newsbriefs.
htm.

Scarlett:1997:NNPt

[Sca97h] Kane Scarlett. News and new product briefs (12/2/97). JavaWorld: IDG’s magazine for the Java community, 2 (12):??, December 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-12-1997/jw-12-newsbriefs2.
htm.

Scarlett:1997:NNPd

[Sca97i] Kane Scarlett. News and new product briefs (2/15/97). JavaWorld: IDG’s magazine for the Java community, 2 (3):??, March 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-03-1997/jw-03-newsbriefs.
htm.

Scarlett:1997:NNPo

[Sca97j] Kane Scarlett. News and new product briefs (8/15/97). JavaWorld: IDG’s magazine for the Java community, 2 (9):??, September 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1997/jw-09-newsbriefs.
htm.

Scarlett:1997:NNPq

[Sca97k] Kane Scarlett. News and new product briefs (9/15/97). JavaWorld: IDG’s magazine for the Java community, 2 (10):??, October 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1997/jw-10-newsbriefs.
htm.

Scarlett:1997:NNPn

[Sca97l] Kane Scarlett. News and new product briefs (9/3/97). JavaWorld: IDG’s magazine for the Java community, 2 (9):??, September 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1997/jw-09-newsbriefs2.
htm.

Scarlett:1997:NNNd

[Sca97m]  Kane Scarlett.  News:

News and new product briefs (1/15/97). Java-World: IDG’s magazine for the Java community, 2(2): ??, February 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-02-1997/jw-02-newsbriefs.
htm.

Scarlett:1997:NNNa

[Sca97n]  Kane Scarlett.  News:

News and new product briefs (1/1/97). Java-World: IDG’s magazine for the Java community, 2(1): ??, January 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1997/jw-01-newsbriefs2.
htm.

Scarlett:1997:NNNb

[Sca97o]  Kane Scarlett.  News:

News and new product briefs (12/15/96). Java-World: IDG’s magazine for the Java community, 2(1): ??, January 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1997/jw-01-newsbriefs.
htm.

Scarlett:1997:NNNc

[Sca97p]  Kane Scarlett.  News:

News and new product briefs (2/1/97). Java-World: IDG’s magazine for the Java community, 2(2): ??, February 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-02-1997/jw-02-newsbriefs2.
htm.

Scarlett:1997:NNNf

[Sca97q]  Kane Scarlett.  News:

News and new product briefs (3/15/97). Java-World: IDG’s magazine for the Java community, 2 (4):??, April 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-04-1997/jw-04-newsbriefs.
htm.

Scarlett:1997:NNNh

[Sca97r] Kane Scarlett. News: News and new product briefs (4/15/97) — 25 stories! JavaWorld: IDG’s magazine for the Java community, 2(5):??, May 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-05-1997/jw-05-newsbriefs.
htm.

Scarlett:1997:NNNe

[Sca97s] Kane Scarlett. News: News and new product briefs (4/1/97). JavaWorld: IDG’s magazine for the Java community, 2(4):??, April 1997. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-04-1997/jw-04-newsbriefs2.
htm.

Scarlett:1997:NNNg

[Sca97t] Kane Scarlett. News: News and new product briefs (5/1/97) — 33 new stories! JavaWorld: IDG’s magazine for the Java community, 2(5):??, May 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-05-1997/jw-05-newsbriefs2.
htm.

Scarlett:1997:NNNk

[Sca97u]  Kane Scarlett.  News:

News and new product briefs (6/15/97). Java-World: IDG’s magazine for the Java community, 2 (7):??, July 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1997/jw-07-newsbriefs.
htm.

Scarlett:1997:NNNi

[Sca97v] Kane Scarlett. News: News and new product briefs (6/1/97). JavaWorld: IDG’s magazine for the Java community, 2(6):??, June 1997. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/

jw-06-1997/jw-06-newsbriefs2.
htm.

Scarlett:1997:NNNm

[Sca97w]  Kane Scarlett.  News:

News and new product briefs (7/15/97). Java-World: IDG’s magazine for the Java community, 2(8): ??, August 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-08-1997/jw-08-newsbriefs.
htm.

Scarlett:1997:NNNj

[Sca97x] Kane Scarlett. News: News and new product briefs (7/1/97). JavaWorld: IDG’s magazine for the Java community, 2(7):??, July 1997. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1997/jw-07-newsbriefs2.
htm.

Scarlett:1997:NNNl

[Sca97y]  Kane Scarlett.  News:

News and new product briefs (8/4/97). Java-World: IDG’s magazine for the Java community, 2(8): ??, August 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-08-1997/jw-08-newsbriefs2.
htm.

Scannell:1998:IEJ

[Sca98a] Ed Scannell. IBM eyes Java marketplace with AS/400.

JavaWorld: IDG’s magazine for the Java community, 3(11):??, 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-11-1998/jw-11-iw-as400.
htm.

Scannell:1998:NIW

[Sca98b] Ed Scannell. News: Internet world: Java’s real role. JavaWorld: IDG’s magazine for the Java community, 3 (1):??, January 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1998/jw-01-infoworld.
javarole.htm.

Scarlett:1998:NNPc

[Sca98c] Kane Scarlett. News and new product briefs. Java-World: IDG’s magazine for the Java community, 3(1): ??, January 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1998/jw-01-newsbriefs2.
htm.

Scarlett:1998:NNPd

[Sca98d] Kane Scarlett. News and new product briefs. Java-World: IDG’s magazine for the Java community, 3(1): ??, January 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1998/jw-01-newsbriefs.
htm.

Scarlett:1998:NNPe

[Sca98e] Kane Scarlett. News and new product briefs. Java-World: IDG’s magazine for the Java community, 3(2): ??, February 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-02-1998/jw-02-newsbriefs2.
htm.

Scarlett:1998:NNPf

[Sca98f] Kane Scarlett. News and new product briefs. Java-World: IDG’s magazine for the Java community, 3(2): ??, February 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-02-1998/jw-02-newsbriefs.
htm.

Scarlett:1998:NNPg

[Sca98g] Kane Scarlett. News and new product briefs. Java-World: IDG’s magazine for the Java community, 3 (3):??, March 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-03-1998/jw-03-newsbriefs2.
htm.

Scarlett:1998:NNPh

[Sca98h] Kane Scarlett. News and new product briefs. Java-World: IDG’s magazine for the Java community, 3 (3):??, March 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-03-1998/jw-03-newsbriefs.
                                        htm.

Scarlett:1998:NNPj

[Sca98i] Kane Scarlett. News and new product briefs. Java-World: IDG’s magazine for the Java community, 3 (4):??, April 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-04-1998/jw-04-newsbriefs.
                                        htm.

Scarlett:1998:NNPw

[Sca98j] Kane Scarlett. News and new product briefs (10/15/98). JavaWorld: IDG’s magazine for the Java community, 3(11):??, 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-11-1998/jw-11-newsbriefs.
                                        htm.

Scarlett:1998:NNPt

[Sca98k] Kane Scarlett. News and new product briefs (10/1/98). JavaWorld: IDG’s magazine for the Java community, 3(10):??, 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-10-1998/jw-10-newsbriefs2.
                                        htm.

Scarlett:1998:NNPy

[Sca98l] Kane Scarlett. News and new product briefs (11/15/98).

JavaWorld: IDG’s magazine for the Java community, 3(12):??, 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-12-1998/jw-12-newsbriefs.
htm.

Scarlett:1998:NNPv

[Sca98m] Kane Scarlett. News and new product briefs (11/2/98). JavaWorld: IDG’s magazine for the Java community, 3(11):??, 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-11-1998/jw-11-newsbriefs2.
htm.

Scarlett:1998:NNPx

[Sca98n] Kane Scarlett. News and new product briefs (12/2/98). JavaWorld: IDG’s magazine for the Java community, 3(12):??, 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-12-1998/jw-12-newsbriefs2.
htm.

Scarlett:1998:NNPl

[Sca98o] Kane Scarlett. News and new product briefs (4/15/98). JavaWorld: IDG’s magazine for the Java community, 3(5):??, May 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-05-1998/jw-05-newsbriefs.
htm.

Scarlett:1998:NNPi

[Sca98p] Kane Scarlett. News and new product briefs (4/1/98). JavaWorld: IDG’s magazine for the Java community, 3 (4):??, April 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-04-1998/jw-04-newsbriefs2.
                                        htm.

Scarlett:1998:NNPk

[Sca98q] Kane Scarlett. News and new product briefs (5/4/98). JavaWorld: IDG’s magazine for the Java community, 3(5):??, May 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-05-1998/jw-05-newsbriefs2.
                                        htm.

Scarlett:1998:NNPq

[Sca98r] Kane Scarlett. News and new product briefs (7/15/98). JavaWorld: IDG’s magazine for the Java community, 3 (8):??, August 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-08-1998/jw-08-newsbriefs.
                                        htm.

Scarlett:1998:NNPo

[Sca98s] Kane Scarlett. News and new product briefs (7/2/98). JavaWorld: IDG’s magazine for the Java community, 3(7):??, July 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1998/jw-07-newsbriefs2.
htm.

Scarlett:1998:NNPs

[Sca98t] Kane Scarlett. News and new product briefs (8/15/98). JavaWorld: IDG’s magazine for the Java community, 3 (9):??, September 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1998/jw-09-newsbriefs.
htm.

Scarlett:1998:NNPp

[Sca98u] Kane Scarlett. News and new product briefs (8/3/98). JavaWorld: IDG’s magazine for the Java community, 3 (8):??, August 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-08-1998/jw-08-newsbriefs2.
htm.

Scarlett:1998:NNPu

[Sca98v] Kane Scarlett. News and new product briefs (9/15/98). JavaWorld: IDG’s magazine for the Java community, 3(10):??, 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1998/jw-10-newsbriefs.
htm.

Scarlett:1998:NNPr

[Sca98w] Kane Scarlett. News and new product briefs (9/1/98).

JavaWorld: IDG’s magazine for the Java community, 3 (9):??, September 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1998/jw-09-newsbriefs2.
htm.

Scarlett:1998:NGR

[Sca98x] Kane Scarlett. News: Get ready to Swing (1.0). Java-World: IDG’s magazine for the Java community, 3 (3):??, March 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-03-1998/jw-03-swinggui.
htm.

Scarlett:1998:NNNd

[Sca98y]  Kane Scarlett.  News:

News and new product briefs (5/15/98). Java-World: IDG’s magazine for the Java community, 3 (6):??, June 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-06-1998/jw-06-newsbriefs.
htm.

Scarlett:1998:NNNe

[Sca98z]  Kane Scarlett.  News:

News and new product briefs (6/15/98). Java-World: IDG’s magazine for the Java community, 3 (7):??, July 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/

jw-07-1998/jw-07-newsbriefs.
htm.

Scarlett:1998:NNNc

[Sca98-27] Kane Scarlett. News: News and new product briefs (6/2/98). JavaWorld: IDG’s magazine for the Java community, 3(6):??, June 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-06-1998/jw-06-newsbriefs2.
htm.

Scarlett:1998:NPI

[Sca98-28] Kane Scarlett. News profile: ILOG JViews: Backed by programming experience. JavaWorld: IDG’s magazine for the Java community, 3 (4):??, April 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-04-1998/jw-04-ilog.
htm.

Scarlett:1998:NRT

[Sca98-29]  Kane Scarlett.  News:

Roaster Technologies suddenly decides to end it all. JavaWorld: IDG’s magazine for the Java community, 3 (1):??, January 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1998/jw-01-roaster.
htm.

Scarlett:1999:NNP

[Sca99] Kane Scarlett. News and new product briefs (12/18/98).

JavaWorld: IDG’s magazine for the Java community, 4 (1):??, January 1999. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1999/jw-01-newsbriefs.
htm.

Schwartz:1996:SSB

[Sch96a] Rebecca Schwartz. Stock-Watch: Strange brew: Branding, premiums, and America Online. Java-World: IDG’s magazine for the Java community, 1(9): ??, November 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1996/jw-10-stockwatch.
htm.

Schwartz:1996:SNM

[Sch96b] Rebecca Schwartz. Stock-Watch: The next Microsoft. JavaWorld: IDG’s magazine for the Java community, 1(5):??, July 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1996/jw-07-stockwatch.
htm.

Schwartz:1996:SWI

[Sch96c] Rebecca Schwartz. Stock-Watch: When ‘intel inside’ no longer matters. Java-World: IDG’s magazine for the Java community, 1(6): ??, August 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.

javaworld.com/javaworld/
jw-08-1996/jw-08-stockwatch.
htm.

Schwartz:1997:SHI

[Sch97a] Rebecca Schwartz. Stock-Watch: How IBM’s January effect will impact its stock’s value. Java-World: IDG’s magazine for the Java community, 2(2): ??, February 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-02-1997/jw-02-stockwatch.
htm.

Schwartz:1997:SCK

[Sch97b] Rebecca Schwartz. Stock-Watch: The Cisco kid — friend of mine? Java-World: IDG’s magazine for the Java community, 2 (4):??, April 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-04-1997/jw-04-stockwatch.
htm.

Scherer:1998:RCM

[SD98] Jason Scherer and Rinaldo Di Giorgio. Restricted-channel multicast in Java. Java-World: IDG’s magazine for the Java community, 3(9): ??, September 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1998/jw-09-javadev.
htm.

Sykes:1998:SIP

[SE98] Rebecca Sykes and Kristi Essick. Sun, IBM pitch JavaOS for NCs as PC complement. JavaWorld: IDG’s magazine for the Java community, 3(5):??, May 1998. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-05-1998/jw-05-idgns.
                                        javaos.htm.

Shah:1996:YBS

[Sha96a]  Rawn Shah.  Are you

being served? Java-World: IDG’s magazine for the Java community, 1(9): ??, November 1996. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-10-1996/jw-10-jeeves.
                                        htm.

Shah:1996:ATS

[Sha96b] Rawn Shah. Authoring tools survey. JavaWorld: IDG’s magazine for the Java community, 1(5):??, July 1996. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-07-1996/jw-07-authortools.
                                        htm.

Shah:1996:BGJ

[Sha96c]  Rawn Shah.  A begin-

ner’s guide to JavaScript. JavaWorld: IDG’s magazine for the Java community, 1 (1):??, March 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-03-1996/jw-03-javascript.
intro.htm.

Shah:1996:BBM

[Sha96d]  Rawn Shah.  Bending

over backwards to make JavaScript work on 14 platforms. JavaWorld: IDG’s magazine for the Java community, 1(2):??, April 1996. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-04-1996/jw-04-jsinterview.
htm.

Shah:1996:DOC

[Sha96e] Rawn Shah. Distributed object computing with Joe and NEO. JavaWorld: IDG’s magazine for the Java community, 1(4):??, June 1996. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-06-1996/jw-06-dist.
comp.htm.

Shah:1996:IDJ

[Sha96f]  Rawn Shah.  Integrating

databases with Java via JDBC. JavaWorld: IDG’s magazine for the Java community, 1(3):??, May 1996. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-05-1996/jw-05-shah.
htm.

Shah:1996:JNF

[Sha96g]  Rawn Shah.  JavaScript

and Netscape frames. Java-World: IDG’s magazine for the Java community, 1 (2):??, April 1996. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-04-1996/jw-04-javascript.
                                        htm.

Shah:1996:PHB

[Sha96h] Rawn Shah. A portable hill of beans. JavaWorld: IDG’s magazine for the Java community, 1(7):??, September 1996. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-09-1996/jw-09-javabeans.
                                        htm.

Shah:1997:HJW

[Sha97]  Rawn Shah.  How Java

will bring together the Internet and telephony. Java-World: IDG’s magazine for the Java community, 2(2): ??, February 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-02-1997/jw-02-api.telephony.
                                        htm.

Shah:1998:SJ

[Sha98a] Rawn Shah. The skinny on Jini. JavaWorld: IDG’s magazine for the Java community, 3(8):??, August 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-08-1998/jw-08-jini.
htm.

Shah:1998:SIJ

[Sha98b] Rawn Shah. Sun, IBM join to announce JavaOS for Business. JavaWorld: IDG’s magazine for the Java community, 3(4):??, April 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-04-1998/jw-04-javaos.
htm.

Shoffner:1997:JSSd

[Sho97a] Michael Shoffner. Java step by step: Increase the functionality in your distributed client/server apps. Java-World: IDG’s magazine for the Java community, 2(1): ??, January 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1997/jw-10-step.
html.

Shoffner:1997:JSSc

[Sho97b] Michael Shoffner. Java step by step: Scale an application from two to three tiers with JDBC. JavaWorld: IDG’s magazine for the Java community, 2(6):??, June 1997. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-06-1997/jw-06-step.
htm.

Shoffner:1997:JSSa

[Sho97c] Michael Shoffner. Java step by step: Write your own threaded discussion forum. JavaWorld: IDG’s magazine for the Java community, 2 (2):??, February 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-02-1997/jw-02-step.
                                        htm.

Shoffner:1997:JSSb

[Sho97d] Michael Shoffner. Java step by step: Write your own threaded discussion forum: The communications and server components. Java-World: IDG’s magazine for the Java community, 2 (3):??, March 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-03-1997/jw-03-step.
                                        htm.

Shoffner:1997:JVA

[Sho97e] Michael Shoffner. JavaBeans vs. ActiveX: Strategic analysis. JavaWorld: IDG’s magazine for the Java community, 2(2):??, February 1997. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-02-1997/jw-02-activex-
                                        beans.htm.

Shoffner:1998:JSSb

[Sho98a] Michael Shoffner. Java step by step: Add the power of CORBA to our distributed whiteboard. Java-World: IDG’s magazine for the Java community, 3 (3):??, March 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-03-1998/jw-03-step.
html.

Shoffner:1998:JSSa

[Sho98b] Michael Shoffner. Java step by step: Networking our whiteboard with servlets. JavaWorld: IDG’s magazine for the Java community, 3 (1):??, January 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1998/jw-01-step.
html.

Shoffner:1998:JSSd

[Sho98c]  Michael Shoffner.  Java

step by step: Write a session Enterprise JavaBean. JavaWorld: IDG’s magazine for the Java community, 3(7):??, July 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1998/jw-07-step.
htm.

Shoffner:1998:JSSc

[Sho98d] Michael Shoffner. Java step by step: Write your own MOM! JavaWorld: IDG’s magazine for the Java community, 3(5):??, May 1998. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-05-1998/jw-05-step.
                                        htm.

Shoffner:1998:MFSa

[Sho98e] Michael Shoffner. Making the forum swing. Java-World: IDG’s magazine for the Java community, 3(9): ??, September 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-09-1998/jw-09-step.
                                        htm.

Shoffner:1998:MFSb

[Sho98f] Michael Shoffner. Making the forum swing, Part 2. JavaWorld: IDG’s magazine for the Java community, 3(11):??, 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-11-1998/jw-11-step.
                                        htm.

Shoffner:1998:SRN

[Sho98g] Michael Shoffner. Sun releases the new Java 2 platform in style. JavaWorld: IDG’s magazine for the Java community, 3(12):??, 1998. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jbe/jw-jbe-launch.htm.

Shoffner:1999:ETJ

[Sho99a] Michael Shoffner. Enterprise technologies in the Java 2 platform. Java-World: IDG’s magazine for the Java community, 4(1): ??, January 1999. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jbe/jw-jbe-distributed.
htm.

Shoffner:1999:SRN

[Sho99b]  Michael Shoffner.  Sun

releases the new Java 2 platform in style. Java-World: IDG’s magazine for the Java community, 4(1): ??, January 1999. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jbe/jw-jbe-launch.htm.

Siddalingaiah:1997:NJI

[Sid97a] Madhu Siddalingaiah. News:

Java Internet Business Expo: A step forward for Java. JavaWorld: IDG’s magazine for the Java community, 2 (9):??, September 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1997/jw-09-jibe.
madhu.htm.

Siddalingaiah:1997:PFJ

[Sid97b] Madhu Siddalingaiah. Perspectives: The future of Java — rhetoric or reality? Java-World: IDG’s magazine for the Java community, 2(12): ??, December 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.

                                        javaworld.com/javaworld/
                                        jw-12-1997/jw-12-perspectives.
                                        htm.

Smith:1997:TT

[SKC97] Gregory S. Smith, Rich Kadel, and Phil Inje Chang. Technical topics. Java-World: IDG’s magazine for the Java community, 2 (5):??, May 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        javaone97/java1-97-toc.
                                        html#techtopic.

Scannell:1998:JJ

[SL98] Ed Scannell and Martin La-Monica. Java in jeopardy? JavaWorld: IDG’s magazine for the Java community, 3 (4):??, April 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-04-1998/jw-04-infoworld.
                                        javafocus.htm.

Steenburgen:1998:NAG

[SM98] Stephanie Steenburgen and Robert McMillan. News: Activator gets thumbs up from Java community. Java-World: IDG’s magazine for the Java community, 3(1): ??, January 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-01-1998/jw-01-activator.
                                        htm.

Small:1998:ESA

[Sma98] Steve Small. Escape the sandbox: Access native methods from an applet. JavaWorld: IDG’s magazine for the Java community, 3(10):??, 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1998/jw-10-apptowin32.
htm.

Smith:1997:JNV

[Smi97a] Gregory S. Smith. Java’s new virtual machine. Java-World: IDG’s magazine for the Java community, 2 (5):??, May 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-05-1997/jw-05-jo-vm.
htm.

Smith:1997:SF

[Smi97b] Gregory S. Smith. Sound and fury. JavaWorld: IDG’s magazine for the Java community, 2(5):??, May 1997. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-05-1997/jw-05-jo-media.
htm.

Smith:1998:ACM

[Smi98] Jerry Smith. Avoid ’constructor madness’. Java-World: IDG’s magazine for the Java community, 3(11): ??, 1998. CODEN ???? ISSN 1091-8906. URL http://

                                        www.javaworld.com/javaworld/
                                        javatips/jw-javatip63.
                                        htm.

Smith:1997:JMT

[SO97] Gregory S. Smith and Dave Orchard. JavaBeans make their debut at JavaOne. JavaWorld: IDG’s magazine for the Java community, 2(5):??, May 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-05-1997/jw-05-jo-javabeans.
                                        htm.

Sommers:1996:DAR

[Som96a] Bret Sommers. Distributing applets: Remote objects for Java. JavaWorld: IDG’s magazine for the Java community, 1(4):??, June 1996. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-06-1996/jw-06-remote.
                                        objects.htm.

Sommers:1996:NVCa

[Som96b] Bret Sommers. News and views from the conference. JavaWorld: IDG’s magazine for the Java community, 1(4):??, June 1996. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-06-1996/jw-06-javaone-
                                        sommers.htm.

Sommers:1996:NVCb

[Som96c] Bret Sommers. News and views from the conference.

JavaWorld: IDG’s magazine for the Java community, 1(5):??, July 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1996/jw-07-javaone-
sommers.htm.

Sommers:1997:AJB

[Som97a]  Bret Sommers.  Agents:

Not just for Bond anymore. JavaWorld: IDG’s magazine for the Java community, 2 (4):??, April 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-04-1997/jw-04-agents.
htm.

Sommers:1997:NND

[Som97b]  Bret Sommers.  News:

Netscape DevCon: Netscape coins ‘crossware’. Java-World: IDG’s magazine for the Java community, 2 (7):??, July 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1997/jw-07-devcon.
htm.

Sommers:1997:PPT

[Som97c]  Bret Sommers.  A pro-

grammer’s personal tour. JavaWorld: IDG’s magazine for the Java community, 2(5):??, May 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/

                                        javaone97/java1-97-toc.
                                        html#perspectiv.

Sommers:1998:NIS

[Som98] Bret Sommers. News: Internet Showcase features distinguished Java products. JavaWorld: IDG’s magazine for the Java community, 3 (3):??, March 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-03-1998/jw-03-ishowcase.
                                        htm.

Spencer:1998:NJI

[Spe98]  Gordon Spencer.  News:

JDK 1.2: Immature, but tons of new abilities. Java-World: IDG’s magazine for the Java community, 3(1): ??, January 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-01-1998/jw-01-jdkroadmap.
                                        htm.

Spruit:1997:RJB

[Spr97] Sandor Spruit. Reflections on Java, Beans, and relational databases. Java-World: IDG’s magazine for the Java community, 2(9): ??, September 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-09-1997/jw-09-reflections.
                                        htm.

Scannell:1997:NIRa

[SS97] Ed Scannell and Ephraim Schwartz. News: IBM rushes to build Java-enabled chip for NCs. JavaWorld: IDG’s magazine for the Java community, 2(6):??, June 1997. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-06-1997/jw-06-infoworld.
ibmchip.htm.

Steinberg:1996:NNPa

[Ste96a] Jill Steinberg. News and new product briefs (8/1/96). JavaWorld: IDG’s magazine for the Java community, 1 (6):??, August 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-08-1996/jw-08-newsbriefs2.
htm.

Steinberg:1996:NNPb

[Ste96b] Jill Steinberg. News and new product briefs (9/5/96). JavaWorld: IDG’s magazine for the Java community, 1 (7):??, September 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1996/jw-09-newsbriefs1.
htm.

Steinberg:1996:NJG

[Ste96c]  Jill Steinberg.  News:

Java gets $100 million endorsement. JavaWorld: IDG’s magazine for the Java community, 1(7):??, September 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1996/jw-09-javafund.
update.htm.

Steinberg:1996:NNS

[Ste96d] Jill Steinberg. News: New start-up releases Java application and enabling software. JavaWorld: IDG’s magazine for the Java community, 1 (9):??, November 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1996/jw-10-netit.
htm.

Steinberg:1996:NWR

[Ste96e]  Jill Steinberg.  News:

Wind River Systems licenses JavaOS, helping to make embedded development a breeze. Java-World: IDG’s magazine for the Java community, 1(9): ??, November 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1996/jw-10-windriver.
htm.

Steinberg:1996:TJW

[Ste96f]  Jill Steinberg.  Trends:

Java wins attention of Fortune 1,000. Java-World: IDG’s magazine for the Java community, 1(6): ??, August 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.

javaworld.com/javaworld/
jw-08-1996/jw-08-forrester.
htm.

Steinberg:1996:VFA

[Ste96g] Jill Steinberg. VC firm to announce $100 million ‘Java Fund’. Java-World: IDG’s magazine for the Java community, 1(6): ??, August 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-08-1996/jw-08-javafund.
htm.

Steinberg:1997:NJD

[Ste97a] Daniel H. Steinberg. News:

Java development on Macintosh — is it viable or not? JavaWorld: IDG’s magazine for the Java community, 2(7):??, July 1997. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1997/jw-07-javamac.
htm.

Steinberg:1997:NRJ

[Ste97b] Daniel H. Steinberg. News:

Roaster 4 — not just for Mac developers anymore. Java-World: IDG’s magazine for the Java community, 2(12): ??, December 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-12-1997/jw-12-roaster.
htm.

Steinberg:1997:NAM

[Ste97c] Daniel H. Steinberg. News: The Apple/Microsoft partnership will impact Java … but how? Java-World: IDG’s magazine for the Java community, 2(9): ??, September 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-09-1997/jw-09-partnership.
                                        htm.

Steinberg:1997:NMR

[Ste97d] Daniel H. Steinberg. News: The Macintosh Runtime for Java: What it is and where it’s heading. Java-World: IDG’s magazine for the Java community, 2(9): ??, September 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-09-1997/jw-09-mrj.htm.

Steinberg:1997:JHW

[Ste97e] Jill Steinberg. JavaWorld honors winners of Editors’ Choice Awards. Java-World: IDG’s magazine for the Java community, 2(10): ??, October 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-10-1997/jw-10-jwchoiceawards.
                                        htm.

Steinberg:1997:NCC

[Ste97f] Jill Steinberg. News: Competing components make for prickly panelists. Java-World: IDG’s magazine for the Java community, 2 (3):??, March 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-03-1997/jw-03-component.
web97.htm.

Steinberg:1997:NJS

[Ste97g] Jill Steinberg. News: Java Studio: For those who don’t want to write a line of code. JavaWorld: IDG’s magazine for the Java community, 2 (11):??, November 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-11-1997/jw-11-javastudio.
htm.

Steinberg:1997:NJP

[Ste97h] Jill Steinberg. News: Java-Soft president sums up where Java’s at. JavaWorld: IDG’s magazine for the Java community, 2(3):??, March 1997. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-03-1997/jw-03-baratz.
web97.htm.

Steinberg:1997:NPP

[Ste97i] Jill Steinberg. News: Polese pushes ‘push’ technology. JavaWorld: IDG’s magazine for the Java community, 2 (3):??, March 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/

                                        jw-03-1997/jw-03-polese.
                                        web97.htm.

Steelman:1998:JTP

[Ste98a] Jon Steelman. Java tip 50: On the Parameter Constants pattern. JavaWorld: IDG’s magazine for the Java community, 3(4):??, April 1998. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        javatips/jw-javatip50.
                                        htm.

Steinberg:1998:DJN

[Ste98b] Daniel H. Steinberg. Digging for Java news in Macworld Expo keynote. Java-World: IDG’s magazine for the Java community, 3(8): ??, August 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-08-1998/jw-08-mwkeynote.
                                        htm.

Steinberg:1998:JNS

[Ste98c] Daniel H. Steinberg. Java news from this summer’s Macworld Expo in New York. JavaWorld: IDG’s magazine for the Java community, 3 (8):??, August 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-08-1998/jw-08-mwexpo.
                                        htm.

Steinberg:1998:JB

[Ste98d] Daniel H. Steinberg. Java to be — or not to be — on the BeOS. JavaWorld: IDG’s magazine for the Java community, 3(7):??, July 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1998/jw-07-be.htm.

Steinberg:1998:NAR

[Ste98e] Daniel H. Steinberg. News:

Apple releases MacOS Runtime for Java (MRJ) version 2.0. JavaWorld: IDG’s magazine for the Java community, 3(1):??, January 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1998/jw-01-mrj2.
htm.

Steinberg:1998:NJM

[Ste98f] Daniel H. Steinberg. News:

Java on the MacOS — A report from Macworld Expo San Francisco. Java-World: IDG’s magazine for the Java community, 3(2): ??, February 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-02-1998/jw-02-macworld.
htm.

Steinberg:1998:JGE

[Ste98g] Jill Steinberg. Java goes to the extreme. Java-World: IDG’s magazine for the Java community, 3(8): ??, August 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.

                                        javaworld.com/javaworld/
                                        jw-08-1998/jw-08-srl.htm.

Steinberg:1998:NJF

[Ste98h] Jill Steinberg. News: JavaOS finds its way into 30,000 set-top boxes from Hitachi. JavaWorld: IDG’s magazine for the Java community, 3(7):??, July 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-07-1998/jw-07-hitachi.
                                        htm.

Steinberg:1998:SSM

[Ste98i] Jill Steinberg. Sun’s Scott McNealy and Alan Baratz discuss preliminary ruling in MS/Sun suit. Java-World: IDG’s magazine for the Java community, 3 (4):??, April 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-04-1998/jw-04-lawsuit.
                                        htm.

Sundsted:1996:HJJ

[Sun96a]  Todd Sundsted.  How-to

Java: Java and event handling. JavaWorld: IDG’s magazine for the Java community, 1(6):??, August 1996. CODEN ???? ISSN 1091-8906. URL http://

                                        www.javaworld.com/javaworld/
                                        jw-08-1996/jw-08-event.
                                        htm.

Sundsted:1996:HJO

[Sun96b]  Todd Sundsted.  How-

to Java: Observer and observable. JavaWorld: IDG’s magazine for the Java community, 1(9):??, November 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1996/jw-10-howto.
htm.

Sundsted:1996:HJE

[Sun96c]  Todd Sundsted.  How-

to Java: The effective user interface. JavaWorld: IDG’s magazine for the Java community, 1(7):??, September 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1996/jw-09-userint.
htm.

Sundsted:1996:IA

[Sun96d] Todd Sundsted. Introduction to the AWT. Java-World: IDG’s magazine for the Java community, 1 (5):??, July 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1996/jw-07-awt.htm.

Sundsted:1997:HJCb

[Sun97a]  Todd Sundsted.  How-to

Java: 3D computer graphics: Getting the hang of VRML. JavaWorld: IDG’s magazine for the Java community, 2 (8):??, August 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-08-1997/jw-08-howto.
html.

Sundsted:1997:HJCa

[Sun97b]  Todd Sundsted.  How-to

Java: 3D computer graphics, part 1: Model your world. JavaWorld: IDG’s magazine for the Java community, 2 (1):??, January 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
/javaworld/jw-05-1997/
jw-05-howto.html.

Sundsted:1997:HJB

[Sun97c]  Todd Sundsted.  How-to

Java: Build dynamically extensible applications. Java-World: IDG’s magazine for the Java community, 2(1): ??, January 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1997/jw-09-howto.
html.

Sundsted:1997:HJD

[Sun97d]  Todd Sundsted.  How-to

Java: Draw text the easy way. JavaWorld: IDG’s magazine for the Java community, 2(1):??, January 1997. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1997/jw-01-howto.
html.

Sundsted:1997:HJH

[Sun97e]  Todd Sundsted.  How-to

Java: How Java uses the producer/consumer model to handle images — an insider’s look. JavaWorld: IDG’s magazine for the Java community, 2(3):??, March 1997. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-03-1997/jw-03-howto.
html.

Sundsted:1997:HJL

[Sun97f]  Todd Sundsted.  How-

to Java: Learn how applets load network-based images asynchronously. Java-World: IDG’s magazine for the Java community, 2(2): ??, February 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-02-1997/jw-02-howto.
html.

Sundsted:1997:HJU

[Sun97g]  Todd Sundsted.  How-to

Java: Use the two ‘R’s of Java 1.1 — readers and writers. JavaWorld: IDG’s magazine for the Java community, 2(1):??, January 1997. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-11-1997/jw-11-howto.
html.

Sundsted:1997:HJWa

[Sun97h]  Todd Sundsted.  How-to

Java: Waging war on electronic junk mail. Java-World: IDG’s magazine for the Java community, 2(1): ??, January 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1997/jw-10-howto.
html.

Sundsted:1997:HJWb

[Sun97i]  Todd Sundsted.  How-

to Java: Write worldclass applications. Java-World: IDG’s magazine for the Java community, 2(12): ??, December 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-12-1997/jw-12-howto.
html.

Sundsted:1997:JTI

[Sun97j] Todd Sundsted. Java tips: Improve your applet’s download performance using Navigator’s Java console. Java-World: IDG’s magazine for the Java community, 2 (4):??, April 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javaworld/javatips/jw-
javatip28.html.

Sundsted:1998:ACT

[Sun98a] Todd Sundsted. Agents can think, too! JavaWorld: IDG’s magazine for the Java community, 3(10):??, 1998.

CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1998/jw-10-howto.
htm.

Sundsted:1998:ATA

[Sun98b]  Todd Sundsted.  Agents

talking to agents. Java-World: IDG’s magazine for the Java community, 3(9): ??, September 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1998/jw-09-howto.
htm.

Sundsted:1998:ASI

[Sun98c] Todd Sundsted. Application servers: An introduction. JavaWorld: IDG’s magazine for the Java community, 3(12):??, 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jbe/jw-jbe-appservers.
htm.

Sundsted:1998:AM

[Sun98d]  Todd Sundsted.  How-to

Java: Agents on the move. JavaWorld: IDG’s magazine for the Java community, 3(7):??, July 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1998/jw-07-howto.
htm.

Sundsted:1998:HJIb

[Sun98e]  Todd Sundsted.  How-to

Java: An introduction to agents. JavaWorld: IDG’s magazine for the Java community, 3(6):??, June 1998. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-06-1998/jw-06-howto.
                                        html.

Sundsted:1998:HJIa

[Sun98f]  Todd Sundsted.  How-to

Java: Internationalize dynamic messages. Java-World: IDG’s magazine for the Java community, 3(2): ??, February 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-02-1998/jw-02-howto.
                                        html.

Sundsted:1998:HJL

[Sun98g]  Todd Sundsted.  How-to

Java: Localize this! Java-World: IDG’s magazine for the Java community, 3(1): ??, January 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-01-1998/jw-01-howto.
                                        html.

Sundsted:1998:HJM

[Sun98h]  Todd Sundsted.  How-to

Java: MVC meets Swing. JavaWorld: IDG’s magazine for the Java community, 3 (4):??, April 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-04-1998/jw-04-howto.
html.

Sundsted:1998:HJU

[Sun98i]  Todd Sundsted.  How-

to Java: Unfurling Java’s data transfer API. Java-World: IDG’s magazine for the Java community, 3 (5):??, May 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-05-1998/jw-05-howto.
html.

Sundsted:1998:SDI

[Sun98j] Todd Sundsted. Signed and delivered: An introduction to security and authentication. JavaWorld: IDG’s magazine for the Java community, 3(12):??, 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-12-1998/jw-12-howto.
htm.

Sundsted:1998:ZYD

[Sun98k] Todd Sundsted. Zip your data and improve the performance of your network-based applications. JavaWorld: IDG’s magazine for the Java community, 3(11):??, 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-11-1998/jw-11-howto.
htm.

Sundsted:1999:ASI

[Sun99a] Todd Sundsted. Application servers: An introduction. JavaWorld: IDG’s magazine for the Java community, 4 (1):??, January 1999. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jbe/jw-jbe-appservers.
                                        htm.

Sundsted:1999:JWT

[Sun99b] Todd Sundsted. In Java we trust: A look at the Security API. Java-World: IDG’s magazine for the Java community, 4(1): ??, January 1999. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-01-1999/jw-01-howto.
                                        htm.

Sundsted:1999:JTP

[Sun99c] Todd Sundsted. Java technologies prove crucial to Enterprise solutions. Java-World: IDG’s magazine for the Java community, 4(1): ??, January 1999. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jbe/jw-jbe-javasolutions.
                                        htm.

Sundsted:1999:MHM

[Sun99d]  Todd Sundsted.  Mes-

saging helps move Java into the enterprise. Java-World: IDG’s magazine for the Java community, 4(1): ??, January 1999. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1999/jw-01-jms.htm.

Sykes:1996:PM

[Syk96a] Rebecca Sykes. Puzzle me this. JavaWorld: IDG’s magazine for the Java community, 1(1):??, March 1996. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-03-1996/jw-03-puzzle.
htm.

Sykes:1996:WGJ

[Syk96b]  Rebecca Sykes.  We go

to Java camp. Java-World: IDG’s magazine for the Java community, 1 (1):??, March 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-03-1996/jw-03-javacamp.
htm.

Sykes:1997:NCT

[Syk97] Rebecca Sykes. News: Caravelle throws lot in with Java in WebWatcher 5.0. Java-World: IDG’s magazine for the Java community, 2(10): ??, October 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1997/jw-10-idgns.
webwatcher.htm.

Sykes:1998:NJM

[Syk98] Rebecca Sykes. News: Java’s making inroads but Europe likes its HTML. Java-World: IDG’s magazine for the Java community, 3 (7):??, July 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-07-1998/jw-07-idgns-
                                        inroads.htm.

Taaffe:1998:NBL

[Taa98] Joanne Taaffe. News: Bull to launch Java smart card. JavaWorld: IDG’s magazine for the Java community, 3 (1):??, January 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-01-1998/jw-01-idgns.
                                        smartcard.htm.

Trott:1997:NMS

[TF97] Bob Trott and Andrew Foster. News: Microsoft and Sun post Java licensing contract on web. Java-World: IDG’s magazine for the Java community, 2(11): ??, November 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-11-1997/jw-11-infoworld.
                                        msmoves.htm.

Taaffe:1997:NSL

[TJ97] Joanne Taaffe and Margret Johnston. News: Siemens licenses Java for smart cards.

JavaWorld: IDG’s magazine for the Java community, 2 (8):??, August 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-08-1997/jw-08-idgns.
smartcards.htm.

Trott:1997:NDD

[Tro97a] Bob Trott. News: Developers decry Sun-Microsoft war over Java. Java-World: IDG’s magazine for the Java community, 2(11): ??, November 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-11-1997/jw-11-infoworld.
devs.htm.

Trott:1997:NMC

[Tro97b] Bob Trott. News: Microsoft cancels ‘Java Summit’ at last minute. Java-World: IDG’s magazine for the Java community, 2(12): ??, December 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-12-1997/jw-12-infoworld.
mscancels.htm.

Trott:1998:JDU

[Tro98a] Bob Trott. Justice Department uses Microsoft’s own words against it. Java-World: IDG’s magazine for the Java community, 3 (6):??, June 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.

                                        javaworld.com/javaworld/
                                        jw-06-1998/jw-06-iw-doj.
                                        htm.

Trott:1998:MCS

[Tro98b]  Bob Trott.  Microsoft

claims: Sun knew our Java plans all along. Java-World: IDG’s magazine for the Java community, 3(9): ??, September 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-09-1998/jw-09-iw-sunvsms.
                                        htm.

Trott:1997:NMW

[TS97] Bob Trott and Ed Scannell. News: Microsoft’s Willis says: ‘Java is not a standard’. JavaWorld: IDG’s magazine for the Java community, 2(5):??, May 1997. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-05-1997/jw-05-infoworld.
                                        msjava.standard.htm.

Uimonen:1996:NCA

[UG96] Terho Uimonen and Rob Guth. News: Comdex Asia: Gosling cools Java hype, lauds Microsoft’s efforts. JavaWorld: IDG’s magazine for the Java community, 1 (9):??, November 1996. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-10-1996/jw-10-asia.
                                        htm.

Valais:1996:NIL

[Val96] Alexandre Valais. News:

An inside look at the Java Cup International contest. JavaWorld: IDG’s magazine for the Java community, 1 (6):??, August 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-08-1996/jw-08-javacup.
htm.

vanderLinden:1996:HEW

[van96] Peter van der Linden. How exceptions work. Java-World: IDG’s magazine for the Java community, 1 (4):??, June 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-06-1996/jw-06-exceptions.
htm.

Vanhelsuwe:1997:BRJb

[Van97a] Laurence Vanhelsuwé. Book review: Java game programming books: A comparative review. Java-World: IDG’s magazine for the Java community, 2(9): ??, September 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1997/jw-09-bookreview.
htm.

Vanhelsuwe:1997:BRJc

[Van97b] Laurence Vanhelsuwé. Book review: Java network programming books: A comparative review. Java-World: IDG’s magazine for the Java community, 2(11): ??, November 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-11-1997/jw-11-bookreview.
htm.

Vanhelsuwe:1997:BRL

[Van97c] Laurence Vanhelsuwé. Book review: Lemay’s latest ‘Java 1.1’ title fails to deliver. JavaWorld: IDG’s magazine for the Java community, 2 (8):??, August 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-08-1997/jw-08-bookreview.
htm.

Vanhelsuwe:1997:BRJa

[Van97d] Laurence Vanhelsuwé. Book review: The Java Threads API makes it to print media. JavaWorld: IDG’s magazine for the Java community, 2(7):??, July 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1997/jw-07-threads.
htm.

Vanhelsuwe:1997:CSC

[Van97e] Laurence Vanhelsuwé. Cover story: Create your own supercomputer with Java. JavaWorld: IDG’s magazine for the Java community, 2 (1):??, January 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1997/jw-01-dampp.
htm.

Vanhelsuwe:1997:JPE

[Van97f] Laurence Vanhelsuwé. JavaBeans: properties, events, and thread safety. Java-World: IDG’s magazine for the Java community, 2(9): ??, September 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1997/jw-09-raceconditions.
htm.

Vanhelsuwe:1997:NGS

[Van97g] Laurence Vanhelsuwé. Need a good set of abstract data structures? ObjectSpace’s JGL packs a punch! Java-World: IDG’s magazine for the Java community, 2 (6):??, June 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-06-1997/jw-06-jgl.htm.

Vanhelsuwe:1998:BRD

[Van98a] Laurence Vanhelsuwé. Book review: Data structures and algorithms: A comparative slice and dice — er, review — of 5 Java DS&A books. JavaWorld: IDG’s magazine for the Java community, 3(6):??, June 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/

jw-06-1998/jw-06-bookreview.
htm.

Vanhelsuwe:1998:BRJb

[Van98b] Laurence Vanhelsuwé. Book review: Java virtual machine books — a comparative review. JavaWorld: IDG’s magazine for the Java community, 3(3):??, March 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-03-1998/jw-03-bookreview.
htm.

Vanhelsuwe:1998:BRJa

[Van98c] Laurence Vanhelsuwé. Book review: JDBC and Java database programming books: A comparative review. Java-World: IDG’s magazine for the Java community, 3(1): ??, January 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1998/jw-01-bookreview.
htm.

Vanhelsuwe:1998:IDM

[Van98d] Laurence Vanhelsuwé. It don’t mean a thing if it ain’t got that Swing: A comparative review of 4 JFC/Swing books. Java-World: IDG’s magazine for the Java community, 3(9): ??, September 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1998/jw-09-bookreview.
htm.

Vanhelsuwe:1998:SBF

[Van98e] Laurence Vanhelsuwé. Speed up batch file processing using Java Core Reflection and generic programming. JavaWorld: IDG’s magazine for the Java community, 3(11):??, 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-11-1998/jw-11-batch.
htm.

Vanhelsuwe:1999:BCF

[Van99] Laurence Vanhelsuwé. The battle of the container frameworks: which should you use? JavaWorld: IDG’s magazine for the Java community, 4 (1):??, January 1999. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1999/jw-01-jglvscoll.
htm.

Venners:1996:UHB

[Ven96a] Bill Venners. Under the hood: Bytecode basics. Java-World: IDG’s magazine for the Java community, 1(7): ??, September 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1996/jw-09-bytecodes.
htm.

Venners:1996:FPA

[Ven96b] Bill Venners. Under the hood: Floating-point arithmetic. JavaWorld: IDG’s magazine for the Java community, 1(9):??, November 1996. CODEN ???? ISSN 1091-8906. URL http://

www.javaworld.com/javaworld/
jw-10-1996/jw-10-hood.
htm.

Venners:1996:UHJb

[Ven96c] Bill Venners. Under the hood: Java’s garbagecollected heap. Java-World: IDG’s magazine for the Java community, 1(6): ??, August 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-08-1996/jw-08-gc.htm.

Venners:1996:UHJa

[Ven96d] Bill Venners. Under the hood: The Java class file lifestyle. JavaWorld: IDG’s magazine for the Java community, 1(5):??, July 1996. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1996/jw-07-classfile.
htm.

Venners:1996:UHL

[Ven96e] Bill Venners. Under the hood: The lean, mean, virtual machine. Java-World: IDG’s magazine for the Java community, 1 (4):??, June 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-06-1996/jw-06-vm.htm.

Venners:1997:UHCb

[Ven97a] Bill Venners. Under the hood: Closing the hood — A farewell to a longtime column. JavaWorld: IDG’s magazine for the Java community, 2(12): ??, December 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-12-1997/jw-12-hood.
html.

Venners:1997:UHCa

[Ven97b] Bill Venners. Under the hood: Control flow. Java-World: IDG’s magazine for the Java community, 2 (3):??, March 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-03-1997/jw-03-hood.
htm.

Venners:1997:UHHa

[Ven97c] Bill Venners. Under the hood: How the Java virtual machine handles exceptions. JavaWorld: IDG’s magazine for the Java community, 2 (1):??, January 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1997/jw-01-hood.
htm.

Venners:1997:UHHb

[Ven97d] Bill Venners. Under the hood: How the Java virtual machine handles method invocation and return. Java-World: IDG’s magazine for the Java community, 2 (6):??, June 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-06-1997/jw-06-hood.
htm.

Venners:1997:UHHc

[Ven97e] Bill Venners. Under the hood: How the Java virtual machine performs thread synchronization. Java-World: IDG’s magazine for the Java community, 2 (7):??, July 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1997/jw-07-hood.
htm.

Venners:1997:UHJb

[Ven97f] Bill Venners. Under the hood: Java security: How to install the security manager and customize your security policy. Java-World: IDG’s magazine for the Java community, 2(11): ??, November 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-11-1997/jw-11-hood.
htm.

Venners:1997:UHJa

[Ven97g] Bill Venners. Under the hood: Java’s security architecture. JavaWorld: IDG’s magazine for the Java community, 2(8):??, August 1997. CODEN ???? ISSN 1091-8906. URL http://

www.javaworld.com/javaworld/
jw-08-1997/jw-08-hood.
htm.

Venners:1997:UHSb

[Ven97h] Bill Venners. Under the hood: Security and the class loader architecture. Java-World: IDG’s magazine for the Java community, 2(9): ??, September 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1997/jw-09-hood.
htm.

Venners:1997:UHSc

[Ven97i] Bill Venners. Under the hood: Security and the class verifier. JavaWorld: IDG’s magazine for the Java community, 2(10): ??, October 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1997/jw-10-hood.
htm.

Venners:1997:UHSa

[Ven97j] Bill Venners. Under the hood: Solve your real-world problems with aglets. Java-World: IDG’s magazine for the Java community, 2 (5):??, May 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/

                                        jw-05-1997/jw-05-hood.
                                        htm.

Venners:1997:UHA

[Ven97k] Bill Venners. Under the hood: The architecture of aglets. JavaWorld: IDG’s magazine for the Java community, 2(4):??, April 1997. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-04-1997/jw-04-hood.
                                        htm.

Venners:1997:TFC

[Ven97l] Bill Venners. Under the hood: Try-finally clauses defined and demonstrated. JavaWorld: IDG’s magazine for the Java community, 2 (2):??, February 1997. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-02-1997/jw-02-hood.
                                        htm.

Venners:1998:COI

[Ven98a] Bill Venners. The ’canonical object’ idiom. JavaWorld: IDG’s magazine for the Java community, 3(10):??, 1998. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-10-1998/jw-10-techniques.
                                        htm.

Venners:1998:CSD

[Ven98b] Bill Venners. Cover story: Designing with exceptions. JavaWorld: IDG’s magazine for the Java community, 3(7):??, July 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1998/jw-07-techniques.
htm.

Venners:1998:DTS

[Ven98c]  Bill Venners.  Design

for thread safety. Java-World: IDG’s magazine for the Java community, 3(8): ??, August 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-08-1998/jw-08-techniques.
htm.

Venners:1998:DTD

[Ven98d]  Bill Venners.  Design

techniques: Designing object initialization. Java-World: IDG’s magazine for the Java community, 3 (3):??, March 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-03-1998/jw-03-techniques.
htm.

Venners:1998:DTI

[Ven98e] Bill Venners. Design techniques: Introduction to the new column. Java-World: IDG’s magazine for the Java community, 3(2): ??, February 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-02-1998/jw-02-techniques.
htm.

Venners:1998:DTK

[Ven98f] Bill Venners. Design techniques: Keep focused! designing fields and methods. JavaWorld: IDG’s magazine for the Java community, 3 (4):??, April 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-04-1998/jw-04-techniques.
                                        htm.

Venners:1998:DTO

[Ven98g] Bill Venners. Design techniques: Object finalization and cleanup. Java-World: IDG’s magazine for the Java community, 3 (6):??, June 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-06-1998/jw-06-techniques.
                                        htm.

Venners:1998:WMD

[Ven98h] Bill Venners. Design techniques: What’s a method to do. JavaWorld: IDG’s magazine for the Java community, 3(5):??, May 1998. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-05-1998/jw-05-techniques.
                                        htm.

Venners:1998:DIO

[Ven98i] Bill Venners. Designing with interfaces: One programmer’s struggle. JavaWorld: IDG’s magazine for the Java community, 3(12):??, 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-12-1998/jw-12-techniques.
htm.

Venners:1998:EGI

[Ven98j] Bill Venners. The ‘event generator’ idiom. Java-World: IDG’s magazine for the Java community, 3(9): ??, September 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1998/jw-09-techniques.
htm.

Venners:1998:EJT

[Ven98k] Bill Venners. Exceptions in Java: A tutorial. Java-World: IDG’s magazine for the Java community, 3 (7):??, July 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1998/jw-07-exceptions.
htm.

Venners:1998:IVC

[Ven98l] Bill Venners. Inheritance versus composition: Which one should you choose? JavaWorld: IDG’s magazine for the Java community, 3(11):??, 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-11-1998/jw-11-techniques.
htm.

Venners:1998:JTB

[Ven98m] Bill Venners. Jini technology, out of the box. JavaWorld: IDG’s magazine for the Java community, 3(12):??, 1998. CODEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jbe/jw-jbe-jini.htm.

Venners:1998:NIS

[Ven98n] Bill Venners. News: Impressions of SD ’98. Java-World: IDG’s magazine for the Java community, 3 (3):??, March 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-03-1998/jw-03-sd98.
                                        venners.htm.

Venners:1998:OIJ

[Ven98o] Bill Venners. Object initialization in Java. Java-World: IDG’s magazine for the Java community, 3 (3):??, March 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-03-1998/jw-03-initialization.
                                        htm.

Venners:1998:SON

[Ven98p]  Bill Venners.  Sun out-

lines new licensing model for Java platform standard extensions. JavaWorld: IDG’s magazine for the Java community, 3(12):??, 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.

javaworld.com/javaworld/
jbe/jw-jbe-stand-ext.htm.

Venners:1999:DDE

[Ven99a] Bill Venners. Design with dynamic extension. Java-World: IDG’s magazine for the Java community, 4(1): ??, January 1999. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1999/jw-01-techniques.
htm.

Venners:1999:JTB

[Ven99b] Bill Venners. Jini technology, out of the box. Java-World: IDG’s magazine for the Java community, 4(1): ??, January 1999. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jbe/jw-jbe-jini.htm.

Venners:1999:SJV

[Ven99c] Bill Venners. The state of the Java virtual machine (JVM). JavaWorld: IDG’s magazine for the Java community, 4 (1):??, January 1999. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jbe/jw-jbe-jvm.htm.

Venners:1999:SON

[Ven99d]  Bill Venners.  Sun out-

lines new licensing model for Java platform Standard Extensions. Java-World: IDG’s magazine for the Java community, 4(1): ??, January 1999. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jbe/jw-jbe-stand-ext.htm.

Vizard:1998:SOM

[VF98] Matthew Vizard and Emily Fitzloff. Sun and ObjectSpace make Jini plans. JavaWorld: IDG’s magazine for the Java community, 3 (9):??, September 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1998/jw-09-iw-objectspace.
htm.

vanHoff:1996:JCA

[vW96] Arthur van Hoff and Kathy Walrath. Javamation: Creating animation in Java applets. JavaWorld: IDG’s magazine for the Java community, 1(1):??, March 1996. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-03-1996/jw-03-animation.
htm.

Webb:1998:JJ

[Web98] Todd A. Webb. To jar or not to jar. JavaWorld: IDG’s magazine for the Java community, 3(7):??, July 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1998/jw-07-jar.htm.

Weintraub:1996:BRC

[Wei96]  Jacob Weintraub.  Book

review: Core Java not just for experts. Java-World: IDG’s magazine for the Java community, 1(7): ??, September 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1996/jw-09-corejava.
review.htm.

Weil:1998:MLN

[Wei98]  Nancy Weil.  Marimba

launches new Castanet with a less Java-centric focus. JavaWorld: IDG’s magazine for the Java community, 3(7):??, July 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1998/jw-07-idgns-
marimba.htm.

Westerkamp:1996:ASD

[Wes96] Eric Westerkamp. Automatic software distribution of Java applications. Java-World: IDG’s magazine for the Java community, 1 (3):??, May 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-05-1996/jw-05-westerkamp.
htm.

Willms:1998:GRE

[Wil98] Roland Willms. Get rid of those empty black squares in your text editor. JavaWorld:

IDG’s magazine for the Java community, 3(12):??, 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javatips/jw-javatip64.
htm.

Winslow:1997:SSC

[Win97a] Maria Winslow. 1java step by step: Chart your way to custom graph components. JavaWorld: IDG’s magazine for the Java community, 2(7):??, July 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-07-1997/jw-07-step.
htm.

Winslow:1997:CCC

[Win97b] Maria Winslow. Java step by step: Creating custom components. JavaWorld: IDG’s magazine for the Java community, 2(4):??, April 1997. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-04-1997/jw-04-step.
htm.

Wisniewski:1997:JTT

[Wis97] Thomas Wisniewski. Java tip #37: The trick to using Java networking applets behind firewalls. Java-World: IDG’s magazine for the Java community, 2(9): ??, September 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/

javatips/jw-javatip37.
htm.

Mitchell:1998:ERP

[wJDM98] Nige Warren with John D. Mitchell. Explicit and recursive programming in Java. JavaWorld: IDG’s magazine for the Java community, 3(10):??, 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javatips/jw-javatip62.
htm.

Wah:1997:JTH

[WM97a] Chong Ser Wah and John D.

Mitchell. Java tips: How to play audio in applications. JavaWorld: IDG’s magazine for the Java community, 2 (2):??, February 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javatips/jw-javatip24.
htm.

West:1997:JTI

[WM97b] Jeff West and John D.

Mitchell. Java tip 43: Instruction on loading bitmap files in a Java app. Java-World: IDG’s magazine for the Java community, 2(12): ??, December 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javatips/jw-javatip43.
htm.

Woollacott:1997:NNL

[Woo97] Matthew Woollacott. News: Novell to launch Javabased initiatives. Java-World: IDG’s magazine for the Java community, 2(9): ??, September 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-09-1997/jw-09-infoworld.
novell.htm.

Giorgio:1998:WOC

[wRDG98] Thomas Schaeck with Rinaldo Di Giorgio. Write OpenCard card services for Java Card applets. JavaWorld: IDG’s magazine for the Java community, 3(10):??, 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1998/jw-10-javadev.
htm.

Wendler:1999:HWC

[WwRDGSB99] Mike Wendler and with Rinaldo Di Giorgio Stephan Breideneich. How to write a CardTerminal class for simple and complex readers in an OpenCard environment. JavaWorld: IDG’s magazine for the Java community, 4 (1):??, January 1999. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1999/jw-01-javadev.
htm.

Yourdon:1996:FJN

[You96] Edward Yourdon. Feature:

Java and the new Internet programming paradigm. JavaWorld: IDG’s magazine for the Java community, 1 (6):??, August 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-08-1996/jw-08-yourdonbook.
htm.

Zhang:1997:JTH

[Zha97] Nick Zhang. Java tips: How to improve Java’s I/O performance. JavaWorld: IDG’s magazine for the Java community, 2(3):??, March 1997. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javatips/jw-javatip26.
htm.

Zukowski:1996:JTC

[Zuk96]  John Zukowski.  Java

tips: Creating in-memory images in Java. Java-World: IDG’s magazine for the Java community, 1(9): ??, November 1996. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javatips/jw-javatip16.
htm.

Zukowski:1997:APP

[Zuk97a] John Zukowski. Avoid potential pitfalls of Microsoft’s non-standard SDK for Java. JavaWorld: IDG’s magazine for the Java community, 2 (10):??, October 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1997/jw-10-sdk.htm.

Zukowski:1997:CSH

[Zuk97b]  John Zukowski.  Cover

story: How to avoid potential pitfalls of Microsoft’s non-standard SDK for Java. JavaWorld: IDG’s magazine for the Java community, 2 (11):??, November 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-11-1997/jw-11-pitfalls.
htm.

Zukowski:1997:FHA

[Zuk97c]  John Zukowski.  Follow-

up: ‘how to avoid potential pitfalls of Microsoft’s non-standard SDK for Java’. JavaWorld: IDG’s magazine for the Java community, 2 (12):??, December 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-12-1997/jw-12-pitfalls.
update.htm.

Zukowski:1997:JCD

[Zuk97d] John Zukowski. JavaOne complete: from day 0 to day last +1. JavaWorld: IDG’s magazine for the Java community, 2(5):??, May 1997. CODEN ???? ISSN 1091-

8906. URL http://www.

javaworld.com/javaworld/
jw-05-1997/jw-05-javaone-
zukowski.htm.

Zukowski:1997:PAR

[Zuk97e] John Zukowski. ‘pitfalls’ article receives attention from JavaWorld readers. Java-World: IDG’s magazine for the Java community, 2(12): ??, December 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-12-1997/jw-12-pitfalls.
letters.htm.

Zukowski:1997:WDSa

[Zuk97f] John Zukowski. What does Sun’s lawsuit against Microsoft mean for Java developers? JavaWorld: IDG’s magazine for the Java community, 2(1):??, January 1997. CODEN ???? ISSN 1091-8906. URL http://

www.javaworld.com/javaworld/
jw-10-1997/jw-10-lawsuit.
html.

Zukowski:1997:WDSb

[Zuk97g] John Zukowski. What does Sun’s lawsuit against Microsoft mean for Java developers? JavaWorld: IDG’s magazine for the Java community, 2(11):??, November 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-10-1997/jw-10-lawsuit.
htm.

Zukowski:1998:IKS

[Zuk98a]  John Zukowski.  At in-

prise keynote, Sun’s Alan Baratz makes release-date announcements. Java-World: IDG’s magazine for the Java community, 3(8): ??, August 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-08-1998/jw-08-inprise.
                                        htm.

Zukowski:1998:ICR

[Zuk98b]  John Zukowski.  Inprise

Conference reveals juicy bits of Java news. Java-World: IDG’s magazine for the Java community, 3(9): ??, September 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        jw-09-1998/jw-09-inprisefull.
                                        htm.

Zukowski:1998:JTUb

[Zuk98c] John Zukowski. Java tip 46: Use Java 1.2’s Authenticator class. Java-World: IDG’s magazine for the Java community, 3(2): ??, February 1998. CO-DEN ???? ISSN 1091-

                                        8906. URL http://www.
                                        javaworld.com/javaworld/
                                        javatips/jw-javatip46.
                                        htm.

Zukowski:1998:JTUa

[Zuk98d]  John Zukowski.  Java

tip 47: URL authentication revisited. Java-World: IDG’s magazine for the Java community, 3(2): ??, February 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
javatips/jw-javatip47.
htm.

Zukowski:1998:MSJ

[Zuk98e] John Zukowski. Microsoft sheds its Java skin — again. JavaWorld: IDG’s magazine for the Java community, 3 (8):??, August 1998. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-08-1998/jw-08-pitfalls-
update.htm.

Zukowski:1998:PYW

[Zuk98f] John Zukowski. Prepare yourself for what’s new and different in the forthcoming JDK 1.2 release. JavaWorld: IDG’s magazine for the Java community, 3(11):??, 1998. CODEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-11-1998/jw-11-jdk12.
htm.